### v3

# Reweighted Spectral Partitioning Works: Bounds for Special Graph Classes

Abstract
Spectral partitioning is a method that can be used to compute small sparse cuts or small edge-separators in a wide variety of graph classes, by computing the second-smallest eigenvalue (and eigenvector) of the Laplacian matrix. Upper bounds on this eigenvalue for certain graph classes imply that the method obtains small edge-separators for these classes, usually with a sub-optimal dependence on the maximum degree. In this work, we show that a related method, called reweighted spectral partitioning, guarantees near-optimal sparse vertex-cuts and vertex-separators in a wide variety of graph classes. In many cases, this involves little-to-no necessary dependence on maximum degree.
We also obtain a new proof of the planar separator theorem, a strengthened eigenvalue bound for bounded-genus graphs, and a refined form of the recent Cheeger-style inequality for vertex expansion via a specialized dimension-reduction step.
1 Introduction
Partitioning a graph into smaller pieces is a fundamental type of problem that arises in a variety of fields. Applications of graph partitioning include divide and conquer algorithms, dynamic programming algorithms [RH01], graph drawing [BPP22], geometrically embedding graphs [SSX25], FPGA compilation [CCW21], image processing, VLSI design, compute parallelization, bioinformatics, simulation, road networks, social networks [BMS+16], graph processing [OFS17], and even air traffic control methods [FGBJ17].
Many of these applications have similar ways of measuring partitioning quality, giving rise to some notable well-studied objectives, such as various forms of “expansion” and “separators”, each of which is a class of objectives with a variety of theoretical applications. In their most fundamental form, these objective types are studied in the sense of a single “cut”, partitioning the graph into two parts.
Fix a graph $G=(V,E)$ with $n$ vertices. For a subset $A\subset V$ with $|A|\leq n/2$, let $\delta(A)$ denote the set of edges with one endpoint in $A$ and one endpoint in $V\setminus A$. Likewise, let $\partial A$ denote the set of vertices in $V\setminus A$ adjacent to some vertex in $A$. The edge expansion of $A$ is the ratio $\frac{|\delta(A)|}{|A|}$. The vertex expansion of $A$ is the ratio $\frac{|\partial(A)|}{|A|}$. Due to this form, $A$ is often called a ratio cut in both cases, and a bounded-ratio cut if only an upper bound on the expansion is known. The edge or vertex expansion of $G$ is defined as the minimum edge or vertex expansion over all possible choices of $A$. The edge expansion of $G$ is denoted $\phi(G)$, and the vertex expansion of $G$ is denoted $\psi(G)$. Edge and vertex expansion are both NP-hard to compute exactly [Kai04] (see Appendix B for vertex expansion). Moreover, for a graph with vertex expansion $\psi$ and maximum degree $\Delta$, it is hard to find a subset with vertex expansion less than $C\sqrt{\psi\log\Delta}$ for an absolute constant $C$, assuming the so-called “small-set expansion hypothesis” [LRV13], which was suggested by Raghavendra and Steurer [RS10].
[…]
{TAB}
(5,3pt,3pt)[5pt]—l—c—c—c——c—c—c—c—c—c—c— Graph class Separator Size Time Refs.
Planar $\operatorname{\mathcal{O}}(\sqrt{n})$ $\operatorname{\mathcal{O}}(n)$ [Ung51, LT79]
[…]
$K_{h}$-minor-free $\operatorname{\mathcal{O}}(h\sqrt{n})$ $\operatorname{\mathcal{O}}(n^{1+\varepsilon})$ [^2] Assumes $h$ is a constant, and uses a tower function in terms of $h$. Note that some more tractable algorithms are known with weaker guarantees [WN11]. [KR10]
RIG over $K_{h}$-minor-free $\operatorname{\mathcal{O}}\!\bigl{(}h^{3}\sqrt{m\log h}\bigr{)}$ $\text{poly}(n)$ [^3] Solves a sequence of $O(n)$ linear programs, each with approximately $n^{2}$ variables and $n^{3}$ (sparse) constraints (exact number of each depends on formulation, but is always polynomial). [Lee17]
$k$-ply neighbourhood system in $\mathbb{R}^{d}$ $\operatorname{\mathcal{O}}(dk^{1/d}n^{1-1/d})$ $O(nd^{2})$ [^4] Assumes an embedding of the graph in $\mathbb{R}^{d}$ is given. [MTTV97]
$k$-nearest-neighbour graph in $\mathbb{R}^{d}$ $\operatorname{\mathcal{O}}(dk^{1/d}n^{1-1/d})$ $O(nd^{2})$ [^4] Assumes an embedding of the graph in $\mathbb{R}^{d}$ is given. [MTTV97]
Table 1: Best-known sizes of 2/3 2/32 / 3-vertex-separators for several graph classes, as well as the time required to compute the separator. Many of these results require additional information besides the graph to compute the separator in the stated time complexity (see the footnotes). $n$ is always the number of vertices, and $m$ is always the number of edges.
These two types of objectives are intimately connected. An edge or vertex-separator $S$ itself forms either $S=\delta(A)$ or $S=\partial(A)$ (respectively) for some set $|A|\leq n/2$, so a graph with an edge/vertex-separator $S$ has edge/vertex-expansion at most $O\left(|S|/n\right)$. A converse form of this is also often true. In particular, it can be shown by a standard recursive argument that if every induced subgraph $H\subset G$ has edge/vertex expansion at most $\kappa$, then $G$ has an edge/vertex separator of size $O(\kappa n)$ (see e.g. [ST96, ST07, Lemma A.1]).
1.1 Algorithms for General Graphs
Although all types of expansion and separators we have mentioned are NP-hard to compute in general, a number of different algorithms with guarantees are known. These algorithms use no special information about the graph, but still have guarantees for many graph classes.
1.1.1 Spectral Partitioning
Most relevant to this work is the method of spectral partitioning, which is the algorithm induced by the so-called “Cheeger inequality”, which is a well-known result in Spectral Graph Theory. For an undirected graph $G$ with $n$ vertices, let its adjacency matrix (or weight matrix) be denoted $A$, and the diagonal matrix formed from its degrees (or sums of incident weights) be denoted $D$. The Laplacian matrix of $G$ is $L(G):=D-A$. We denote the eigenvalues of $L(G)$ as $0=\lambda_{1}(G)\leq\lambda_{2}(G)\leq\lambda_{3}(G)\leq\cdots\leq\lambda_{n}(G)$. The study of these eigenvalues, and their relationships to combinatorial quantities, is a primary focus of the field of Spectral Graph Theory. The second-smallest Laplacian eigenvalue in particular ($\lambda_{2}(G)$) was dubbed the algebraic connectivity of a graph by Fiedler [Fie73], and is sometimes called the Fiedler value of a graph. As the name would suggest, algebraic connectivity has a relationship to a combinatorial form of connectivity. Namely, edge expansion:
Theorem 1.1 (Cheeger inequality for edge expansion [AM85, Moh89, JS88]).
For a graph $G$ with maximum degree $\Delta$,
$\frac{\phi(G)^{2}}{2\Delta}\leq\lambda_{2}(G)\leq 2\phi(G).$
It is known that both sides of this inequality are asymptotically tight [HLW06]. A key feature of this theorem’s proof is that it is algorithmically constructive — the eigenvector associated with $\lambda_{2}(G)$ can be converted to a binary vector defining a cut via a threshold test, obtaining a sparse edge cut, giving the upper bound on edge expansion. $\lambda_{2}(G)$ can be computed (approximately) alongside a corresponding eigenvector in near-linear ($\tilde{\operatorname{\mathcal{O}}}(m)$) time [KMP11]. This is the algorithm called spectral partitioning. Applying this successively ($\operatorname{\mathcal{O}}(n)$ times), it can also be used to obtain edge-separators in $\tilde{\operatorname{\mathcal{O}}}(nm)$ time [ST96, ST07]. This algorithm is simple and fast, and, in fact, has guaranteed results (i.e., upper bounds on the sizes of the resulting separators) for many graph classes. These guarantees follow from upper bounds on $\lambda_{2}(G)$ for the classes, which have been found for a large number of graph classes (see Table 2, and see Appendix A for the definition of each graph class). Somewhat surprisingly, these bounds were all found after spectral partitioning had already seen widespread use as a heuristic.
{TAB}
(5,3pt,3pt)[5pt]—l—c—c——c—c—c—c—c—c—c—c— Graph class $\lambda_{2}(G)\lesssim$ Refs.
Planar $\dfrac{\Delta}{n}$ [ST96, ST07]
[…]
Triangulated [^1] Due to a subtle issue, this bound does not apply to general $g$ graphs, which is what the references claim. We will amend this issue as part of our work. $g$ $\dfrac{\text{poly}(\Delta)\cdot g}{n}$ [Kel04, Kel06b, Kel06a]
$K_{h}$-minor-free $\dfrac{\Delta\,(h\log h\log\log h)^{2}}{n}$ [BLR10, CF25, DP25]
RIG over $K_{h}$-minor-free $\dfrac{\Delta^{2}\,h^{6}\log h}{n}$ [Lee17]
$k$-ply neighbourhood system in $\mathbb{R}^{d}$ $\Delta\left(\frac{k}{n}\right)^{\frac{2}{d}}$ [ST96, ST07]
$k$-nearest-neighbour graph in $\mathbb{R}^{d}$ $\Delta\left(\frac{k}{n}\right)^{\frac{2}{d}}$ [ST96, ST07]
Table 2: Best-known asymptotic upper bounds on $\lambda_{2}(G)$ for several graph classes. $n$ is always the number of vertices, and $\Delta$ is always the maximum degree. The notation $a\lesssim b$ means $a=\operatorname{\mathcal{O}}(b)$.
A downside of spectral partitioning is its dependence on the maximum degree $\Delta$ of the given graph. For instance, the resulting edge-separator size in planar graphs has a sub-optimal dependence on $\Delta$ compared to the best-known: it results in an edge-separator of size $\operatorname{\mathcal{O}}(\Delta\sqrt{n})$, but it is known that edge-separators exist of size $\operatorname{\mathcal{O}}(\sqrt{\Delta n})$ [Mil84, DDSV93]. In fact, this weaker bound is identical to what’s implied by the planar separator theorem – one can take the vertex separator of size $\operatorname{\mathcal{O}}(\sqrt{n})$ given by the planar separator theorem, and form an edge separator of size $\operatorname{\mathcal{O}}(\Delta\sqrt{n})$ from the edges incident to each vertex. This (very weakly) suggests that the techniques underpinning the eigenvalue bound could be more directly related to vertex-separators than edge-separators, in a more direct manner than what’s currently known.
1.1.2 Reweighted Spectral Partitioning
Since a larger value of $\lambda_{2}$ is associated with higher connectivity (and some other properties), one interesting question that could be asked is: Given a graph $G$, what is the weight distribution over the edges (under some normalization constraints) that maximizes $\lambda_{2}$? Intuitively, in some cases, this might let us reweigh the edges of a graph to improve its connectivity. This question was first posed by Boyd, Diaconis and Xiao, who recognized it as a variant of the fastest-mixing Markov chain problem [BDX04]. They refer to the extremal value as $\lambda_{2}^{*}(G,\pi)$ for an undirected connected graph $G$, and a probability mass function $\pi:V\to[0,1]$ of the vertices, and it is formulated as the following optimization problem (also called the maximum reweighted spectral gap):
$\begin{array}[]{r c c r c l l}\lambda_{2}^{*}(G,\pi)&:=&\max_{P\geq 0}&\lambda_{2}(I-P)\\ &&\textup{subject to}&P(u,v)&=&0&\forall uv\not\in E\cup\{vv:v\in V(G)\}\\ &&&\sum_{v\in V}P(u,v)&=&1&\forall u\in V\\ &&&\pi(u)P(u,v)&=&\pi(v)P(v,u)&\forall uv\in E\cup\{vv:v\in V(G)\}\end{array}$
$P$ is a reweighted adjacency matrix of $G$ (with self-loops added) normalized to represent a reversible Markov chain with stationary distribution $\pi$. Hence, $I$ is the weighted degree-matrix for the weighted graph, $I-P$ is the Laplacian matrix of the weighted graph, and $\lambda_{2}(I-P)$ is its second-smallest eigenvalue. We will restrict ourselves to discussing the case where $\pi$ is the uniform distribution, which we denote simply as $\lambda_{2}^{*}(G)$. Before moving on, we make an important observation:
Observation 1.2.
For a graph $G$ with maximum degree $\Delta$, $\lambda_{2}(G)\leq\Delta\cdot\lambda_{2}^{*}(G)$.
Boyd, Diaconis, and Xiao also realize their optimization problem as a Semidefinite Program meeting Slater’s condition (therefore allowing it to be approximately solved in polynomial time via e.g. [ZL18]), and in the process they derive a dual formulation (better stated and properly extracted in follow-up work by Roch [Roc05]):
$\begin{array}[]{r c c r c l l}\gamma^{(d)}(G)&:=&\min_{\begin{subarray}{c}f:V\to\mathbb{R}^{d}\\ y:V\to\mathbb{R}_{\geq 0}\end{subarray}}&\sum_{v\in V}y(v)\\ &&\textup{subject to}&\sum_{v\in V}f(v)&=&\hskip 10.00002pt\overline{0}\\ &&&y(u)+y(v)&\geq&\hskip 10.00002pt\frac{||f(u)-f(v)||_{2}^{2}}{\sum_{x\in V}||f(x)||_{2}^{2}}&\forall uv\in E\end{array}$ (1)
A strong duality proof gives that $\lambda_{2}^{*}(G)=\gamma^{(n)}(G)$. Note that we have intentionally added the choice of dimension $d$ (rather than fixing $d=n$) for later use. Other methods of computing $\lambda_{2}^{*}(G)$ include a particularly simple and practical sub-gradient method [BDX04] and a matrix-multiplicative weight-update method that runs in near-linear time if $\lambda_{2}^{*}$ is very large and weighted in a particular way [LTW24, Section 5].
One useful bound on $\lambda_{2}^{*}(G)=\gamma^{(n)}\leq\gamma^{(1)}(G)$ is:
Observation 1.3.
For any graph $G$, $\gamma^{(1)}(G)\lesssim 1$.
Proof.
If $n$ is even, assign $y$ to be $\frac{2}{n}$ for every vertex, and assign $f$ to be 1-1- 1 for half the vertices and 111 for the other half of the vertices, so then $\gamma^{(1)}(G)\leq 2\lesssim 1$. If $n$ is odd (and at least 333), assign $y$ to be $\frac{2}{n-1}$ for every vertex, and assign $f$ to be an even split of 1-1- 1 and 111 for all but one vertex, which is assigned 0, resulting in $\gamma^{(1)}(G)\leq 2+\frac{2}{n-1}\lesssim 1$. ∎
A series of work has recently resulted in a Cheeger-style inequality relating vertex expansion with $\lambda_{2}^{*}$:
Theorem 1.4 (Cheeger inequality for vertex expansion [Roc05, OTZ22, JPV22, KLT22]).
For an undirected graph $G$ with maximum degree $\Delta$, and any probability distribution $\pi$ on $V$,
$\frac{\psi(G)^{2}}{\log\Delta}\lesssim\lambda_{2}^{*}(G)\lesssim\psi(G).$
This is the single most important known theorem involving $\lambda_{2}^{*}$. Like the proof of Cheeger’s inequality, Theorem 1.4 is also algorithmically constructive, and the implied randomized polynomial-time algorithm that converts a solution to $\lambda_{2}^{*}(G)$ (or more specifically, $\gamma^{(n)}(G)$) into a bounded-ratio vertex-cut upper-bounding $\psi(G)$ is the family of methods we call reweighted spectral partitioning. We call this a “family” of methods since a few different methods for each step of the proof (with slightly different constants) have been given. Moreover, we will provide yet another construction for one step with slightly different bounds later in this work to arrive at a “refined” Cheeger inequality for vertex expansion in Section 5.
The “easy” direction Theorem 1.4 was shown by Roch [Roc05]. A weaker form of the hard direction was given by Olesker-Taylor and Zanetti [OTZ22], which was quickly refined by Jain, Pham, and Vuong [JPV22], and independently by Kwok, Lau, and Tung [KLT22] (who also handled the weighted case, which we omit). Our outline will follow the techniques of Kwok, Lau, and Tung. There are two separate results that together imply both sides of the inequality:
Theorem 1.5 ([OTZ22, Theorem 2.10]).
For a graph $G$,
$\psi(G)^{2}\lesssim\gamma^{(1)}(G)\lesssim\psi(G).$
This result is generalized by Kwok, Lau, and Tung for the weighted case [KLT22, Theorem 3.15]. Their proof is notable for our case of uniform weights too, since it implies a very simple and practical algorithm for this step.
Theorem 1.6 ([KLT22, Proposition 3.14]).
For a graph $G$ with maximum degree $\Delta$,
$\gamma^{(n)}(G)\lesssim\gamma^{(1)}(G)\lesssim\gamma^{(n)}(G)\cdot\log\Delta.$
The “hard” direction of Theorem 1.6 essentially uses a (lossy) dimension-reduction step, which is also quite simple algorithmically. We note for later that there is also an easier dimension-reduction step that is useful in some cases for smaller initial dimension $d$: Specifically, it can be shown that $\gamma^{(1)}(G)\lesssim d\gamma^{(d)}(G)$ [OTZ22, Proof of Proposition 2.9].
Unlike $\lambda_{2}$, no non-trivial bounds on $\lambda_{2}^{*}$ are yet known, and hence reweighted spectral partitioning is not yet known to produce good separators. This is the primary focus of our work.
1.1.3 Other Methods
The guarantees obtained by spectral partitioning are of an unusual form compared to typical approximation algorithms. Some more typical results are known. In particular, for a general graph with $n$ vertices, and $m\geq n$ edges, an $\operatorname{\mathcal{O}}(m^{1+\varepsilon})$-time algorithm (for any $\varepsilon>0$) is known that can obtain an $\operatorname{\mathcal{O}}(\sqrt{\log n})$-approximation of either edge or vertex expansion [LTW24]. However, it is quite impractical due to the use of almost-linear time min-cost flow [CKL+22b, CKL+22a]. This result strengthened many other prior results obtaining slightly weaker bounds or slower algorithms, although a slightly more refined approximations is still possible in polynomial time for many cases [FHL05, FHL08] (we will briefly discuss this towards the end of our paper, since we will use some related techniques).
1.2 Results and Organization
We will give bounds on $\lambda_{2}^{*}$, $\gamma^{(1)}$, and $\gamma^{(d)}$ for a large variety of graph classes, summarized in Table 3 and Table 4. In turn, these imply bounds on (vertex) separator sizes, as well as bounds on the resulting separators from reweighted spectral partitioning, summarized in Table 5. This also slightly improves the dependence on dimension for separator bounds in some geometric graph classes. Most of our bounds on $\lambda_{2}^{*}$ will primarily adapt existing machinery of bounds for $\lambda_{2}$. However, for $g$ graphs, we are able to make some notable improvements to the techniques, and our bound on $\lambda_{2}^{*}$ also implies a new bound on $\lambda_{2}$. These techniques are the topic of Section 3.2. The main new tool is the use of something we call “uniform” shallow minors, although these are used in a very different manner compared to previous work leveraging forbidden shallow minors.
{TAB}
(2,1pt,1pt)[4pt]—c—c——c—c—c—c—c— Graph class $\frac{\lambda_{2}}{\Delta}\leq\lambda_{2}^{*}\leq\gamma^{(1)}\lesssim$
Planar $\dfrac{1}{n}$
[…]
$K_{h}$-minor-free $\dfrac{(h\log h\log\log h)^{2}}{n}$
RIG over $K_{h}$-minor-free $\dfrac{\Delta\,h^{6}(\log\log h)^{2}}{n}$
Table 3: Our asymptotic upper bounds on $\gamma^{(1)}$ and $\lambda_{2}^{*}$ for several graph classes given in Corollary 4.9, Theorem 3.5 and Theorem 3.17.
{TAB}
(2,1pt,1pt)[4pt]—c—c——c—c—c— Graph class $\frac{\lambda_{2}}{\Delta}\leq\lambda_{2}^{*}\leq\gamma^{(d)}\lesssim$
$k$-ply neighbourhood system in $\mathbb{R}^{d}$ $\left(\frac{k}{n}\right)^{\frac{2}{d}}$
$k$-nearest-neighbour graph in $\mathbb{R}^{d}$ $\left(\frac{k}{n}\right)^{\frac{2}{d}}$
Table 4: Our asymptotic upper bounds on $\gamma^{(d)}$ and $\lambda_{2}^{*}$ for geometric graph classes, given in Section 3.1.
{TAB}
(3,1pt,1pt)[4pt]—c—c—c——c—c—c—c—c—c—c— Graph class Separator size $\lesssim$ Poly-time separator size $\lesssim$
Planar $\sqrt{n}$ $\sqrt{n}$
[…]
$K_{h}$-minor-free $(h\log h\log\log h)\sqrt{n}$ $\min\{\log h,\sqrt{\log\Delta}\}(h\log h\log\log h)\sqrt{n}$
RIG over $K_{h}$-minor-free $h^{3}\log\log h\sqrt{n\Delta}$ $h^{3}\log\log h\sqrt{n\Delta\log\Delta}$
$k$-ply neighbour -hood system in $\mathbb{R}^{d}$ $\sqrt{\min\{d,\log\Delta\}}\cdot k^{1/d}n^{1-1/d}$ $\sqrt{\log\Delta}\cdot k^{1/d}n^{1-1/d}$
$k$-nearest-neigh -bour graph in $\mathbb{R}^{d}$ $\sqrt{\min\{d,\log\Delta\}}\cdot k^{1/d}n^{1-1/d}$ $\sqrt{d+\log k}\cdot k^{1/d}n^{1-1/d}$
Table 5: From the bounds on $\lambda_{2}^{*}$ stated in Table 3 and Table 4, we obtain (asymptotic) separator sizes via Theorem 1.5, and polynomial-time computable (asymptotic) separator sizes obtainable via reweighted spectral partitioning as stated in Theorem 5.3.
Overall, the techniques for these bounds can be categorized into “geometric intersection graph” techniques, which we consider in Section 3, and “metric” techniques, which we consider in Section 4. These categories also apply to the existing bounds on $\lambda_{2}$, so we discuss the history of these bounds and techniques in Section 2.
In Section 5, we will give a refined Cheeger-style inequality for vertex expansion that can be specialized for a number of graph classes. This refinement will involve the “padded partition modulus”, which we will define in the next section. The most interesting specialized (constructive) implication of the yet-unstated refined Cheeger inequality is:
Corollary 1.7 (Refined Cheeger Inequality for Vertex Expansion in Minor-free Graphs).
For a graph $G$ with no $K_{h}$-minor,
$\frac{\psi(G)^{2}}{\min\{\log\Delta,(\log h)^{2}\}}\lesssim\lambda_{2}^{*}(G)\lesssim\psi(G).$
For the class of $K_{h}$-minor-free graphs, where $h$ is fixed (which includes a superclass of planar graphs), the dependence on the maximum degree $\Delta$ is completely eliminated. Hence, when combined with the upper bound on $\lambda_{2}^{*}$ for $K_{h}$-minor-free graphs, it is possible to use reweighted spectral partitioning (in polynomial time) to obtain separators with only a slightly sub-optimal dependence on $h$. Analogous forms of this statement for other graph classes are used to obtain the “poly-time separator size” column in Table 5.
In Section 6, we will give a high-level view displaying some surprising symmetries between quantities related to $\lambda_{2}^{*}$, and quantities used to obtain the “metric” bounds in Section 4, which suggest promising avenues for future work.
1.3 Parallel Work
In parallel to this work, Kam Chuen Tung independently derived the same bound on $\lambda_{2}^{*}$ and $\gamma^{(1)}$ for planar graphs stated in Theorem 3.5, as well as the bound on $\lambda_{2}^{*}$ for bounded-genus graphs stated in Corollary 4.9 (which forms one part of the bound given in Table 3), and a slightly weaker bound on $\lambda_{2}^{*}$ for $K_{h}$-minor-free graphs compared to the one stated in Corollary 4.9. The gap for forbidden-minor graphs is due to improvements external to both works. The technique for planar graphs is identical, while the techniques for the other two bounds differ very slightly (specifically, we bound $\gamma^{(1)}$ first, rather than bounding $\lambda_{2}^{*}$ directly). These results have appeared in his recent PhD thesis [Tun25, Chapter 7], which also includes some interesting generalizations. Both projects originated from simultaneous independent course projects in 2022.
2 Detailed Background
In this section, we give some more detailed background that is important to our methods. We will first discuss the history of bounds on $\lambda_{2}(G)$ for a few graph classes in more detail. Then we will discuss two groups of tools used for some of those bounds.
2.1 Bounds on $\lambda_{2}(G)$ for Special Graph Classes
Bounds on $\lambda_{2}(G)$ are known for a wide variety of graph classes. One of the first, and essentially the only to date that is known to be tight (up to constant factors), is the one for planar graphs, discovered by Spielman and Teng [ST96, ST07]:
Theorem 2.1 ([ST96, ST07, Theorem 3.3]).
For a planar graph $G$ of $n$ vertices and maximum degree $\Delta$, $\lambda_{2}(G)\leq\frac{8\Delta}{n}$.
Combined with Theorem 1.1, this obtains the weaker form of the planar separator theorem. Spielman and Teng also conjectured several bounds for $\lambda_{2}$ in several other classes of graphs, including bounded-genus graphs and minor-free graphs. Of particular note (since we will propose a generalized form) is their conjecture for bounded-genus graphs:
Conjecture 2.2 ([ST96, ST07, Conjecture 1]).
For a graph $G$ of $n$ vertices, genus $g$ and maximum degree $\Delta$, the second smallest eigenvalue of the Laplacian matrix of $G$ has $\lambda_{2}(G)\lesssim\frac{\Delta g}{n}$.
We use the notation $a\lesssim b$ throughout the paper to denote that $a\leq\operatorname{\mathcal{O}}(b)$.
Progress towards this conjecture has been the subject of several follow-up papers [Kel04, Kel06b, BLR10, KLPT09]. The conjecture was first partly answered by Kelner [Kel04, Kel06b], who showed the following:
Theorem 2.3 ([Kel04, Kel06b, Theorem 2.3], expanded proof in [Kel06a, Chapters 11 and 12]).
For a triangulated $g$ graph $G$ with $n$ vertices and maximum degree $\Delta$, the second smallest eigenvalue of the Laplacian matrix of $G$ has $\lambda_{2}(G)\lesssim\frac{\text{poly}(\Delta)\cdot g}{n}$, where $\text{poly}(\Delta)$ denotes a polynomial in $\Delta$.
This result is of particular interest since the proof is entirely geometric, and it was essentially the last result of this kind to have a geometric proof. As mentioned earlier, the references for this result claim that it applies to general $g$ graphs. This is due to the fact that any $g$ graph can be triangulated by adding edges (i.e., without adding new vertices), and adding edges only increases $\lambda_{2}$. The issue with this logic is that adding edges naively could also very significantly increase $\Delta$, even to the order of $n$. A method of triangulation by adding edges that asymptotically preserves $\Delta$ is known for planar graphs [KB92, Kan93, KB97], but, to the best of our knowledge, no such result is known for $g$ graphs. We will later bypass this issue with a careful triangulation that also adds vertices.
Orthogonal progress on the conjecture was separately made by Biswal, Lee and Rao, who gave a bound polynomial in $g$ instead of $\Delta$, by making use of metric embeddings [BLR10]:
Theorem 2.4 ([BLR10, Theorem 5.2]).
For a graph $G$ of $n$ vertices with genus $g$, maximum degree $\Delta$, and at least $n\gtrsim g$ vertices, the second smallest eigenvalue of the Laplacian matrix of $G$ has $\lambda_{2}(G)\lesssim\frac{\Delta\cdot g^{3}}{n}$.
Note that the restriction of $n\gtrsim g$ is actually unnecessary: It follows from Geršgorin circle theorem [Ger31] that $\lambda_{2}(G)\leq 2\Delta$ for all graphs. So, if $n\lesssim g$, then $\lambda_{2}(G)\lesssim\Delta\lesssim\frac{\Delta g}{n}\lesssim\frac{\Delta\cdot g^{3}}{n}$.
Their methods also allowed them to give a bound for $K_{h}$-minor-free graphs:
Theorem 2.5 ([BLR10, Theorem 5.3]).
For a graph $G$ of $n$ vertices which contains no $K_{h}$ minor, has maximum degree $\Delta$, the second smallest eigenvalue of the Laplacian matrix of $G$ has $\lambda_{2}(G)\lesssim\frac{\Delta h^{7}}{n}$. If $G$ has at least $n\gtrsim h\sqrt{\log h}$ vertices, then $\lambda_{2}(G)\lesssim\frac{\Delta\cdot h^{6}\log h}{n}$.
Note that the second of these bounds is stronger (and more generally applies without the restriction $n\gtrsim h\sqrt{\log h}$) by the same argument as above.
The bounds of Biswal, Lee, and Rao have since been slightly tightened due to improvements of some mathematical tools used to devise them, resulting in the improved bounds given in Table 2. In particular, one of the key steps in their method is to apply a certain form of metric embedding. This metric embedding induces a $g^{2}$ factor in the bounded-genus bound, and an (asymptotic) $h^{4}$ factor in both forbidden minor bounds. These metric embedding results have since been improved to $(\log g)^{2}$ [LS10] and $(\log h)^{2}$ [CF25], respectively. Note that the latter bound is more general. Moreover, the existing $\sqrt{\log h}$ factor in the vertex count requirement in the second forbidden minor bound (which appears squared in the corresponding bound on $\lambda_{2}$) arises from a weak form of Hadwiger’s conjecture, and has since been improved to $\log\log h$ [DP25]. We discuss these improvements in a bit more detail later in this section.
A similar theorem is also known for region intersection graphs:
Theorem 2.6 ([Lee17]).
Let $G$ be a $K_{h}$-minor-free graph. For a graph $\hat{G}\in\text{rig}(G)$ with maximum degree $\Delta$ and $n$ vertices,
$\lambda_{2}(G)\lesssim\frac{\Delta^{2}\cdot h^{6}\log h}{n}.$
Compared to the bounds for minor-free and bounded-genus graphs, this bound for region intersection graphs is weaker, and does not generalize the bound for minor-free graphs (even for constant $h$) as we might expect. To do so for constant $h$ would require replacing one of the $\Delta$ (the maximum degree) factors by $m/n$ (the average degree). Unfortunately, it is not clear if this could be done.
2.2 Extremal Spread
To obtain their bounds on $\lambda_{2}$, Biswal, Lee, and Rao [BLR10] studied the “$L_{2}$-extremal spread”. In slightly more generality:
Definition 2.7.
For a graph $G$ and a non-negative vertex weighting function $\omega:V(G)\to\mathbb{R}_{\geq 0}$, let $d_{G}^{\omega}:V\times V\to\mathbb{R}_{\geq 0}$ be the vertex-weighted shortest-path semi-metric through $G$ with vertex-weights given by $\omega$ (the vertex-weighted length of a path is the sum of the weights of the vertices along it, halving the contributions of the first and last vertex). The spread of $\omega$ is
$\sum_{u,v\in V(G)}d_{G}^{\omega}(u,v).$
The $L^{p}$-extremal spread of $G$ is
$\overline{s}_{p}(G):=\sup_{\omega:V(G)\to\mathbb{R}_{\geq 0},||\omega||_{p}\leq 1}\sum_{u,v\in V(G)}d_{G}^{\omega}(u,v).$
{TAB}
(2,1pt,1pt)[4pt]—c—c—c——c—c—c—c—c— Graph class $\overline{s}_{2}(G)\gtrsim$ Refs.
[…]
$K_{h}$-minor-free $n\geq 4h$ $\dfrac{n^{2}}{h^{3/2}}$ [BLR10, Theorems 3.9 and 3.11]
$K_{h}$-minor-free $n\gtrsim h\sqrt{\log h}$ $\dfrac{n^{2}}{h\sqrt{\log h}}$ [BLR10, Theorems 3.9 and 3.11]
RIG over $K_{h}$-minor-free $\dfrac{n^{2}}{h\sqrt{\Delta\log h}}$ [Lee17, Corollary 5.4]
Table 6: Asymptotic lower bounds on the $L^{2}$-extremal spread $\overline{s}_{2}(G)$ for various graph classes. $n$ is always the number of vertices, and $\Delta$ is the maximum degree.
Note that, by definition, $\overline{s}_{2}(G)\geq\overline{s}_{1}(G)$. A number of bounds for $\overline{s}_{2}$ are known, which we will review. A summary can be found in Table 6.
By studying the dual of the $L_{2}$-extremal spread, and applying crossing number bounds, Biswal, Lee, and Rao were able to prove the following results for special graph classes:
Theorem 2.8 ([BLR10, Theorem 3.1]).
For a $g$ graph $G$ with $n$ vertices, if $n\geq 3\sqrt{g}$, then
$\overline{s}_{2}(G)\gtrsim\frac{n^{2}}{\sqrt{g}}.$
Theorem 2.9 ([BLR10, Theorems 3.9 and 3.11]).
For a $K_{h}$-minor-free graph $G$ on $n$ vertices, where $n\geq 4h$,
$\overline{s}_{2}(G)\gtrsim\frac{n^{2}}{h^{\frac{3}{2}}}.$
Furthermore, if $n\gtrsim h\sqrt{\log h}$, then
$\overline{s}_{2}(G)\gtrsim\frac{n^{2}}{h\sqrt{\log h}}.$
The bound for bounded-genus graphs is essentially optimal, while the second bound for $K_{h}$-minor-free graphs is essentially optimal up to the $\sqrt{\log h}$ factor. Two recent results actually improve the $\sqrt{\log h}$ in the second bound: Norin, Postle, and Song [NPS23] first showed that it may be replaced with $(\log h)^{\frac{1}{4}+\varepsilon}$, for any $\varepsilon>0$. Delcourt and Postle [DP25] further improved this to $\log\log h$. The possibility of reducing this factor to $\operatorname{\mathcal{O}}(1)$ is closely related to Hadwiger’s conjecture, which is a major open problem in graph theory [Die25].
For region intersection graphs, Lee [Lee17] instead studied the dual of the $L^{1}$-extremal spread, generalizing results of Matoušek for string graphs [Mat14]. In particular, Lee obtained the following result based on an argument of Matoušek’s:
Theorem 2.10 ([Lee17, Theorem 2.2]).
For graphs $G$ and $\hat{G}$, where $\hat{G}\in\text{rig}(G)$: If $G$ has no $K_{h}$-minor, then $\overline{s}_{1}(\hat{G})\gtrsim\frac{n^{2}}{h\sqrt{m\log h}}$.
Note that our definition of $L^{1}$-extremal spread differs from Lee’s by a factor of $n$. Using essentially the same argument, Lee also obtained a relatively weak bound on $\overline{s}_{2}$:
Theorem 2.11 ([Lee17, Corollary 5.4]).
For graphs $G$ and $\hat{G}$, where $\hat{G}\in\text{rig}(G)$, so that $\Delta$ is the maximum degree of $\hat{G}$: If $G$ has no $K_{h}$-minor, then $\overline{s}_{2}(\hat{G})\gtrsim\frac{n^{2}}{h\sqrt{\Delta\log h}}$.
Similarly to Theorem 2.9, the $\sqrt{\log h}$ factor in these two theorems can also be replaced by a $\log\log h$ factor.
2.3 Non-Expansive Metric Embeddings
In proving their respective bounds on $\lambda_{2}$, Biswal, Lee, and Rao [BLR10], and Lee and Sidiropoulos [LS10] both made use of a certain forms of metric embeddings as a key step. We discuss such embeddings here, along with some other relevant literature. Importantly, this is the step that produces most of the loss for the bounds on $\lambda_{2}$.
Definition 2.12.
For two metric spaces $(X,d)$ and $(Y,d^{\prime})$ with a metric embedding function $f:X\to Y$, we say that $f$ is non-expansive if $d^{\prime}(f(x),f(y))\leq d(x,y)$ for all $x,y\in X$.
Definition 2.13.
For two metric spaces $(X,d)$ and $(Y,d^{\prime})$ with a metric embedding function $f:X\to Y$, we say the average $p$-distortion of $f$ is
$\frac{\sum_{x,x^{\prime}\in X}d(x,x^{\prime})^{p}}{\sum_{x,x^{\prime}\in X}d^{\prime}(f(x),f(x^{\prime}))^{p}}.$
It’s worth noting here that $\sum_{x,x^{\prime}\in X}\frac{d(x,x^{\prime})^{p}}{d^{\prime}(f(x),f(x^{\prime}))^{p}}\leq\frac{\sum_{x,x^{\prime}\in X}d(x,x^{\prime})^{p}}{\sum_{x,x^{\prime}\in X}d^{\prime}(f(x),f(x^{\prime}))^{p}}$, so an upper bound on the average $p$-distortion also provides an upper bound on this quantity.
This notion of average distortion was studied by Rabinovich [Rab08]. This is a bit different than most typical contexts for metric embeddings, which usually consider “worst-case distortion”. In fact, we can often obtain better bounds for this kind of “average” distortion.
The primary method for obtaining bounds on these average distortion values is by studying yet another quantity:
Definition 2.14.
[…]
$\begin{array}[]{c c c c}\sup_{x\in X}&\Pr_{P\sim\mu}\left[B\left(x,\frac{\Delta}{\alpha}\right)\subset P(x)\right]&\geq&\frac{1}{2},\end{array}$
where $B(x,r)$ denotes the metric ball of radius $r$ centred at $x$.
There are some other definitions of this quantity as well, but this form is suitable for our purposes. In particular, we will not need to use the structure of the modulus of padded decomposability directly – all the relevant results to this work have already been derived.
Theorem 2.15 ([BLR10, Theorem 4.4]).
[…]
We are particularly interested in the modulus of padded decomposability for shortest-path metrics on graphs in special classes. We use the following notation:
Definition 2.16.
For a graph $G=(V,E)$, the extremal edge and vertex moduli of padded decomposability on $G$ are defined as the maximum values of $\alpha(V,d_{\omega})$ across all edge-weight functions $\omega:E\to\mathbb{R}_{\geq 0}$ and vertex-weight functions $\omega:V\to\mathbb{R}_{\geq 0}$ (respectively) inducing shortest-path metrics $d_{\omega}$ on $G$. We denote the edge-weighted quantity as $\alpha(G)$, and the vertex-weighted quantity as $\overline{\alpha}(G)$. Note that $\overline{\alpha}(G)\leq\alpha(G)$.
There is a long sequence of results bounding $\alpha(G)$ for various graph classes. A very recent result has subsumed most of them:
Theorem 2.17 ([CF25, Theorem 2]).
Let $G$ be a $K_{h}$-minor-free graph. Then $\alpha(G)\in\operatorname{\mathcal{O}}(\log h)$. Consequently, for any set of edge or vertex weights $\omega$ and an induced shortest-path distance metric $d_{\omega}$, there exists a non-expansive metric embedding function $f:(V(G),d_{\omega})\to\mathbb{R}$ with average 222-distortion at most $\operatorname{\mathcal{O}}((\log h)^{2})$.
Conroy and Filtser outline the recent history of the results subsumed by the above [CF25]. Importantly, the above result also implies that $\alpha(G)=\operatorname{\mathcal{O}}(\log g)$ for a graph of genus $g$.
Lee [Lee17] proved a similar theorem for region intersection graphs that is not immediately subsumed by the above:
Theorem 2.18 ([Lee17]).
For graphs $G$ and $\hat{G}$, where $\hat{G}\in\text{rig}(G)$: If $G$ has no $K_{h}$-minor, then $\overline{\alpha}(G)\in\operatorname{\mathcal{O}}(h^{2})$.
It seems likely that the $\operatorname{\mathcal{O}}(h^{2})$ here could also be reduced to $\operatorname{\mathcal{O}}(\log h)$ as well, but the techniques of both papers are quite involved, so it would take some careful examination. It is also not immediately clear if the $\overline{\alpha}(G)$ could be replaced by $\alpha(G)$.
3 Bounds via Geometric Intersection Graphs
In this section, we will show that the dimension-restricted dual program $\gamma^{(d)}$ to $\lambda_{2}^{*}$ is intimately related to a certain kind of $d$-dimensional geometric intersection graphs. Using these geometric techniques, we will obtain bounds on both quantities for planar graphs, $k$-nearest neighbour graphs, and bounded-genus graphs. As a consequence, we will recover a new proof of the planar separator theorem. The arguments in this section adapt methods by Spielman and Teng [ST96, ST07], as well as Kelner [Kel04, Kel06b, Kel06a].
We start by defining a modified form of $\gamma^{(d)}$:
$\begin{array}[]{r c c r c l l}\dot{\gamma}^{(d)}(G)&:=&\min_{\begin{subarray}{c}f:V\to\mathbb{R}^{d}\\ s:V\to\mathbb{R}_{\geq 0}\end{subarray}}&\frac{\sum_{v\in V}\left(s(v)\right)^{2}}{\sum_{x\in V}||f(x)||_{2}^{2}}\\ &&\textup{subject to}&\sum_{v\in V}f(v)&=&\hskip 10.00002pt\overline{0}\\ &&&s(u)+s(v)&\geq&\hskip 10.00002pt||f(u)-f(v)||_{2}&\forall uv\in E\end{array}$ (2)
Lemma 3.1.
For any graph $G$, and any value $d\geq 1$, $\frac{1}{2}\gamma^{(d)}\leq\dot{\gamma}^{(d)}(G)\leq\gamma^{(d)}(G)$.
Proof.
For the first inequality, let $s,f$ be the optimal solution to $\dot{\gamma}^{(d)}(G)$, and take $y(v):=\frac{2s(v)^{2}}{\sum_{x\in V}||f(x)||_{2}^{2}}$ for each $v\in V$. Then, for each edge $uv\in E$,
$y(u)+y(v)=\frac{2s(u)^{2}+2s(v)^{2}}{\sum_{x\in V}||f(x)||_{2}^{2}}\geq\frac{\left(s(u)+s(v)\right)^{2}}{\sum_{x\in V}||f(x)||_{2}^{2}}\geq\frac{||f(u)-f(v)||_{2}^{2}}{\sum_{x\in V}||f(x)||_{2}^{2}},$
so the constraints are satisfied. Moreover $\sum_{v\in V}y(v)=2\cdot\dot{\gamma}^{(d)}(G)$, so we get the inequality $\gamma^{(d)}(G)\leq\sum_{v\in V}y(v)=2\cdot\dot{\gamma}^{(d)}(G)$.
For the second inequality, let $y,f$ be the optimal solution to $\gamma^{(d)}(G)$, and take
$s(v):=\sqrt{y(v)\cdot\sum_{x\in V}||f(x)||_{2}^{2}}$
for each $v\in V$. We perform a similar argument for each $uv\in E$:
$\left[s(u)+s(v)\right]^{2}\geq(s(u))^{2}+(s(v))^{2}=\left[y(u)+y(v)\right]\cdot\sum_{x\in V}||f(x)||_{2}^{2}\geq||f(u)-f(v)||_{2}^{2},$
so the constraints are satisfied after taking the square root of each side. Moreover, $\sum_{v\in V}s(v)^{2}=\gamma^{(d)}(G)$, so we get the inequality $\dot{\gamma}^{(d)}(G)\leq\sum_{v\in V}s(v)^{2}=\gamma^{(d)}(G)$. ∎
Importantly, $\dot{\gamma}^{(d)}(G)$ gives us a new geometric interpretation of the problem: A solution satisfying the constraints corresponds to a representation of $G$ as a subgraph of a geometric intersection graph: For each vertex $v$, create a $d$-dimensional ball $B_{v}$ of radius $s(v)$ centered at $f(v)$. Construct a new graph $H=(V,F)$ so that $uv\in F$ if and only if $B_{u}\cap B_{v}\neq\emptyset$. Then $H\supset G$. As a consequence, if we can bound the sum of squared radii for such a geometric representation, with the normalization constraints $\sum_{x\in V}||f(x)||_{2}^{2}=1$ and $\sum_{x\in V}f(x)=\overline{0}$, then we can bound $\gamma^{(d)}(G)$.
One particularly important natural case of a graph class that admits such representations are planar graphs, in which such a representation corresponds to special form of the well-celebrated circle packing theorem. We start with a definition:
Definition 3.2.
A planar circle packing $\mathcal{P}$ of a simple undirected graph $G$ is a set of (possibly overlapping) circles $\{C\}_{v\in V(G)}$ in the complex plane $\mathbb{C}$ such that $uv\in E(G)$ if and only if the circles $C_{u}$ and $C_{v}$ are mutually tangent (with disjoint interiors). If all pairs of circles also have pairwise disjoint interiors, then we say that $\mathcal{P}$ is univalent.
Figure 1: Two examples planar circle packings. The left planar circle packing is univalent, while the right is not.
See Figure 1 for examples of univalent and non-univalent circle packings. We can now state the theorem:
Theorem 3.3 (Planar Circle Packing Theorem/Koebe-Andreev-Thurston Theorem [Koe36, And70, TM79]).
Let $G$ be a simple undirected graph. Then $G$ has a univalent planar circle packing if and only if $G$ is a planar graph. Moreover, if $G$ is a maximal planar graph, then this circle packing is unique up to Möbius transformations.
Spielman and Teng [ST96, ST07] made use of this theorem to bound $\lambda_{2}(G)$ for planar graphs $G$, hence giving a proof of a weaker version of the planar separator theorem via Theorem 1.1. We will use a similar argument to bound $\lambda_{2}^{*}(G)$, allowing us to give a new proof of the full planar separator theorem via Theorem 1.4. We will make use of one of their theorems to accomplish this:
Theorem 3.4 ([ST96, ST07, Theorem 4.5]).
[…]
This theorem statement is slightly weaker than the statement used by Spielman and Teng, who also described the structure of the homeomorphism as a stereographic projection. However, it is sufficient for our purposes. In particular, it gives us a method of “normalizing” a geometric intersection graph of balls in $\mathbb{R}^{d}$ to be a geometric intersection graph of (geodesic) balls in $S^{d}$. Importantly, since the theorem gives a homeomorphism, it preserves the ply of each individual point, which is the total number of balls containing that point. From this, we can now obtain the desired bound on $\lambda_{2}^{*}$ for planar graphs:
Theorem 3.5.
Let $G$ be a planar graph with $n$ vertices and maximum degree $\Delta$. Then $\frac{\lambda_{2}(G)}{\Delta}\leq\lambda_{2}^{*}(G)\leq\gamma^{(3)}(G)\leq\frac{8}{n}$. Hence, $\psi(G)^{2}\lesssim\frac{1}{n}$.
Note that this theorem is a strengthening of both Theorem 2.1 and the planar separator theorem. The proof method will also be analogous to that of Spielman and Teng [ST96, ST07, proof of Theorem 3.3].
Proof.
If we can show that $\gamma^{(3)}(G)\leq\frac{8}{n}$, the other inequalities in the statement follow from previous theorem statements.
By Theorem 3.3 and Theorem 3.4, there exists a representation of $G$ as a circle packing on the sphere $S^{2}\subset\mathbb{R}^{3}$. Let the centers be given by $f:V\to S^{2}\subset\mathbb{R}^{3}$ (i.e., $||f(v)||_{2}^{2}=1$ for each $v\in V$) and let the radii be given by $s:V\to\mathbb{R}_{\geq 0}$. We claim first that $f,s$ form a feasible solution to $\dot{\gamma}^{(3)}(G)$: The constraint $\sum_{v\in V}f(v)=\overline{0}$ is exactly the statement that the centroid of the centres is the origin, which is given by Theorem 3.4. The other constraints follow from the facts that two balls intersect if and only if they share an edge, and that $||f(u)-f(v)||_{2}$ is bounded above by the geodesic distance along $S^{2}$ between $f(u)$ and $f(v)$. Next, we bound the objective value: It follows from the statement of Theorem 3.3 that the ply of all points in $S^{2}$ except for a set of measure 0 is at most 111. Hence, the total area of all balls is bounded by the area of $S^{2}$ itself, which is known to be $4\pi$. The area of the ball on $S^{2}$ for vertex $v$ is bounded above by $\pi(s(v))^{2}$, since $s(v)$ is the (geodesic) radius of the ball, so $\sum_{v\in V}\pi(s(v))^{2}\leq 4\pi$. Therefore, we get a bound on the objective value of
$\frac{\sum_{v\in V}s(v)^{2}}{\sum_{x\in V}||f(x)||_{2}^{2}}\leq\frac{4}{n}.$
Finally, the result follows from Lemma 3.1. ∎
3.1 Geometric Intersection Graphs and Nearest-Neighbour Graphs
Spielman and Teng [ST96, ST07] also used higher-dimensional geometric intersection graphs with similar properties to bound $\lambda_{2}$ for a few other graph classes. In particular, they obtained bounds for $d$-dimensional $k$-ply neighbourhood systems, and $d$-dimensional $k$-nearest-neighbour graphs.
We can strengthen another key result of theirs:
Theorem 3.6.
[…]
$\frac{\lambda_{2}(G)}{\Delta}\leq\lambda_{2}^{*}(G)\leq\gamma^{(d+1)}(G)\leq 2\cdot\left(\frac{A_{d}}{V_{d}}\cdot\frac{k}{n}\right)^{\frac{2}{d}}<2\cdot\left(2\pi\cdot\left(d+2\right)\right)^{\frac{1}{d}}\cdot\left(\frac{k}{n}\right)^{\frac{2}{d}}\lesssim\left(\frac{k}{n}\right)^{\frac{2}{d}}.$
Consequently, $\psi(G)\lesssim\sqrt{\min\{d,\log\Delta\}}\cdot\left(\frac{k}{n}\right)^{\frac{1}{d}}$. Moreover, reweighted spectral partitioning will find, in polynomial time, a sparse vertex cut $S$ with $\frac{|\partial S|}{|S|}\lesssim\sqrt{\log\Delta}\cdot\left(\frac{k}{n}\right)^{\frac{1}{d}}$.
Note that this theorem is essentially a further generalization of Theorem 3.5.
Proof.
[…]
$\sum_{i=1}^{n}V_{d}\cdot r_{i}^{\prime\prime d}\leq\sum_{i=1}^{n}\text{volume}(B_{i}^{\prime})\leq k\cdot A_{d}.$
[…]
$\dot{\gamma}^{(d)}(G)\leq\frac{\sum_{v\in V}r_{i}^{\prime\prime 2}}{n}\leq\left(\frac{\sum_{v\in V}r_{i}^{\prime\prime d}}{n}\right)^{\frac{2}{d}}\leq\left(\frac{A_{d}\cdot k}{V_{d}\cdot n}\right)^{\frac{2}{d}},$
where the second inequality follows from the power-mean inequality. ∎
Figure 2: An example of a geodesic ball $B_{i}^{\prime}$ on $S^{1}$ embedded in $\mathbb{R}^{2}$, with $c_{i}^{\prime},r_{i}^{\prime},r_{i}^{\prime\prime}$ all labelled.
As previously mentioned, we will also apply this result to $k$-nearest neighbour graphs, again in a similar manner to Spielman and Teng [ST96, ST07, Corollary 5.2]. Let $\tau_{d}$ denote the kissing number in $d$ dimensions, which is the maximum number of non-overlapping unit balls in $\mathbb{R}^{d}$ arranged to all touch a central unit ball. It is known that as $d\to\infty$, $2^{0.2075d(1+\operatorname{o}(1))}\leq\tau_{d}\leq 2^{0.401d(1+\operatorname{o}(1))}$ [Wyn65, KL78]. Hence, $\log\tau_{d}\gtrsim d$ and $\tau_{d}^{\frac{1}{d}}\lesssim 1$. Miller, Teng, Thurston, and Vavasis [MTTV97] observed that every $k$-nearest neighbour graph is the subgraph of a $\tau_{d}k$-ply neighbourhood system, and moreover that every $k$-nearest neighbour graph has maximum degree bounded by $\tau_{d}k$. Hence, we obtain the following corollary:
Corollary 3.7.
Let $G$ be a $d$-dimensional $k$-nearest neighbour graph with $n$ vertices. Then,
$\frac{\lambda_{2}(G)}{\Delta}\leq\lambda_{2}^{*}(G)\leq\gamma^{(d+1)}(G)<2\cdot\left(2\pi\cdot\left(d+2\right)\right)^{\frac{1}{d}}\cdot\left(\frac{\tau_{d}k}{n}\right)^{\frac{2}{d}}\lesssim\left(\frac{k}{n}\right)^{\frac{2}{d}}.$
Consequently, $\psi(G)\lesssim\sqrt{\min\{d,\log\Delta\}}\cdot\left(\frac{k}{n}\right)^{\frac{1}{d}}$. Moreover, reweighted spectral partitioning will find, in polynomial time, a sparse vertex cut $S$ with $\frac{|\partial S|}{|S|}\lesssim\sqrt{\log\Delta}\cdot\left(\frac{k}{n}\right)^{\frac{1}{d}}\lesssim\sqrt{d+\log k}\cdot\left(\frac{k}{n}\right)^{\frac{1}{d}}$.
Note that the bound on $\frac{|\partial S|}{|S|}$ is simplified due to the bound $\log\Delta\lesssim\log\tau_{d}k\lesssim d+\log k$.
The above two statements apply algorithmically to reweighted spectral partitioning. We note that, in some applications, the geometric information may actually be known, in which case the constructions above can be used more directly:
Corollary 3.8.
Let $G$ be a $d$-dimensional $k$-ply neighbourhood system with maximum degree $\Delta$, provided as a set of $d$-dimensional coordinates and radii for each vertex. Then, in polynomial time, we can compute a balanced 2/32/32 / 3-vertex-separator of size $\operatorname{\mathcal{O}}\left(\sqrt{\min\{d,\log\Delta\}}\cdot\left(\frac{k}{n}\right)^{\frac{1}{d}}\right)$.
Corollary 3.9.
Let $G$ be a $d$-dimensional $k$-nearest neighbour graph with maximum degree $\Delta$, provided as a set of $d$-dimensional coordinates and a value $k$. Then, in polynomial time, we can compute a balanced 2/32/32 / 3-vertex-separator of size $\operatorname{\mathcal{O}}\left(\sqrt{\min\{d,\log\Delta\}}\cdot\left(\frac{k}{n}\right)^{\frac{1}{d}}\right)$.
3.2 Bounded-Genus Graphs
Kelner’s bound on $\lambda_{2}$ (stated in Theorem 2.3) for triangulated $g$ graphs also makes use of circle packings [Kel04, Kel06b, Kel06a], and, as we will see, can also be generalized to a bound on $\lambda_{2}^{*}$ and $\gamma^{(1)}$. However, the resulting bound also has a (large) polynomial dependence on the maximum degree $\Delta$. Rather than directly applying this bound, we will apply a new transformation that allows us to almost fully tighten this dependence. Specifically, our new transformation will take a $g$ graph with $n$ vertices and maximum degree $\Delta$, and produce a graph with $O(\Delta\cdot n)$ vertices and constant maximum degree. We will then relate the values of $\gamma^{(1)}$ for the two graphs. In the case of $\lambda_{2}^{*}$ and $\gamma^{(1)}$, this relationship will result in a greatly improved $O(\log\Delta)$ factor for the upper bound.
Recall that a graph $G$ has genus $g$ if it has a combinatorial embedding (represented as a rotation system) onto a surface of genus $g$. That is, each vertex has a circular ordering on its vertices, forming a locally-planar embedding. Such an embedding is cellular if each face induced by the rotation system is simple. A cellular embedding is a triangulation if each face defined by the rotation system is a triangle.
It is notable that Kelner’s bound on $\lambda_{2}$ does not apply to non-triangulated $g$ graphs, despite the paper claiming otherwise. The claimed argument is as follows:
  • Adding edges to a graph can only increase $\lambda_{2}$.
  • Add edges until the graph is maximal.
  • Apply the bound on triangulated $g$ graphs that Kelner proves in the rest of his paper. Note that this bound relies on a generalized circle packing theorem of He and Schramm [HS93] that applies only to simple triangulations of $g$ surfaces.
This argument has two separate issues:
  • A maximal $g$ graph is not always a triangulation. In particular, $K_{5}$ is a maximal genus-111 graph that is not triangulated in any embedding into the torus. We may prove this by contradiction: Suppose this was not the case, and that it had a triangulated embedding in the torus. Then, by the Euler characteristic, such an embedding must have 555 faces. We may count the number of edge-face incidences. Since the embedding is triangulated (and hence cellular), each edge is incident to exactly two faces, and there are 101010 edges and hence 210=202$\cdot$ 10=202 ⋅ 10 = 20 incidences in total. Since the embedding is triangulated, there are exactly 333 edges incident to each face, and hence there are 35=153$\cdot$ 5=153 ⋅ 5 = 15 incidences in total – a contradiction.
  • Naively adding edges may increase the maximum degree, possibly up to $\Omega(n)$, and the bound that Kelner obtains for triangulations has a dependence on the maximum degree.
It is worth noting that neither of these issues arise for planar graphs: Maximal planar graphs are always planar triangulations, and there is a known result that triangulates a planar graph while asymptotically preserving the maximum degree [KB92, KB97]. For genus $g>0$, we will present a method to bypass these issues with carefully chosen Steiner vertices while simultaneously tightening and generalizing the bound to $\lambda_{2}^{*}$ and $\gamma^{(1)}$.
Henceforth, we will fix a rotation system for our initial graph $G$. This rotation system need not be cellular, but we will assume that $G$ is simple (as we do throughout this paper). Importantly for our purposes, this will allow us to make certain local transformations to a graph without changing its genus, so long as the rotation system is respected in these transformations. These transformations will make use of the following structure:
Definition 3.10.
Let $G=(V,E)$ be a graph with $n$ vertices, and let $H=(V^{\prime},E^{\prime})$ be a graph with $r\cdot n$ vertices. Suppose there is a function $p:V^{\prime}\to V$ so that $|p^{-1}(v)|=r$ for each $v\in V$, so that $p^{-1}(v)|$ is also a connected subset of vertices with diameter $L$. Suppose furthermore that for each edge $uv\in E^{\prime}$, $p(u)p(v)\in E$. Then we say that $G$ is a uniform shallow minor of $H$ with depth $L$.
Usually when studying minors, we usually care about “forbidden” minors of a particular given graph. However, we will use uniform shallow minors in the opposite sense: Given a graph $G$, we will usually aim to find a graph $H$ for which $G$ is a uniform shallow minor. In particular, their use is characterized by the following technical lemma:
Lemma 3.11.
Let $G=(V,E)$ be a graph with $n$ vertices, and let $H=(V^{\prime},E^{\prime})$ be a graph with $|V^{\prime}|=r\cdot n$ vertices. Suppose $G$ is a uniform shallow minor of $H$ with depth $L$. Then $\gamma^{(1)}(G)\lesssim r\cdot L\cdot\gamma^{(1)}(G)$.
This lemma will be useful in the following sense: If we have a graph $G$ of genus $g$, we may be able to find a graph $H$ of genus $g$ by performing local transformations at each vertex, so that $G$ is a uniform shallow minor of $H$. In particular, we may be able to find an $H$ with useful properties, such as a reduced maximum degree or faces that are easier to triangulate. Then, a bound on $\gamma^{(1)}(H)$ will imply a bound on $\gamma^{(1)}(G)$, with “loss” $L$.
Proof.
We start with a trivial case whose elimination we will use later: If $\gamma^{(1)}(H)\cdot r\cdot L\gtrsim 1$, then the result holds since $\gamma^{(1)}(G)\lesssim 1$ (see 1.3).
[…]
Let $\pi(v)$ be a random variable that is uniformly chosen over $p^{-1}(v)$, so that each $\pi(v)$ is independent. Let $m(v)$ be samples of all such $\pi(v)$ so that
$\sum_{u\in V}\sum_{v\in V}|f_{H}(m(u))-f_{H}(m(v))|^{2}\geq\operatorname{\mathbb{E}}\left[\sum_{u\in V}\sum_{v\in V}|f_{H}(\pi(u))-f_{H}(\pi(v))|^{2}\right].$
Let $f_{G}(v):=f_{H}(m(v))-\frac{1}{n}\sum_{x\in V}f_{H}(m(x))$, and let $y_{G}(v):=4r\cdot(2L+1)\cdot\sum_{v^{\prime}\in p^{-1}(v)}y_{H}(v)$.
Under these choices, the objective claim is satisfied since $\sum_{v\in V}y_{G}(v)=4r\cdot(2L+1)\cdot\sum_{v^{\prime}\in V^{\prime}}y_{H}(v)$, and hence we need only check that the constraints of $\gamma^{(1)}(G)$ are satisfied by $f_{G},y_{G}$.
The first constraint is satisfied since $\sum_{v\in V}f_{G}(v)=\sum_{v\in V}f_{H}(m(v))-\sum_{x\in V}f_{H}(m(x))=\overline{0}$. Next, note that for any vertex $v\in V$, and any two vertices $u^{\prime},v^{\prime}\in p^{-1}(v)$, there is a path $u^{\prime}=v_{1}^{\prime},\dots,v_{l+1}^{\prime}=v^{\prime}$ of length $l\leq L$ contained entirely within $p^{-1}(v)$. Hence, by the triangle inequality and Cauchy-Schwarz,
$\begin{array}[]{r c l}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}&\leq&\left[|f_{H}(v_{1}^{\prime})-f_{H}(v_{2}^{\prime})|+\cdots+|f_{H}(v_{l}^{\prime})-f_{H}(v_{l+1}^{\prime})|\right]^{2}\\[4.0pt] &\leq&l\left[|f_{H}(v_{1}^{\prime})-f_{H}(v_{2}^{\prime})|^{2}+\cdots+|f_{H}(v_{l}^{\prime})-f_{H}(v_{l+1}^{\prime})|^{2}\right]\\[4.0pt] &\leq&L\left[|f_{H}(v_{1}^{\prime})-f_{H}(v_{2}^{\prime})|^{2}+\cdots+|f_{H}(v_{l}^{\prime})-f_{H}(v_{l+1}^{\prime})|^{2}\right]\\[4.0pt] &\leq&L\left[y(v_{1}^{\prime})+2y(v_{2}^{\prime})+2y(v_{3}^{\prime})+\cdots+2y(v_{l}^{\prime})+y(v_{l+1}^{\prime})\right],\end{array}$
and so for a fixed $v\in V$,
$\begin{array}[]{r l}&\sum_{u^{\prime}\in p^{-1}(v)}\sum_{v^{\prime}\in p^{-1}(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}\\[4.0pt] \leq&\sum_{u^{\prime}\in p^{-1}(v)}\sum_{v^{\prime}\in p^{-1}(v)}L\left[y(v_{1}^{\prime})+2y(v_{2}^{\prime})+2y(v_{3}^{\prime})+\cdots+2y(v_{l}^{\prime})+y(v_{l+1}^{\prime})\right]\\[4.0pt] \leq&2L\sum_{u^{\prime}\in p^{-1}(v)}\sum_{v^{\prime}\in p^{-1}(v)}\sum_{w^{\prime}\in p^{-1}(v)}y(w^{\prime}).=2Lr^{2}\sum_{w^{\prime}\in p^{-1}(v)}y(w^{\prime})\\[4.0pt] \end{array}$
Using this bound, we can also get a lower bound on $\sum_{v\in V}|f_{G}(v)|^{2}$:
$\begin{array}[]{r c l}2|V|\cdot\sum_{v\in V}|f_{G}(v)|^{2}&=&\sum_{u\in V}\sum_{v\in V}|f_{G}(u)-f_{G}(v)|^{2}\\[12.0pt] &=&\sum_{u\in V}\sum_{v\in V}|f_{H}(m(u))-f_{H}(m(v))|^{2}\\[12.0pt] &\geq&\operatorname{\mathbb{E}}\left[\sum_{u\in V}\sum_{v\in V}|f_{H}(\pi(u))-f_{H}(\pi(v))|^{2}\right]\\[12.0pt] &=&\frac{1}{r^{2}}\sum_{u\in V}\sum_{v\in V\setminus\{u\}}\sum_{u^{\prime}\in p^{-1}(u)}\sum_{v^{\prime}\in p^{-1}(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}\\[12.0pt] &=&\frac{1}{r^{2}}\left[\sum_{u^{\prime}\in V^{\prime}}\sum_{v^{\prime}\in V^{\prime}}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}\right.\\ &&\left.-\sum_{v\in V}\sum_{u^{\prime}\in p^{-1}(v)}\sum_{v^{\prime}\in p^{-1}(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}\right]\\[16.0pt] &\geq&\frac{1}{r^{2}}\left[2|V|\cdot r\cdot\sum_{v^{\prime}\in V^{\prime}}|f_{H}(v^{\prime})|^{2}\right.\\ &&\left.-2Lr^{2}\cdot\left(\sum_{v^{\prime}\in V^{\prime}}y_{H}(v^{\prime})\right)\cdot\left(\sum_{v^{\prime}\in V^{\prime}}|f_{H}(v^{\prime})|^{2}\right)\right]\\[12.0pt] &=&\frac{2\sum_{v^{\prime}\in V^{\prime}}|f_{H}(v^{\prime})|^{2}}{r}\left[|V|-r\cdot(L/2)\cdot\gamma^{(1)}(H)\right]\\[12.0pt] &\geq&\frac{2\sum_{v^{\prime}\in V^{\prime}}|f_{H}(v^{\prime})|^{2}}{r}\left[|V|-1\right],\end{array}$
where the last step follows from our early assumption on the value of $\gamma^{(1)}(H)$. Thus, $\sum_{v\in V}|f_{G}(v)|^{2}\geq\frac{\sum_{v^{\prime}\in V^{\prime}}|f_{H}(v^{\prime})|^{2}}{2r}$.
It remains to show the constraint for each edge is satisfied. Let $e=uv\in E$. Similarly to our earlier discussion, there is a path of length $k\leq 2L+1$ in $H$ from $m(u)$ to $m(v)$ using only vertices in $p^{-1}(\{u,v\})$. Denote this path as $m(u)=v_{1},\dots,v_{k+1}=m(v)$. Then,
$\begin{array}[]{r c l}y_{G}(u)+y_{G}(v)&=&4r(2L+1)\sum_{u^{\prime}\in p^{-1}(u)}y_{H}(u^{\prime})+\sum_{v^{\prime}\in p^{-1}(v)}y_{H}(v^{\prime})\\[8.0pt] &\geq&4r(2L+1)\left[y_{H}(v_{1})+\cdots+y_{H}(v_{k+1})\right]\\[8.0pt] &\geq&2r(2L+1)\sum_{i=1}^{k}\left[y_{H}(v_{i})+y_{H}(v_{i+1})\right]\\[8.0pt] &\geq&2r(2L+1)\frac{\sum_{i=1}^{k}\left|f_{H}(v_{i})-f_{H}(v_{i+1})\right|^{2}}{\sum_{x\in V^{\prime}}|f_{H}(x)|^{2}}\\[8.0pt] &\geq&(2L+1)\frac{\sum_{i=1}^{k}\left|f_{H}(v_{i})-f_{H}(v_{i+1})\right|^{2}}{\sum_{x\in V}|f_{G}(x)|^{2}}\\[8.0pt] &\geq&(2L+1)\frac{\left[\sum_{i=1}^{k}\left|f_{H}(v_{i})-f_{H}(v_{i+1})\right|\right]^{2}}{k\sum_{x\in V}|f_{G}(x)|^{2}}\\[8.0pt] &\geq&(2L+1)\frac{\left|f_{H}(v_{1})-f_{H}(v_{k+1})\right|^{2}}{k\sum_{x\in V}|f_{G}(x)|^{2}}\\[8.0pt] &\geq&\frac{\left|f_{G}(u)-f_{G}(v))\right|^{2}}{\sum_{x\in V}|f_{G}(x)|^{2}},\end{array}$
where the antepenultimate inequality follows from Cauchy-Schwarz, and the penultimate inequality follows from the triangle inequality. ∎
Lemma 3.12.
Let $G$ be a graph with genus $g$, $n$ vertices, and maximum degree $\Delta$. Then there exists a graph $H$ with $n\Delta$ vertices and maximum degree 444 with genus $g$ so that $\gamma^{(1)}(G)\lesssim\gamma^{(1)}(H)\cdot\Delta\cdot\log\Delta$.
Proof.
[…]
We claim that $H$ is of genus exactly $g$: $G$ is a minor of $H$ (by contracting all edges in $E^{\prime\prime}$), so clearly the genus of $H$ is at least $g$. Furthermore, a rotation system of $G$ can be extended to a rotation system of $H$ since the trees forming $E^{\prime\prime}$ are planar respecting the ordering induced by the rotation system (see Figure 3), so the genus of $H$ is also at most $g$.
Figure 3: Two examples of the degree-reduction construction for a $g$ graph, respecting a rotation system, local to a vertex. The vertex is replaced with $\Delta$ vertices, at least one vertex for each incident edge, and these new vertices are connected with a near-perfect binary tree. In the left example, the displayed vertex is of maximum degree in its graph, while in the right example it is not (resulting in some degree-111 vertices).
It remains to show that $\gamma^{(1)}(G)\lesssim\gamma^{(1)}(H)\cdot\Delta\cdot\log\Delta$. In fact, $G$ is exactly a uniform shallow minor of $H$ with depth $2\log_{2}\Delta$, so we simply apply Lemma 3.11. ∎
We will be able to leverage this lemma in order to turn a bound on $\lambda_{2}^{*}$ (and $\gamma^{(1)}$) for bounded-genus graphs of constant maximum degree into a bound on $\lambda_{2}^{*}$ (and $\gamma^{(1)}$) for bounded-genus graphs of arbitrary maximum degree at only a $\log\Delta$ loss. In particular, we will obtain a bound on $\lambda_{2}^{*}$ (and $\gamma^{(1)}$) for triangulated bounded-genus graphs of constant degree by adapting Kelner’s techniques [Kel04, Kel06b, Kel06a] for bounding $\lambda_{2}$. Before exploring these techniques, we first show how to reduce to the case of triangulated bounded-genus graphs, which will use a similar technique:
Lemma 3.13.
Let $G$ be a graph of genus $g$ with $n$ vertices and maximum degree $\Delta$. Then there is a triangulated $g$ graph $H$ with $(\Delta+1)\cdot n$ vertices and maximum degree $O(\Delta)$, so that $\gamma^{(1)}(G)\lesssim(\Delta+1)\cdot\gamma^{(1)}(H)$.
Proof.
We will assume for simplicity that $G$ is connected. $O(1)$ edges can be added to each vertex to make this the case if it is not true.
Figure 4: A demonstration of the triangulation steps on a face. From left to right, top to bottom: $G$, $H_{1}$, $H_{2}$, $H_{3}$, $H$.
Let $H_{1}$ be the graph that adds $\Delta$ vertices around each vertex to $G$, and ensure that between any two edges in the rotation system around a vertex in $G$, at least one such edge is added. Observe that $H_{1}$ is a $g$ graph with $(\Delta+1)n$ vertices and maximum degree $2\Delta$. Observe also that $G$ is a uniform shallow minor of $H_{1}$ of depth 222, and hence Lemma 3.11 shows that $\gamma^{(1)}(G)\lesssim(\Delta+1)\gamma^{(1)}(H_{1})$.
Now, we will add edges to $H_{1}$. In particular, we will construct a sequence of graphs that add edges to $H_{1}$, but never add vertices, and hence $\gamma^{(1)}$ will only increase for this sequence of graphs. Specifically, within each face of $H_{1}$, we create a cycle of all the newly added vertices in order around the face. The result is a graph we call $H_{2}$ (with genus $g$, $(\Delta+1)n$ vertices, and maximum degree $\max\{3,2\Delta\}$) that has no adjacent pair of faces both with size >4>4> 4. Moreover, any vertex incident to a face of size >4>4> 4 (i.e., those in the created cycles) is incident exactly two other faces, and has degree exactly 333. For each face of exactly size 444 in $H_{2}$, triangulate it arbitrarily to get another new graph $H_{3}$ (with genus $g$, $(\Delta+1)n$ vertices, and maximum degree $\max\{5,4\Delta\}$). See Figure 4 for examples of each of these graphs around a particular face. It’s worth noting that for non-cellular embeddings of higher genus graphs, it is possible that a non-simple face may include the same edge twice in the same “direction” during a traversal around the face (as opposed to the example in the figure, where the doubly-included edge is included in opposite directions). Fortunately, this does not cause issues for the construction described here either. It is also worth noting that these steps are necessary even if all the faces are simple: Specifically, we would like to handle the remaining non-triangular faces independently, and the separation ensures that we are not able to later produce parallel edges (see [KB97, Figure 6.3(b)] for a demonstration of the issues of adjacent simple faces).
It remains to triangulate the remaining non-triangular faces of $H_{3}$, none of which are adjacent, and all of which are simple. This can be accomplished by adding at most 222 edges per vertex. In particular, Kant and Bodlaender [KB92, KB97] gave an algorithm called “zig-zagging” that adds edges to a single face using a sequence of ear-cuts, triangulating the face. This algorithm applies in the case of faces on any rotation system, not just a rotation system on a planar graph. Call the final graph $H$, which is a triangulated $g$ graph with maximum degree $\max\{7,4\Delta\}$. See the last diagram in Figure 4 for a demonstration of this algorithm.
Since $H$ has the same vertex set as $H_{1}$, and a strict superset of its edges, $\gamma^{(1)}(H)\geq\gamma^{(1)}(H_{1})$, so the lemma follows from the earlier argument. ∎
For the remaining lemmas, we need only consider triangulated $g$ graphs.
A $g$ graph naturally embeds into a $g$ surface. Moreover, a $g$ graph in fact has a circle packing on a $g$ surface. Both the surface and packing is unique up to a certain set of transformations [HS93]. However, to bound $\gamma^{(3)}$ in a similar manner to Theorem 3.5, we would need a circle packing on the unit sphere $S^{2}$, likely of ply $O(g)$. Unfortunately, as Kelner [Kel04, Kel06b, Kel06a] pointed out, such circle packings need-not exist. Instead, in order to bound $\lambda_{2}$, Kelner made use of two steps: First, a specific subdivision procedure is applied to the graph, and it is shown that a bound for the graph obtained by a sequence of subdivision procedures also induces a bound on the initial graph. Second, by applying this procedure a sufficient number of times, it is shown that a continuous analogue of the circle packing theory can be used to obtain a bound. We will apply analogues of each of these steps for $\gamma^{(3)}$ (and $\gamma^{(1)}$, which differs only by a constant factor of at most 333) instead of $\lambda_{2}$.
The subdivision procedure used by Kelner is the hexagonal subdivision of a triangulation $G$, which produces a graph $G^{\prime}$ that by bisecting every edge in $E(G)$, and then connecting all three bisection vertices around each resulting face to form a triangle with new edges. See Figure 5 for an example of this operation.
Refer to caption    Refer to caption    Refer to caption
Figure 5: The hexagonal subdivision procedure applied to small example triangulation subgraphs. The left and centre drawings show a single application of the procedure, while the right drawing shows many.
For a triangulation $G$, let $G^{(k)}$ denote the triangulation that results from $k$ successive applications of hexagonal subdivision. The first step is to show that we can relate $\gamma^{(1)}(G)$ and $\gamma^{(1)}(G^{(k)})$:
Lemma 3.14.
Let $G$ be a triangulation of genus $g$ with maximum degree $\Delta$. Then there is some (universal) constant $c$ so that $|V(G)|\cdot\gamma^{(1)}(G)\lesssim\Delta^{c}\cdot|V(G^{(k)})|\cdot\gamma^{(1)}(G^{(k)})$.
We defer the proof of Lemma 3.14 to Appendix C since it is quite similar to that of Lemma 3.11, and will primarily make use of an argument of Kelner  [Kel06a, Kel04, Kel06b] that randomly “embeds” $G$ into $G^{(k)}$. Compared to the proof of Lemma 3.11, this randomized argument includes random paths in addition to random vertices.
The second step is to show that a sufficient number of subdivisions allows us to obtain a bound. This step will make use of the following lemma essentially proved by Kelner:
Lemma 3.15 ([Kel06b, Lemma 5.3, Lemma 5.4, proof of Theorem 2.3]).
Let $G$ be a graph of genus $g$. For each $k$, let $F^{(k)}$ denote a $g$ surface on which $G^{(k)}$ admits a circle packing. For each vertex $v\in G^{(k)}$, let $C_{v}$ denote the disk in the packing corresponding to $v$, and let $r_{v}$ denote its radius. Let $A(C)$ and $D(C)$ denote the area (measure) and diameter (longest geodesic path between a point pair), respectively, of a compact and connected region $C$ either in $F^{(k)}$ or $S^{2}$. Then there is a sequence $f^{(k)}:F^{(k)}\to S^{2}$ of analytic maps so that, for any $\varepsilon>0$, there exists a threshold $N$ such that for any $k\geq N$ a partition of $S^{(k)}$ into $S^{(k)}_{1}\cup S^{(k)}_{2}$ exists with the following properties:
  • For any vertex $v\in G^{(k)}$ where $C_{v}\subset S^{(k)}_{1}$,
    $r_{v}^{2}\lesssim D\left(f^{(k)}(C_{v})\right)^{2}\lesssim A\left(f^{(k)}(C_{v})\right).$
  • The projection of $S^{(k)}_{1}$ with $f^{(k)}$ has ply $O(g)$. That is: For each point $p\in S^{2}$,
    $\left|\left(f^{(k)}\right)^{-1}(p)\right|\lesssim g,$
    almost surely.
  • $\sum_{C_{v}:C_{v}\cap S^{(k)}_{2}\neq\emptyset}D\left(f^{(k)}(C_{v})\right)^{2}\lesssim\varepsilon.$
  • The mapping of the centres of each $C_{v}$ ($v\in V(G^{(k)})$) under $f^{(k)}$, for $S^{2}$ embedded in $\mathbb{R}^{3}$ in the standard manner, is exactly the origin.
This is a significant simplification of the sequence of results used by Kelner, tailored for our purposes. We note, for the interested reader, that Kelner gives a complete exposition of the required steps that can be used to prove the above statement only in the journal version of his work [Kel06b] and his thesis [Kel06a, Chapter 11]. Only a high-level outline is given in the original paper [Kel04].
This lemma gives us an analogous approximate circle packing construction for bounded-genus graphs after a sufficient number of hexagonal subdivisions. We will now use it to prove the following lemma:
Lemma 3.16.
Let $G$ be a triangulation of genus $g$. Then, there is some threshold $N\geq 0$ so that for all $k\geq N$, $\gamma^{(1)}(G^{(k)})\lesssim\gamma^{(3)}(G^{(k)})\lesssim\frac{g}{|V(G^{(k)})|}$.
Proof.
Apply Lemma 3.15 with $\varepsilon=O(1)$ to get $N$. Fix any $k\geq N$. Denote $n^{(k)}:=|V(G^{(k)})|$. We will now choose the values $y,f$ for the formulation of $\dot{\gamma}^{(3)}$ to obtain a similar argument to the proof of Theorem 3.5. For each $v\in V(G^{(k)})$, choose $s(v):=D(f^{(k)}(C_{v}))$ for some constant $\rho\in\Theta(1)$ to be chosen later. Choose $f(v)$ as the image of the centre of $C_{v}$ under $f^{(k)}$. Note that under this choice, $\sum_{v\in V(G^{(k)})}f^{(k)}(v)$ is exactly the origin.
For each $uv\in E(G^{(k)})$,
$||f(v)-f(u)||_{2}\leq D(f^{(k)}(C_{v}))+D(f^{(k)}(C_{u}))=s(u)+s(v).$
Hence, all the constraints are satisfied.
Note that $\sum_{v\in V(G^{(k)})}|f^{(k)}(v)|^{2}=n^{(k)}$. Let $V_{1}=\{v\in V(G^{(k)}):C_{v}\subset S_{1}\}$ and $V_{2}=\{v\in V(G^{(k)}):C_{v}\cap S_{2}\neq\emptyset\}$. Then,
$n^{(k)}\cdot\dot{\gamma}^{(3)}(G^{(k)})=\sum_{v\in V(G^{(k)})}D(f^{(k)}(C_{v}))^{2}=\sum_{v\in V_{1}}D(f^{(k)}(C_{v}))^{2}+\sum_{v\in V_{2}}D(f^{(k)}(C_{v}))^{2}$
$\lesssim\sum_{v\in V_{1}}A(f^{(k)}(C_{v}))+\varepsilon\lesssim g\cdot A(S^{2})+\varepsilon\lesssim g,$
so $\dot{\gamma}^{(3)}(G^{(k)})\lesssim\frac{g}{n^{(k)}}$. ∎
The combination of Lemma 3.12, Lemma 3.13, Lemma 3.14 and Lemma 3.16 together imply the following theorem:
Theorem 3.17.
Let $G$ be a graph with $n$ vertices, maximum degree $\Delta$, and genus at most $g$. Then,
$\frac{\lambda_{2}(G)}{\Delta}\leq\lambda_{2}^{*}(G)\leq\gamma^{(1)}(G)\lesssim\frac{g\log\Delta}{n}.$
Proof.
[…]
$\frac{\lambda_{2}(G)}{\Delta}\leq\lambda_{2}^{*}(G)\leq\gamma^{(1)}(G)\lesssim\gamma^{(1)}(G^{\prime})\cdot\Delta\log\Delta\leq\gamma^{(1)}(G^{\prime\prime})\cdot\Delta\log\Delta\lesssim\frac{g\log\Delta}{n},$
which implies the stated result. ∎
4 Bounds via 111-dimensional Extremal $L_{p}$-Embedded Spread
In this section, we will study a quantity we call the $d$-dimensional extremal $L_{2}$-embedded spread (with a focus on the 111-dimensional case). In more generality:
Definition 4.1.
For a graph $G$, we define the $d$-dimensional extremal $L^{p}$-embedded spread as:
$\begin{array}[]{r l}\overline{Q}^{(d)}_{p}(G):=&\begin{array}[]{c l}\max_{\begin{subarray}{c}y:V(G)\to\mathbb{R}_{\geq 0},\\ f:V(G)\to\mathbb{R}^{d}\end{subarray}}&\sum_{u,v\in V(G)}\left|\left|f(u)-f(v)\right|\right|_{p}^{p}\\ \textup{subject to}&\begin{array}[]{r c l l}||y||_{1}&\leq&1\\ y(u)+y(v)&\geq&\left|\left|f(u)-f(v)\right|\right|_{p}^{p}&\forall uv\in E(G).\end{array}\end{array}\end{array}$ (3)
This will be a key intermediate quantity that will allow us to relate $\overline{s}_{2}(G)$, $\gamma^{(1)}$, and (consequently) $\lambda_{2}^{*}$.
The case $\overline{Q}^{(1)}_{1}$ specifically corresponds to a notion directly related to a certain form of vertex separators used by Feige, Hajiaghayi, and Lee [FHL05, FHL08] (called “observable spread” by Lee [Lee17]). We will discuss this connection briefly later. However, $\overline{Q}_{2}^{(d)}$ will be more relevant for our methods. In particular, there is a relationship between the $d$-dimensional extremal $L^{2}$-embedded spread and $\gamma^{(d)}$:
Lemma 4.2.
For a graph $G$, $\gamma^{(d)}(G)=\frac{2n}{\overline{Q}^{(d)}_{2}(G)}$
Proof.
We will make a sequence of transformations to each optimization problem.
First, observe that the following is an equivalent formulation for $\gamma^{(d)}$:
$\begin{array}[]{r c c r c l l}\gamma^{(d)}(G)&:=&\min_{\begin{subarray}{c}f:V\to\mathbb{R}^{d}\\ y:V\to\mathbb{R}_{\geq 0}\end{subarray}}&\frac{\sum_{v\in V}y(v)}{\sum_{x\in V}||f(x)||_{2}^{2}}\\ &&\textup{subject to}&\sum_{v\in V}f(v)&=&\hskip 10.00002pt\overline{0}\\ &&&y(u)+y(v)&\geq&\hskip 10.00002pt||f(u)-f(v)||_{2}^{2}&\forall uv\in E\end{array}$
We claim that any function $f:V\to\mathbb{R}^{d}$ for which $\sum_{v\in V}f(v)=\overline{0}$, it holds that $2n\sum_{v\in V}||f(v)||_{2}^{2}=\sum_{u,v\in V}||f(u)-f(v)||_{2}^{2}$. To prove this claim, let $f_{i}(v)$ denote the $i$th coordinate of the vector $f(v)$. Then,
$\sum_{u,v\in V}||f(u)-f(v)||^{2}$ $=\sum_{u,v\in V}\sum_{i\in[n]}(f_{i}(u)-f_{i}(v))^{2}$
$=\sum_{u,v\in V}\sum_{i\in[n]}\left[f_{i}(u)^{2}+f_{i}(v)^{2}-2f_{i}(u)f_{i}(v)\right]$
$=\sum_{i\in[n]}\sum_{u\in V}\sum_{v\in V}\left[f_{i}(u)^{2}+f_{i}(v)^{2}-2f_{i}(u)f_{i}(v)\right]$
$=\sum_{i\in[n]}\sum_{u\in V}\left[\sum_{v\in V}\left(f_{i}(u)^{2}+f_{i}(v)^{2}\right)-2f_{i}(u)\sum_{v\in V}f_{i}(u)\right]$
$=\sum_{i\in[n]}\sum_{u\in V}\sum_{v\in V}\left[f_{i}(u)^{2}+f_{i}(v)^{2}\right]$
$=\sum_{i\in[n]}2n\sum_{x\in V}f_{i}(x)^{2}$
$=2n\sum_{x\in V}||f(v)||^{2}.$
We obtain:
$\begin{array}[]{r c c r c l l}\gamma^{(d)}(G)&:=&2n\cdot\min_{\begin{subarray}{c}f:V\to\mathbb{R}^{d}\\ y:V\to\mathbb{R}_{\geq 0}\end{subarray}}&\frac{\sum_{v\in V}y(v)}{\sum_{u,v\in V}||f(u)-f(v)||_{2}^{2}}\\ &&\textup{subject to}&\sum_{v\in V}f(v)&=&\hskip 10.00002pt\overline{0}\\ &&&y(u)+y(v)&\geq&\hskip 10.00002pt||f(u)-f(v)||_{2}^{2}&\forall uv\in E.\end{array}$
No vector $f(u)-f(v)$ is affected by translation, so the constraint $\sum_{v\in V}f(v)=\overline{0}$ is unnecessary, and hence the reciprocal of the inner fractional program here is exactly $\overline{Q}_{2}^{(d)}$, and the result follows. ∎
At this point, we’d like to relate $\overline{Q}_{2}^{(1)}(G)$ to $\overline{s}_{2}(G)$ (and consequently also $\overline{s}_{1}(G)$). There are two different natural ways to accomplish this: We could relate $\overline{s}_{1}(G)$ to $\overline{Q}_{1}^{(1)}(G)$ via a metric embedding, and $\overline{Q}_{1}^{(1)}(G)$ to $\overline{Q}_{2}^{(1)}(G)$ via Cauchy-Schwarz, getting a “loss” from each step. Alternatively, we can instead apply Cauchy-Schwarz directly to $\overline{s}_{2}(G)$ first, and then apply a metric embedding to the resulting quantity. This second method is analogous to some steps used by Biswal, Lee, and Rao to bound $\lambda_{2}$ [BLR10, proof of Theorem 5.1]. Each of these methods will result in the same bound, but we will present both since the final proof structure will be of independent interest. We start by relating $\overline{Q}_{1}^{(d)}$ and $\overline{Q}_{2}^{(d)}$:
Lemma 4.3.
For a graph $G$ with $n$ vertices, $\frac{(Q_{1}^{(d)}(G))^{2}}{2dn^{2}}\leq Q_{2}^{(d)}(G)$.
Proof.
Let $y,f$ be the optimal solution to $Q_{1}^{(d)}(G)$. We will construct a (possible sub-optimal) solution $y^{\prime},f^{\prime}$ to $Q_{2}^{(d)}(G)$. Let $y^{\prime}(v):=y(v)^{2}$, and let $f^{\prime}:=f/\sqrt{2}$. Then, by two applications of Cauchy-Schwarz,
$\begin{array}[]{r c l}\sum_{u,v\in V(G)}||f^{\prime}(u)-f^{\prime}(v)||_{2}^{2}&\geq&\sum_{u,v\in V(G)}\frac{1}{d}||f^{\prime}(u)-f^{\prime}(v)||_{1}^{2}\\ &\geq&\frac{\left(\sum_{u,v\in V(G)}||f^{\prime}(u)-f^{\prime}(v)||_{1}\right)^{2}}{dn^{2}}\\ &=&\frac{\left(\sum_{u,v\in V(G)}||f(u)-f(v)||_{2}\right)^{2}}{2dn^{2}}.\end{array}$
so it remains only to show the constraints are satisfied. For any $u,v\in V$,
$\begin{array}[]{r c l}||f^{\prime}(u)-f^{\prime}(v)||_{2}^{2}&\leq&||f^{\prime}(u)-f^{\prime}(v)||_{1}^{2}\\ &=&||f(u)-f(v)||_{1}^{2}/2\\ &\leq&(y(u)+y(v))^{2}/2\\ &\leq&y(u)^{2}+y(v)^{2}=y^{\prime}(u)+y^{\prime}(v).\end{array}$
And finally,
$||y^{\prime}||_{1}=||y||_{2}\leq||y||_{1}\leq 1,$
so $f^{\prime}$ and $y^{\prime}$ form a valid solution to $Q_{2}^{(d)}$ with the claimed objective value. ∎
Next, we show the alternative step in which we could apply Cauchy-Schwarz:
Lemma 4.4.
For a graph $G$,
$\frac{\left[\overline{s}_{2}(G)\right]^{2}}{n^{2}}\leq\sup_{\omega:V(G)\to\mathbb{R}_{\geq 0},||\omega||_{2}\leq 1}\sum_{u,v\in V(G)}\left[d_{G}^{\omega}(u,v)\right]^{2}.$
Proof.
The result follows directly from Cauchy-Schwarz. ∎
We now highlight the metric embedding steps of each of the two methods:
Lemma 4.5.
For a graph $G$,
$\left[\overline{\alpha}(G)\right]^{2}\cdot\overline{Q}_{2}^{(1)}(G)\gtrsim\sup_{\omega:V(G)\to\mathbb{R}_{\geq 0},||\omega||_{2}\leq 1}\sum_{u,v\in V(G)}\left[d_{G}^{\omega}(u,v)\right]^{2}.$
Proof.
Apply Theorem 2.15 to get a non-expansive embedding $f^{\prime}:V\to\mathbb{R}$ from $d_{G}^{\omega}$ to $\mathbb{R}$ with average 222-distortion at most $C_{2}\cdot 2\cdot\left[\overline{\alpha}(G)\right]^{2}$ for a universal constant $C_{2}$. Using $y(v):=\omega(v)^{2}$ and $f(v):=f^{\prime}(v)/2$, we claim the constraints of $\overline{Q}_{2}^{(1)}(G)$ are satisfied: For the first constraint, $||y||_{1}=\sum_{v\in V}\omega(v)^{2}=||\omega||_{2}^{2}\leq 1$. For the remaining constraints,
$y(u)+y(v)=\omega(u)^{2}+\omega(v)^{2}\geq\frac{1}{2}\left[\omega(u)+\omega(v)\right]^{2}\geq\frac{1}{2}|f^{\prime}(u)-f^{\prime}(v)|^{2}\geq|f(u)-f(v)|^{2},$
and so $y,f$ forms a feasible solution to $\overline{Q}_{2}^{(1)}(G)$. Finally, it follows from the 222-distortion bound that
$\sum_{u,v\in V(G)}\left[d_{G}^{\omega}(u,v)\right]^{2}\lesssim\left[\overline{\alpha}(G)\right]^{2}\cdot\sum_{u,v\in V(G)}|f^{\prime}(u)-f^{\prime}(v)|^{2}\lesssim\left[\overline{\alpha}(G)\right]^{2}\cdot\sum_{u,v\in V(G)}|f(u)-f(v)|^{2}.$
The alternative place to employ the metric embedding step is to jump from $\overline{s}_{2}(G)$ to $\overline{Q}_{1}^{(1)}$:
Lemma 4.6.
For a graph $G$,
$\left[\overline{\alpha}(G)\right]^{2}\cdot\overline{Q}_{1}^{(1)}(G)\gtrsim\cdot\overline{s}_{2}(G).$
Proof.
The proof is essentially the same as that of Lemma 4.5, except that a non-expansive embedding with bounded 111-distortion is used. ∎
By combining either Lemma 4.6 and Lemma 4.3, or Lemma 4.4 and Lemma 4.5, we obtain the following result:
Lemma 4.7.
Let $G$ be a graph,
$\left[\overline{\alpha}(G)\right]^{2}\cdot\overline{Q}_{2}^{(1)}(G)\gtrsim\frac{\left[\overline{s}_{2}(G)\right]^{2}}{n^{2}}.$
We obtain a sequence of important corollaries from combining Lemma 4.7 and Lemma 4.2. Most importantly, we obtain the following theorem:
Theorem 4.8.
For a graph $G$ with maximum degree $\Delta$,
$\frac{\lambda_{2}(G)}{\Delta}\leq\lambda_{2}^{*}(G)\leq\gamma^{(n)}(G)\leq\gamma^{(1)}(G)\lesssim\left[\overline{\alpha}(G)\right]^{2}\cdot\frac{n^{3}}{\left[\overline{s}_{2}(G)\right]^{2}}.$
Proof.
The first three inequalities follow from 1.2, the strong duality of $\lambda_{2}^{*}(G)$ and $\gamma^{(n)}(G)$, and a simple relaxation. The last one follows from the combination of Lemma 4.7 and Lemma 4.2. ∎
All of Theorem 2.1, Theorem 2.4, Theorem 2.5, and Theorem 2.6 are strengthened by this result. Using the results in Section 2.2 and Section 2.3 we obtain:
Corollary 4.9.
Let $G$ be a graph with $n$ vertices. If $G$ has genus at most $g$, then
$\frac{\lambda_{2}(G)}{\Delta}\leq\lambda_{2}^{*}(G)\leq\gamma^{(1)}(G)\lesssim\frac{g(\log g)^{2}}{n}.$
If $G$ has no $K_{h}$ minor, then
$\frac{\lambda_{2}(G)}{\Delta}\leq\lambda_{2}^{*}(G)\leq\gamma^{(1)}(G)\lesssim\frac{(h\log h\log\log h)^{2}}{n}.$
If $G\in\text{rig}(\hat{G})$ for a graph $\hat{G}$ with no $K_{h}$ minor, then
$\frac{\lambda_{2}(G)}{\Delta}\leq\lambda_{2}^{*}(G)\leq\gamma^{(1)}(G)\lesssim\frac{\Delta h^{6}(\log\log h)^{2}}{n}.$
Moreover, this also induces a yet another proof of the planar separator theorem via Theorem 1.5, although this proof does not imply that reweighted spectral partitioning can be used as a polynomial time algorithm for the planar separator theorem without the result in the next section.
Note that we can avoid the constraints on $n$ for the bounds in Section 2.2 since $\gamma^{(1)}(G)\lesssim 1$ (see 1.3).
5 A Refined Cheeger Inequality
Almost all the bounds we have found apply directly to $\gamma^{(1)}$, save for those in Section 3.1. Hence, all such bounds also imply the existence of small separators, but the proofs are non-constructive. We are also interested in (efficient) algorithms for computing such separators. Unfortunately, naively applying the construction for Theorem 1.4 would result in a larger separator by a factor of $\sqrt{\log\Delta}$. Sometimes this is quite good, but when $\Delta$ is large we can often do better. One of our goal was to eliminate the necessary dependence on maximum degree, and so ideally we would like to further reduce this factor in cases where $\Delta$ is large. In this section, we (briefly) show that this dimension reduction step can be further mitigated for special graph classes. As a consequence, we are also able to recover a refined Cheeger inequality.
Lemma 5.1.
For any graph $G$, and any $d,p\geq 1$, $\overline{Q}_{p}^{(d)}(G)\leq C_{p}\cdot\left[\alpha(G)\right]^{p}\cdot\overline{Q}_{p}^{(1)}(G),$ where $C_{p}$ is a constant dependent only on $p$.
The proof of this lemma will be analogous to that of Lemma 4.5, which made use of the non-expansive embedding given by Theorem 2.15. However, there are two key differences: First, the edge-weight modulus $\alpha$ is used rather than the vertex-weight modulus $\overline{\alpha}$. Second, the embedding given by Theorem 2.15 is non-expansive for every pair of vertices, but the proof of Lemma 4.5 only makes use of the fact that it applies to adjacent pairs. Hence, a similar a bound could have also applied to a modified optimization problem with a larger objective value. However, here it is critical that the constraints in the optimization problem of $\overline{Q}_{p}^{(1)}(G)$ are only for adjacent pairs.
Proof.
Let $f,y$ be the optimal solution to $Q_{p}^{(d)}$. Let $\omega:E\to\mathbb{R}_{\geq 0}$ be edge weights so that $\omega(uv)=||f(u)-f(v)||_{p}$. Let $d_{\omega}$ be the shortest-path metric over $G$ induced by the edge weights $\omega$. Apply Theorem 2.15 to $d_{\omega}$ to obtain a non-expansive embedding $f^{\prime}:V\to\mathbb{R}$ of $(V,d_{\omega})$ with $p$-distortion at most $C_{p}\cdot\left[\alpha(G)\right]^{p}$. We claim that $f^{\prime},y$ forms a feasible solution to $Q_{p}^{(1)}$. Since $y$ is unchanged, it suffices to only check the constraint on each edge. For each edge $uv\in E$, $d_{\omega}(u,v)=||f(u)-f(v)||_{p}$. Moreover, $d_{\omega}(u,v)\geq|f^{\prime}(u)-f^{\prime}(v)|$ by the non-expansive guarantee of the embedding, so we obtain $|f^{\prime}(u)-f^{\prime}(v)|\leq y(u)+y(v)$ as desired. Finally, by the $p$-distortion guarantee, we obtain the bound $\overline{Q}_{p}^{(d)}(G)\leq C_{p}\cdot\left[\alpha(G)\right]^{p}\cdot\sum_{u,v\in V(G)}||f(u)-f(v)||_{p}^{p}$. ∎
By applying Lemma 4.2, we can immediately obtain the following dimension-reduction theorem:
Theorem 5.2.
For a graph $G$, and all $d\geq 1$, $\gamma^{(1)}(G)\lesssim\left[\alpha(G)\right]^{2}\cdot\gamma^{(d)}(G)$.
Finally, we can use this to refine Theorem 1.4:
Theorem 5.3 (Refined Cheeger Inequality for Vertex Expansion).
For a graph $G$ with maximum degree $\Delta$,
$\frac{\psi(G)^{2}}{\min\{\log\Delta,\alpha(G)^{2}\}}\lesssim\lambda_{2}^{*}(G)\lesssim\psi(G).$
It should be noted that Theorem 1.4 is inherently constructive. However, Theorem 5.3 is not: There is not any known method that can approximate $\alpha(G)$. Fortunately, in the case of $K_{h}$-minor-free graphs, the recent bound of $\alpha(G)\in\operatorname{\mathcal{O}}(\log h)$ is constructive and provides a polynomial-time algorithm, even if the value of $h$ is not known [CF25]. In other words, Corollary 1.7 is constructive (in the sense of implying a polynomial-time algorithm) without parametrization.
6 A High-Level View of Congestion-based Bounds
The results in the previous sections are built on combinations of results and techniques from a variety of previous works, as well as a couple new methods. In this section, we will show that a high-level overview of the whole sequence of proof methods results in some key insight towards answering 2.2 (and similar tightenings for other bounds), as well as a deeper understanding of the structure of all these quantities. For this section, fix a graph $G=(V,E)$ with $n$ vertices. We will primarily focus on the methods in Section 4, and the case of bounded-genus graphs. Analogous observations hold for $K_{h}$-minor-free graphs and the methods in Section 3. Henceforth, assume $G$ has genus at most $g$.
As a first step, we define the other main quantities used by Biswal, Lee, and Rao [BLR10] for their eigenvalue bounds:
Definition 6.1.
[…]
There are several important properties of these quantities shown by Biswal, Lee, and Rao [BLR10]. In particular, they showed the following results:
Theorem 6.2 ([BLR10, proof of Theorem 3.1]).
For a graph $G$ with genus at most $g$, $\text{con}_{2}^{*}(G)\gtrsim\frac{n^{2}}{\sqrt{g}}$.
This theorem is proven using a standard crossing number inequality argument.
Theorem 6.3 ([BLR10, Lemma 2.1]).
For a graph $G$ with $n$ vertices, $\text{con}_{2}^{*}(G)\leq\text{con}_{2}(G)+n^{\frac{3}{2}}$.
This theorem is proven using a probabilistic argument.
Theorem 6.4 ([BLR10, Theorem 2.2]).
For a graph $G$, $\text{con}_{2}(G)=\overline{s}_{2}(G)$.
This theorem is proven using Lagrangian duality and Slater’s condition for convex optimization.
We are now ready to give the full overview of the techniques used to obtain the bounds on $\lambda_{2}^{*}$(G) and $\gamma^{(1)}(G)$:
$\text{cr}_{g}(K_{n})\gtrsim\frac{n^{4}}{g}$ $\text{con}_{2}^{*}(G)\gtrsim\frac{n^{2}}{\sqrt{g}}$ $\text{con}_{2}(G)\gtrsim\frac{n^{2}}{\sqrt{g}}$ $\lambda_{2}^{*}(G)\lesssim(\log g)^{2}\frac{g}{n}$ $\overline{s}_{2}(G)\gtrsim\frac{n^{2}}{\sqrt{g}}$ $\sup_{\begin{subarray}{c}\omega:V\to\mathbb{R}_{\geq 0}\\ ||\omega||_{2}\leq 1\end{subarray}}\sum_{u,v\in V}\left[d_{G}^{\omega}(u,v)\right]^{2}\gtrsim\frac{n^{2}}{g}$ $\gamma^{(n)}(G)\lesssim(\log g)^{2}\frac{g}{n}$ $\overline{Q}_{1}^{(1)}(G)\gtrsim\frac{1}{\sqrt{g}\log g}$ $\overline{Q}_{2}^{(1)}(G)\gtrsim\frac{n^{2}}{g(\log g)^{2}}$ $\gamma^{(1)}(G)\lesssim(\log g)^{2}\frac{g}{n}$ (Theorem 6.2) Integrality gap (Theorem 6.3) Strong duality (Theorem 6.4) Metric embedding (Lemma 4.6) Cauchy-Schwarz (Lemma 4.3) (Lemma 4.2) Cauchy-Schwarz (Lemma 4.4) Metric embedding (Lemma 4.5) Strong duality Relaxation
Recall also that to obtain a cut from $\lambda_{2}^{*}(G)$ requires using a lossy metric embedding (with $\min(\alpha(G)^{2},\log\Delta)$ loss) to get back from $\gamma^{(n)}(G)$ to $\gamma^{(1)}(G)$. This gives us a second diagram for polynomial-time computable quantities, starting from $\lambda_{2}^{*}(G)$:
$\lambda_{2}^{*}(G)\lesssim(\log g)^{2}\cdot\frac{g}{n}$$\gamma^{(n)}(G)\lesssim(\log g)^{2}\cdot\frac{g}{n}$$\gamma^{(1)}(G)\lesssim(\log g)^{4}\cdot\frac{g}{n}$$\psi(G)\lesssim\frac{(\log g)^{2}\sqrt{g}}{\sqrt{n}}$trong duality Metric embedding/dimension reduction (Theorem 5.2) 1 11-dimensional Cheeger inequality for $\psi$ (Theorem 1.5)
There are a number of symmetries in these diagrams. In particular, $\lambda_{2}^{*}$ is related to $\gamma^{(n)}$ by a form of strong duality, just as $\text{con}_{2}$ and $\overline{s}_{2}$ are related by strong linear programming duality. Likewise, $\overline{s}_{2}$ is related to $\overline{Q}_{2}^{(1)}$ by two different methods that each involve a metric embedding, just as $\gamma^{(n)}$ can be related to $\gamma^{(1)}$ via the same metric embedding. It is possible these are all coincidental, but they do suggest the possibility of a deeper relationship between all these quantities than what we have discovered so far.
One notable feature of these symmetries is that we incur two identical metric embedding losses when we actually compute a cut. Both come from essentially the same source, so this seems quite unnecessary. One of these losses is likely necessary: A (general) constant-factor dimension-reduction from $\gamma^{(n)}$ to $\gamma^{(1)}$ would violate the small-set expansion conjecture [LRV13], so we expect some kind of loss at this step. The loss in the upper bound seems unnecessary though. In fact, it seems as if it comes from the fact that we pass through the low-dimensional $\gamma^{(1)}$. All current methods for bounding $\lambda_{2}$ (and $\lambda_{2}^{*}$) in bounded-genus graphs make use of a low-dimensional quantity. Moreover, some methods start by giving a bound on a polynomial-time computable quantity like $\overline{s}_{2}$. If any polynomial-time quantity could be used to approximate $\gamma^{(1)}$ within a constant factor in both directions, then the small-set expansion conjecture would be disproved. Although we have only focused on bounding $\gamma^{(1)}$ in one direction, this (in combination with the symmetry of the current bounds) may suggest that resolving 2.2 (and strengthening it to $\lambda_{2}^{*}$) could require using a higher-dimensional construction. Unfortunately, jumping straight from $\overline{s}_{2}$ to $\gamma^{(n)}$ with some kind of higher-dimensional non-expansive metric embedding seems to require some non-trivial average distortion in general [Rab08], so different methods for this would likely need to be developed. Alternatively, it is quite possible that a bound on $\gamma^{(1)}$ could be found using some structure that also is not believed to be approximable. With these observations in mind, we propose two new conjectures:
Conjecture 6.5.
For a graph $G$ of $n$ vertices with genus $g$, the second smallest eigenvalue of the Laplacian matrix of $G$ has $\lambda_{2}^{*}(G)\lesssim\frac{g}{n}$.
Conjecture 6.6.
For a graph $G$ of $n$ vertices with no $K_{h}$-minor, so that $n\gtrsim h$, the second smallest eigenvalue of the Laplacian matrix of $G$ has $\lambda_{2}^{*}(G)\lesssim\frac{h^{2}}{n}$.
We are specifically not conjecturing that the same bounds hold for $\gamma^{(1)}$, although we have not ruled out this possibility either. Determining if such a bound is even possible is an interesting open question as well.
The first of these conjectures is a slight strengthening of 2.2, and the second is a strengthening of another conjecture of Spielman and Teng [ST07, Conjecture 2]. However, neither of these would completely get rid of the metric embedding/dimension reduction loss in polynomial-time computable cuts that arises during reweighted spectral partitioning via the refined Cheeger-style inequality stated in Theorem 5.3.
An interesting observation is that the purely metric embedding-based methods of bounding $\gamma^{(1)}$ result in vertex expansion certifications no smaller than those obtained by the algorithm of Feige, Hajiaghayi, and Lee [FHL05, FHL08] (who showed that vertex expansion is directly related to $Q_{1}^{(1)}$, more explicitly stated by Lee [Lee17, Theorem 1.7]) and that this would remain true of the polynomial-time computable bounds even if the conjectures above held. This suggests that one of the strengths of reweighted spectral partitioning (and even traditional spectral partitioning) is that it is able to (optionally) replace this dependence on metric embeddings with a (possibly better) dependence on the maximum degree. Granted, it is also possible that an analogous bound involving $\Delta$ could be obtained via more direct methods (i.e., direct approximations of vertex expansion) in the future.
7 Conclusion and Future Work
We have given bounds for $\lambda_{2}^{*}$ (and often, $\gamma^{(1)}$) for a large number of graph classes, as well as a new relationship between $\lambda_{2}^{*}$ and $\gamma^{(1)}$. There are a number of natural questions left open, besides those previously discussed. Some of the most far-reaching questions are as follows:
  • The new relationship between $\lambda_{2}^{*}(G)$ and $\gamma^{(1)}(G)$ incorporating $\alpha(G)$, stated in Theorem 5.3, is in some ways similar to an approximation algorithm of Feige, Hajiaghayi, and Lee [FHL05, FHL08]. In particular, it is natural to ask if vertex expansion may be $\operatorname{\mathcal{O}}(\alpha(G))$-approximated in almost-linear time by strengthening the main result of Lau, Tung, and Wang [LTW24], which gives an almost-linear time $\operatorname{\mathcal{O}}(\sqrt{\log n})$-approximation. In particular, both results build on the metric rounding method Arora, Rao, and Vazirani [ARV08, ARV09].
  • Similarly, it would be interesting to determine if Theorem 5.3 could be made directly constructive, ideally with a simple algorithm. Specifically, this would remove the algorithmic requirements of bounds on $\alpha(G)$. Such a construction would not necessarily need to compute $\alpha(G)$.
  • While the $\log\Delta$ factor in Theorem 5.3 is likely essentially necessary in general graphs [LRV13], we have shown that at least one other possible factor may be used ($\alpha(G)$) that is stronger in some cases. It would be interesting to know if other factors may be used in this dimension-reduction bound, such as the logarithm of average degree, or the logarithm of the square root of the average squared degree. It would also be interesting to know if there is a natural way to unify all these bounds in some manner, particularly if such a unification could result in one simple algorithm.
  • Can these methods be leveraged to efficiently obtain near-optimal balanced separators? The methods of Orecchia and Vishnoi [OV11], and Orecchia, Sachdeva, and Vishnoi [OSV12] seem to be promosing avenues for this possibility.
In addition, there is the practical question of whether or not reweighted spectral partitioning produces better results than other techniques in real-world settings. The methods for reweighted spectral partitioning are still quite new, and has primarily been studied primarily from a theoretical perspective. Hence, this significantly differs from the history of spectral partitioning, which was shown to have practical significance before being shown to have algorithmic guarantees.
Experimentally testing the practicality of the method would first require a complete implementation, so we outline the simplest-known methods for each step: Boyd, Diaconis and Xiao [BDX04] provided a practical and simple subgradient method for computing $\lambda_{2}^{*}$ that can quite easily be implemented with GPU-acceleration in a modern library. Unfortunately, their method as-stated is a primal method only, and does not obtain a dual solution, so a slightly modified form would be necessary. The simplest dimension-reduction technique is that of Kwok, Lau, and Tung [KLT22], stated in Theorem 1.6. Unfortunately, the best-known bounds for $\alpha(G)$ (particularly for forbidden-minor graphs) are not currently as algorithmically simple or practical, nor is their application in Theorem 5.2 and Theorem 5.3. Finally, the simplest cut (rounding) algorithm is also likely that of Kwok, Lau, and Tung [KLT22, Section 3.2.3], although its constants could certainly use some improvement.
8 Acknowledgements
Thank you to Therese Biedl for pointing out that not every embedded graph in the torus can be simply triangulated, which led to the $K_{5}$ example discussed at the start of Section 3.2.
Appendix A Graph Class Definitions
There are a number of graph classes that are relevant to our results. The most notable is the well-known class of planar graphs, which are graphs that can be drawn on the plane without crossing edges. Most of the other classes we will consider generalize planar graphs in some manner.
Definition A.1.
A graph $G$ is said to have genus $g$ if it can be embedded into an orientable surface of genus $g$, and it cannot be embedded into an orientable surface of smaller genus. Planar graphs are exactly those that have genus 0.
Up to homeomorphism, there is only one orientable surface of genus $g$, for any particular $g$. For $g=0$, it is the sphere. The orientable surface of genus $g>0$ can be obtained from the orientable surface of genus $g-1$ by attaching a “handle”. Alternatively, it can be expressed as $g$ “doughnuts” merged together. See Figure 6 for examples of these surfaces, and see Figure 7 for an example of a graph drawn on a genus-333 surface.
Figure 6: Examples of (left to right) a genus-111 surface, a genus-222 surface, a genus-333 surface, and a genus 101010 surface.
Figure 7: An example of a graph (left) and a drawing of the graph on a genus-333 surface (right).
Definition A.2.
For a graph $G$ and a graph $H$, we say that $H$ is a minor of $G$ if it can be obtained from $G$ by a sequence of edge deletions, vertex deletions, and edge contractions (that is, merging the two vertices incident to an edge into one). Equivalently, $H$ is a minor of $G$ if there exists a mapping $f:V(G)\to V(H)$ so that if $uv\in E(H)$, then $\left(f^{-1}(u)\times f^{-1}(v)\right)\cap E(G)\neq\emptyset$.
We say a graph $G$ is $H$-minor-free if it does not contain $H$ as a minor.
Definition A.3.
For graphs $G$ and $\hat{G}$, $\hat{G}$ is said to be a region intersection graph over $G$ if there is a mapping $f:V(\hat{G})\to 2^{V(G)}$ so that $G[f(v)]$ is connected for all $v\in V(\hat{G})$, and $uv\in V(\hat{G})$ if and only if $f(u)\cap f(v)\neq\emptyset$. We denote the set of region intersection graphs over $G$ as $\text{rig}(G)$. When $G$ is a planar graph, $\hat{G}\in\text{rig}(G)$ is said to be a string graph.
Figure 8: An example of strings in the plane (left), and their string graph (right).
A string graph is normally defined in a slightly different way, as the “geometric intersection graph” (a graph with vertices defined by geometric objects and edges defined by intersections) of “strings” (curves in the plane). See Figure 8 for an example of this construction. This is equivalent to a region intersection graph over a plane graph whose regions are specifically paths, which itself is equivalent to the above definition via some basic transformations.
The term “region intersection graph” in the literature usually refers to cases in which the “latent” graph $G$ has a forbidden $K_{h}$ minor, for some $h$. This generalizes the cases of bounded-genus and forbidden-minor graphs: If $G$ has no $K_{h}$ minor, and $\dot{G}$ is the barycentric subdivision of $G$ (which divides each edge with one additional vertex), then $G\in\text{rig}(\dot{G})$, and $\dot{G}$ also has no $K_{h}$ minor.
We will also consider two much more fundamentally geometric graph classes:
Definition A.4.
A graph $G$ is said to be a $d$-dimensional $k$-ply neighbourhood system if it is the geometric intersection graph of a collection of $d$-dimensional balls $B_{1},\dots,B_{n}$ with ply at most $k$ for all points almost surely.
Definition A.5.
A $d$-dimensional $k$-nearest neighbour graph is a graph $G$ with $n$ vertices $v_{1},\dots,v_{n}$ corresponding to points $p_{1},\dots,p_{n}\in\mathbb{R}^{d}$ so that an edge $v_{i}v_{j}$ exists if and only if $p_{i}$ is among the $k$-nearest neighbours of $p_{j}$, or $p_{j}$ is among the $k$-nearest neighbours of $p_{i}$.
Appendix B Vertex Expansion is NP-hard
Although vertex-expansion is widely-known to be NP-hard (see e.g. [LRV13]), we have not been able to find an explicit proof in the literature, so we will provide one here for completeness. Kaibel [Kai04, Theorem 2] faced a similar issue with edge expansion, and presented a proof of hardness of edge expansion based on the hardness of maximum cut.
In this section, we will give a proof of the hardness of vertex expansion based on the hardness of edge expansion. Specifically, we consider the decision problem variants of each, where we ask if there is a subset $S$ of the vertices with edge or vertex expansion at most $\alpha$.
Theorem B.1.
The decision problem form of vertex expansion is NP-hard.
Proof.
We will show that edge expansion can be reduced to vertex expansion. Let $G=(V,E)$ be a graph with $n=|V|$ vertices, and let $\alpha>0$. These form the “input” to the edge expansion problem. We will show that there is a graph $G^{\prime}$ and a value $\beta$ so that $G$ has edge expansion at most $\alpha$ if and only if $G^{\prime}$ has vertex expansion at most $\beta$. At a high-level, $G^{\prime}$ will be formed by bisecting each edge of $G$, and then replacing each vertex that was not a bisection vertex with a clique of size $k$. For a sufficiently large value of $k$ (which is still polynomial in $n$), we will be able to show that an upper bound on the vertex expansion of $G^{\prime}$ can be “rounded” to an upper bound on the edge expansion of $G$. In particular, for any $\varepsilon>0$, we may assume $k$ to be a sufficiently large polynomial in $n$ so that for any $S\subset V$, $|S|k\leq|S|k+|E[S]|\leq|S|k+|E|\leq(|S|+\varepsilon)k$, We will place further requirements on $k$ (including the choice of $\varepsilon$) later.
We now give the explicit construction of $G^{\prime}$ and $\beta$, for a parameter $k$. where $E[S]$ denotes the set of edges in the induced subgraph of $G$ with the vertices $S$. Let $V_{1}$ be a set with $k$ copies of each $v\in V$ labelled $v^{1},\dots,v^{k}$, for a value $k$ to be chosen later. Let $V_{2}=E$. Let $E^{\prime}=\{ev^{i}:e\in E,v\in V\}$. Let $G^{\prime}$ be a graph with vertices $V^{\prime}=V_{1}\cup V_{2}$ and edges $E^{\prime}$. We choose $\beta:=\frac{\alpha}{k}$.
Before proving that $\phi(G)\leq\alpha$ if and only if $\psi(G^{\prime})\leq\beta$, we will prove some useful structure.
We first claim that if there is a subset $S\subset V$, then there is a subset $S^{\prime}\subset V^{\prime}$ with $|\partial(S^{\prime})|=|\delta(S)|$ and $|S^{\prime}|=|S|k+|E[S]|$, where $E[S]$ denotes the edges in the induced subgraph $G[S]$. In particular, $S^{\prime}$ is exactly the subset of $V^{\prime}$ formed by $E[S]\subset V_{2}$ and $\{v^{i}:v\in S\}\subset V_{1}$.
We next claim that if there is a subset $S^{\prime}\subset V^{\prime}$, then there is a second subset $S^{\prime\prime}\subset V^{\prime}$ with with $\frac{|\partial(S^{\prime\prime})|}{|S^{\prime\prime}|}\leq\frac{|\partial(S^{\prime})|}{|S^{\prime}|}$, and a subset $S\subset V$ so that $|\delta(S)|=|\partial(S^{\prime})|$ and $|S|k+|E[S]|=|S^{\prime\prime}|$. Specifically, we choose $S^{\prime\prime}$ to be the set that contains, for each $v\in V$, every element of $\{v^{1},\dots,v^{k}\}$ if and only if $S$ contains any element of this set, and furthermore $S^{\prime\prime}$ contains $e=uv\in V_{2}$ if and only if $S$ contains some $u^{i}$ and $v^{j}$. Note that this corresponds exactly to the sets that can be mapped back to sets $S\subset S$ with the desired correspondence via the inverse of the transformation in the previous paragraph.
We now prove that $\frac{|\partial(S^{\prime\prime})|}{|S^{\prime\prime}|}\leq\frac{|\partial(S^{\prime})|}{|S^{\prime}|}$: We can make all the modifications for vertices in $V_{1}$ “first”, and count the changes to the fraction, and then make the modifications for the vertices in $V_{2}$ and count the remaining changes. For the vertices $v^{i}\in V_{2}$ in $S^{\prime\prime}$ but not in $S^{\prime}$, they contributed at least 111 to the numerator and 0 to the denominator under $S^{\prime}$, and under $S^{\prime\prime}$ they contribute 0 to the numerator and 111 to the denominator. Henceforth, we may assume that $S^{\prime}$ does not “cut” individual sets $\{v_{1},\dots,v_{k}\}$. For the vertices $e=uv\in V_{2}$ not in $S^{\prime}$ but in $S^{\prime\prime}$, such vertices must have all their neighbours included in $S^{\prime\prime}$. Hence, they contributed at least 111 to the numerator and 0 to the denominator under $S^{\prime}$, and under $S^{\prime\prime}$ they contribute 0 to the numerator and 111 to the denominator. For the vertices $e=uv\in V_{2}$ in $S^{\prime}$ but not in $S^{\prime\prime}$, such vertices contributed 111 to the denominator and $k$ to the numerator under $S^{\prime}$, and under $S^{\prime\prime}$ they contribute 111 to the numerator and 0 to the denominator. We claim that each of these replacements (whose sequence eventually results in $S^{\prime\prime}$) decreases the value of the fraction. Let $a$ be the total size of $|\partial(S^{\prime})|$ excluding a clique of vertices in $V_{1}$, and let $b\geq 1$ be the total size of $|S^{\prime}|$ excluding a bisection vertex in $V_{2}$. We need to prove that $\frac{a+k}{b+1}\geq\frac{a+1}{b}$. If we assume $k\geq n^{2}+n+1$ (which is a polynomial in $n$) then, $a\leq n^{2}$ and $b\leq n$, so $k\geq n^{2}+n+1\implies kb\geq n^{2}+n+1\geq a+b+1\implies ab+kb\geq ab+a+b+1\implies(a+k)b\geq(a+1)(b+1)\implies\frac{a+k}{b+1}\geq\frac{a+1}{b}$.
Will will now finally prove that $\phi(G)\leq\alpha$ if and only if $\psi(G^{\prime})\leq\beta=\frac{\alpha}{k}$, In particular, we claim that, for every $\alpha\geq 0$, there exists $S\subset V$ so that $\frac{|\delta(S)|}{|S|}\leq\alpha$ in $G$ if and only if there exists $S^{\prime}$ so that $\frac{|\partial(S^{\prime})|}{|S^{\prime}|}\leq\frac{\alpha}{k}$ in $G^{\prime}$. Note that there are only a polynomial number of possible values of $\frac{|\delta(S)|}{|S|}$ (specifically, there are at most $|E|\cdot|V|$ possible values), so we may assume $\alpha$ is one such value, and that $k$ is chosen so that $\alpha\cdot(1+\varepsilon)$ is strictly less than any larger such value.
In the forward direction, assume $\frac{|\delta(S)|}{|S|}\leq\alpha$. Use $S^{\prime}$ so that $|\partial(S^{\prime})|=|\delta(S)|$ and $|S^{\prime}|=|S|k+|E[S]|$. Then, $\frac{|\partial(S^{\prime})|}{|S^{\prime}|}=\frac{|\delta(S)|}{|S|k+|E[S]|}\leq\frac{|\delta(S)|}{|S|k}\leq\frac{\alpha}{k}$.
In the backward direction, assume $\frac{|\partial(S^{\prime})|}{|S^{\prime}|}\leq\frac{\alpha}{k}$. Then, use the prior construction to find $S$ so that $\frac{|\delta(S)|}{|S|k+|E[S]|}\leq\frac{|\partial(S^{\prime})|}{|S^{\prime}|}\leq\alpha k$. Recall that $|S|k+|E[S]|\leq(|S|+\varepsilon)k$, so $\frac{|\delta(S)|}{(|S|+\varepsilon)k}\leq\alpha k$ and thus $\frac{|\delta(S)|}{|S|}\leq\alpha(1+\varepsilon)$. Hence, by the earlier choice of $k$, we also get $\frac{|\delta(S)|}{|S|}\leq\alpha$.
Appendix C Deferred Proof of Subdivision Lemma
We now prove Lemma 3.14 by adapting an argument of Kelner [Kel04, Kel06b, Kel06a]. The proof will be similar to, but not quite the same as, that of Lemma 3.11. The primary differences will be that the paths corresponding to edges will be randomly sampled, and that the “uniformity” is only approximate. It seems quite plausible that a more general form of Lemma 3.11 could be extracted from the below proof, but the most straightforward proof method using these random paths results in an extra factor of $\Delta$. This factor is perfectly fine for proving Lemma 3.14, but poses issues for a generalized form of Lemma 3.11.
Proof.
Denote $G=(V,E)$, $n:=|V|$, $H:=G^{(k)}=(V^{\prime},E^{\prime})$, and $n^{\prime}:=|V^{\prime}|$. For the remainder of this proof, we will use the notation $\operatorname{\mathcal{O}}_{\Delta}()$ and $\Theta_{\Delta}()$ to hide polynomial factors in $\Delta$. Note that the maximum degree of $H$ is $\max\{6,\Delta\}$, since no new vertices of degree >6>6> 6 are added from a hexagonal subdivision, nor does any existing vertex increase have its degree changed. Each edge in $G$ is split into $2^{k}$ pieces in $H$, and each triangle in $G$ is partitioned into $4^{k}$ triangles in $H$. The number of triangles incident to any vertex of degree >6>6> 6 remains constant during subdivision. Hence, $\frac{n^{\prime}}{n}\in\Theta_{\Delta}(4^{k})$. Let $y_{H},f_{H}$ denote the optimal solution to $\gamma^{(1)}(H)$. Assume without loss of generality that $\sum_{x^{\prime}\in V^{\prime}}|f_{H}|^{2}=1$. This assumption will allow us to slightly simplify some later steps.
Kelner [Kel04, Kel06a, Proof of Lemma 5.2] is able to show that there exists random variables $\pi_{V}:V\to V^{\prime}$ and $\pi_{E}:E\to\{\text{paths through $H$}\}$ with the following properties:
  1. 1.
    For each $uv\in E$, $\pi_{E}(uv)$ is a path in $H$ from $\pi_{V}(u)$ to $\pi_{V}(v)$.
  2. 2.
    For each $v\in V$, $\pi_{V}(v)$ is a uniform distribution over its support, which we denote $p(v)$. Moreover, every vertex $v^{\prime}\in V^{\prime}$ is contained in some $p(v)$ for a $v\in V$.
  3. 3.
    For each $u\neq v\in V$, $\pi_{V}(u)$ and $\pi_{V}(v)$ are independent and have disjoint supports.
  4. 4.
    For each $uv\in E$, $\pi_{E}(uv)$ is dependent only on $\pi_{V}(u)$ and $\pi_{V}(u)$.
  5. 5.
    Each path $\pi_{E}(e)$ (for $e\in E$) has length at most $\operatorname{\mathcal{O}}_{\Delta}\left(2^{k}\right)$.
  6. 6.
    Each vertex in $H$ appears in the image of $\pi_{V}$ with probability $\Theta_{\Delta}(1/4^{k})$. That is, $v^{\prime}\in V^{\prime}$ appears in the image of $\pi_{V}(p(v^{\prime}))$ with this probability.
  7. 7.
    Each edge in $H$ appears in the image of $\pi_{E}$ with probability $\operatorname{\mathcal{O}}_{\Delta}(1/2^{k})$, and moreover each edge (and hence also vertex endpoint) in $H$ appears in the support of $\operatorname{\mathcal{O}}_{\Delta}(1)$ random variables $\pi_{E}(e)$ for $e\in E$. Denote the support of edges in $\pi_{E}(e)$ as $p(e)$.
  8. 8.
    Each vertex $v^{\prime}\in V^{\prime}$ appears as an endpoint in the support of some $\pi_{E}(uv)$ only if it is contained in $p(u)\cup p(v)$ (i.e., the vertices incident to elements of $p(uv)$ are a subset of $p(u)\cup p(v)$).
Essentially, this construction is similar to our earlier uniform shallow minors, except that the uniformity is approximate and ignores factors of $\Delta$. Additionally, this construction also uses randomized paths that use only a small number of vertices in expectation. In contrast, our earlier construction for uniform shallow minors used an upper bound that accounted for every vertex in every path.
Using these properties, we obtain that for each pair $u,v\in V$,
$\operatorname{\mathbb{E}}\left[|f_{H}(\pi_{V}(u))-f_{H}(\pi_{V}(v))|^{2}\right]\in\Omega_{\Delta}\left(\frac{1}{16^{k}}\sum_{u^{\prime}\in p(u),v^{\prime}\in p(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}\right),$
and if $uv\in E$ then
$\operatorname{\mathbb{E}}\left[\sum_{u^{\prime}v^{\prime}\in\pi_{E}(uv)}y_{H}(u^{\prime})+y_{H}(v^{\prime})\right]\in\operatorname{\mathcal{O}}_{\Delta}\left(\frac{1}{2^{k}}\sum_{u^{\prime}v^{\prime}\in p(uv)}y_{H}(u^{\prime})+y_{H}(v^{\prime})\right).$
Hence, there exists some deterministic choices $\pi_{V}^{*},\pi_{E}^{*}$ so that
$\frac{\sum_{uv\in E}\sum_{u^{\prime}v^{\prime}\in\pi_{E}^{*}(uv)}y_{H}(u^{\prime})+y_{H}(v^{\prime})}{\sum_{u,v\in V}|f_{H}(\pi_{V}^{*}(u))-f_{H}(\pi_{V}^{*}(v))|^{2}}\leq\frac{\operatorname{\mathbb{E}}\left[\sum_{uv\in E}\sum_{u^{\prime}v^{\prime}\in\pi_{E}(uv)}y_{H}(u^{\prime})+y_{H}(v^{\prime})\right]}{\operatorname{\mathbb{E}}\left[\sum_{u,v\in V}|f_{H}(\pi_{V}(u))-f_{H}(\pi_{V}(v))|^{2}\right]}$
$\in\operatorname{\mathcal{O}}_{\Delta}\left(8^{k}\frac{\sum_{uv\in E}\sum_{u^{\prime},v^{\prime}\in p(uv)}y_{H}(u^{\prime})+y_{H}(v^{\prime})}{\sum_{u,v\in V}\sum_{u^{\prime}\in p(u),v^{\prime}\in p(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}}\right)$
$=\operatorname{\mathcal{O}}_{\Delta}\left(8^{k}\frac{\sum_{v^{\prime}\in V^{\prime}}y_{H}(v^{\prime})}{\sum_{u,v\in V}\sum_{u^{\prime}\in p(u),v^{\prime}\in p(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}}\right).$
Let $\rho:=\max_{uv\in E}|\pi_{E}^{*}(uv)|$ be the maximum length (in terms of edges) of a sampled path, so $\rho\in\operatorname{\mathcal{O}}_{\Delta}(2^{k})$. Choose $f_{G}(v):=f_{H}(\pi_{V}^{*}(v))-\frac{1}{n}\sum_{x\in V}f_{H}(\pi_{V}^{*}(x))$, and $y_{G}(v):=2\rho\frac{\sum_{u:uv\in E}\sum_{x\in p(v)\cap\pi_{E}^{*}(uv)}y_{H}(v)}{\sum_{x\in V}|f_{G}(x)|^{2}}$.
We start by showing that the constraints are satisfied. First, note that $\sum_{v\in V}f_{G}(v)=\sum_{v\in V}f_{H}(v)-\sum_{v\in V}f_{H}(v)=\overline{0}$, so the normalization constraint is satisfied. Next, consider some edge $uv\in E$. By Cauchy-Schwarz and the triangle inequality (similar to the argument in Lemma 3.11),
$y_{G}(u)+y_{G}(v)\geq\frac{|f_{G}(u)-f_{G}(v)|^{2}}{\sum_{x\in V}|f_{G}(x)|^{2}},$
so the remaining constraints are satisfied.
It remains only to check the objective value. First, note that
$2n\sum_{x\in V}|f_{G}(x)|^{2}=\sum_{u,v\in V}|f_{G}(u)-f_{G}(v)|^{2}=\sum_{u,v\in V}|f_{H}(\pi_{V}^{*}(u))-f_{H}(\pi_{V}^{*}(v))|^{2}.$
Moreover,
$\sum_{u,v\in V}\sum_{u^{\prime}\in p(u),v^{\prime}\in p(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}$
$=\sum_{u^{\prime},v^{\prime}\in V^{\prime}}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}-\sum_{v\in V}\sum_{u^{\prime},v^{\prime}\in p(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}.$
Next,
$\sum_{v\in V}\sum_{u^{\prime},v^{\prime}\in p(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}\leq\sum_{v\in V}\sum_{u^{\prime},v^{\prime}\in p(v)}y_{H}(u^{\prime})+y_{H}(v^{\prime})$
$\leq\sum_{v\in V}2|p(v)|\sum_{v^{\prime}\in p(v)}y_{H}(v^{\prime})\in\operatorname{\mathcal{O}}_{\Delta}(4^{k}\cdot\sum_{v^{\prime}\in V^{\prime}}y_{H}(v^{\prime}))=\operatorname{\mathcal{O}}_{\Delta}(4^{k}\cdot\gamma^{(1)}(H))$
We may assume without loss of generality that $\operatorname{\mathcal{O}}_{\Delta}(4^{k}\cdot\gamma^{(1)}(H))\leq 1$, since otherwise the lemma statement follows from the fact that $\gamma^{(1)}(G)\leq 1$. Hence, $\sum_{v\in V}\sum_{u^{\prime},v^{\prime}\in p(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}=\Theta_{\Delta}(\sum_{u^{\prime},v^{\prime}\in V^{\prime}}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2})$. Combining our bounds together, we obtain
$\sum_{v\in V}y_{G}(v)=\sum_{v\in V}2\rho\frac{\sum_{u:uv\in E}\sum_{x\in p(v)\cap\pi_{E}^{*}(uv)}y_{H}(v)}{\sum_{x\in V}|f_{G}(x)|^{2}}$
$=4\rho n\frac{\sum_{v\in V}\sum_{u:uv\in E}\sum_{x\in p(v)\cap\pi_{E}^{*}(uv)}y_{H}(v)}{\sum_{u,v\in V}|f_{G}(u)-f_{G}(v)|^{2}}\in\operatorname{\mathcal{O}}_{\Delta}\left(8^{k}\rho n\frac{\sum_{v^{\prime}\in V}y_{H}(v^{\prime})}{\sum_{u,v\in V}|f_{H}(\pi_{V}^{*}(u))-f_{H}(\pi_{V}^{*}(v))|^{2}}\right)$
$=\operatorname{\mathcal{O}}_{\Delta}\left(8^{k}\rho n\frac{\sum_{v^{\prime}\in V}y_{H}(v^{\prime})}{\sum_{u^{\prime},v^{\prime}\in V^{\prime}}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}}\right)=\operatorname{\mathcal{O}}_{\Delta}\left(8^{k}\rho\frac{n}{n^{\prime}}\sum_{v^{\prime}\in V}y_{H}(v^{\prime})\right).$
Using the bounds we have on $\rho$ and $\frac{n}{n^{\prime}}$, we obtain that $\sum_{v\in V}y_{G}(v)\in\Theta_{\Delta}\left(4^{k}\gamma^{(1)}(H)\right)$, as desired. ∎