### v4

# Reweighted Spectral Partitioning Works: A Simple Algorithm for Vertex Separators in Special Graph Classes

Abstract
We establish that a simple polynomial-time algorithm that we call reweighted spectral partitioning obtains small $\frac{2}{3}$-balanced vertex-separators for a number of graph classes, including $\operatorname{\mathcal{O}}(\sqrt{n})$-sized separators for planar graphs, $\operatorname{\mathcal{O}}(\min\{(\log g)^{2},\log\Delta\}\cdot\sqrt{gn})$-sized separators for genus-$g$ graphs of maximum degree $\Delta$, and $\operatorname{\mathcal{O}}(\min\{\log h,\sqrt{\log\Delta}\}(h\log h\log\log h)\sqrt{n})$-sized separators for $K_{h}$-minor-free graphs of maximum degree $\Delta$.
To accomplish this, we first obtain a refined form of a Cheeger-style inequality relating the vertex expansion of a graph and the solution to a semidefinite program defined over the graph. Then, to obtain the guarantees for specific graph classes, we derive direct bounds on the value of the semidefinite program.
We also obtain several other results of independent interest, including an improved separator theorem for the intersection graphs of $d$-dimensional balls with bounded ply, a new bound on the Fiedler value of genus-$g$ graphs, and a new “spectral” proof of the planar separator theorem.
1 Introduction
In this work, we study simple algorithms for computing small balanced vertex separators.
Let $G=(V,E)$ be a graph with $n$ vertices and $m$ edges. For a parameter $\alpha\in(\frac{1}{2},1)$, a subset of the vertices $S\subset V$ is called an $\alpha$-balanced vertex-separator of $G$ if every connected component of the induced subgraph of $G$ on $V\setminus S$ has at most $\alpha\cdot n$ vertices. An $\alpha$-balanced edge-separator of $G$ is a subset of edges meeting the same criteria. For simplicity, we will sometimes use the term separator to refer to a $\frac{2}{3}$-balanced vertex-separator.
Small separators are useful for a vast number of applications, including divide and conquer algorithms and dynamic programming [RH01]. Moreover, small separators are known to exist for a large number of common graph classes (see Appendix A for definitions of various graph classes). Most notably, the planar separator theorem states that a planar graph with $n$ vertices has a separator consisting of $\operatorname{\mathcal{O}}(\sqrt{n})$ vertices [Ung51, LT79]. Such a separator can also be found in linear time. Similar existential results are also known for other graph classes, such as (oriented) genus-$g$ graphs always having separators consisting of $\operatorname{\mathcal{O}}(\sqrt{g\,n})$ vertices [GHT84], and $K_{h}$-minor-free graphs always having separators consisting of $\operatorname{\mathcal{O}}(h\sqrt{n})$ vertices [KR10]. Both of these example bounds are also tight. Unfortunately, while all of these existential results also imply algorithms, these algorithms also require time exponential in the class parameter ($g$ or $h$).
Since the implicit algorithms for producing these separators are slow, it is natural to consider more general approximation algorithms. However, approximating the minimum-size $\alpha$-balanced vertex-separator of a graph is NP-hard even for graphs of maximum degree $3$, and even up to an $\operatorname{\mathcal{O}}(n^{\frac{1}{2}-\epsilon})$ additive approximation [BJ92] (for any $\epsilon>0$). Fortunately, pseudo-approximations that relax the balance $\alpha$ are just as useful for most applications: For a graph admitting a $2/3$-balanced vertex-separator of size $k$, an algorithm of Feige, Hajiaghayi, and Lee [FHL05, FHL08] can be used to produce a $3/4$-balanced separator of size $\operatorname{\mathcal{O}}(k\cdot\sqrt{\log n})$ in polynomial time. In fact, they can obtain a separator of size $\operatorname{\mathcal{O}}(k\cdot\sqrt{\log k})$.
We study a slightly different approach. Rather than aim for general (pseudo-)approximation algorithms or specialized per-class algorithms, we will study a polynomial-time algorithm for producing $\frac{2}{3}$-balanced vertex-separators in an arbitrary graph. We call this algorithm reweighted spectral partitioning due to its numerous similarities with the popular “spectral-partitioning” algorithm [ST96, ST07]. This algorithm is sufficiently simple that we have implemented it, [^1] We have not implemented an optional partition oracle used by the algorithm. and examples of separators that it produces can be found in Figure 1.
Figure 1: Examples of vertex separators produced by reweighted spectral partitioning: A $100$-vertex grid graph (left), a $50$-vertex random planar graph (centre), and $50$-vertex 22-dimensional 55-nearest-neighbour graph (right). The separator vertices are the large red ones.
Compared to the pseudo-approximations mentioned above, reweighted spectral partitioning has relatively weak guarantees for general graphs. However, the primary goal of this paper is to show that it admits quite strong guarantees for a number of specific graph classes, including planar graphs and bounded-genus graphs. In fact, we will show that the separators it produces nearly match the per-class separator sizes for a number of important graph classes.
1.1 The Method of Reweighted Spectral Partitioning
The reweighted spectral partitioning algorithm that we consider is a rounding algorithm for a semidefinite program. In particular, it is the combination of algorithms implicit in the proofs of Kwok, Lau, and Tung [KLT22] and Biswal, Lee, and Rao [BLR10]. Critically, it makes use of the following optimization problem parametrized by a value $d\in\mathbb{N}$:
[…]
Several previous works have studied this problem for $d=n$ or $d=1$ [BDX04, Roc05, OTZ22, JPV22, KLT22]. In particular, when $d=n$, this is a semidefinite program and it is known to satisfy Slater’s condition, so it can be approximately solved in polynomial time with general methods [JKL+20].
The reweighted spectral partitioning algorithm we consider then consists of four high-level steps for a graph $G$:
  1. 1.
    Approximately solve the semidefinite program $\gamma^{(n)}(G)$, obtaining variables $f(v)\in\mathbb{R}^{n}$ and $y(v)\in\mathbb{R}_{\geq 0}$ for each vertex $v\in V$.
  2. 2.
    Dimension-reduce the $f(v)$ variables to get one-dimensional values $f^{\prime}(v)\in\mathbb{R}$. Several methods/samples for dimension-reduction are applied, and the best result is used. One of the methods takes, as a parameter, an optional oracle that partitions the vertices of the graph based on a vertex-weighting.
  3. 3.
    Sort the vertices by their $f^{\prime}(v)$ values to get an ordering $v_{1},\dots,v_{n}$, and consider partitions of the form $A_{k}=\{v_{1},\dots,v_{k}\},B_{k}=\{v_{k+1},\dots,v_{n}\}$. For each crossing edge in $A_{k}\times B_{k}\cap E$, select its incident vertex $v$ maximizing $y(v)$, and let $S_{k}$ be the set of selected vertices. Return $(A,S,B):=(A_{k},S_{k},B_{k})$ minimizing $\frac{|S_{k}|}{\min\{|A_{k}|,|B_{k}|\}}$.
  4. 4.
    Repeat on the subgraph of $G$ induced by larger of the sets $A$ and $B$ until the union of all the computed sets $S$ forms a balanced separator.
Importantly, the algorithm requires no information about the graph other than its vertices and edges. The optional partitioning oracle could make use of such information (if desired), but the useful partitioning oracles we will suggest do not. This makes the algorithm suitable for very general applications.
1.2 Main Results
We will show that the reweighted spectral partitioning algorithm expressed in the previous subsection produces provably good results in a number of different senses. Some of these results will make use of a specialized random partition oracle as a subroutine, while others will not.
We will use the notation $a\lesssim b$ to mean $a=\operatorname{\mathcal{O}}(b)$.
1.2.1 Refined Cheeger-Style Inequality
The first main result is an inequality relating $\gamma^{(n)}$ and $\gamma^{(1)}$ with several important quantities that we will now define. The most important of these is a quantity related to vertex separators:
Definition 1.1.
For a subset $A\subset V$ with $|A|\leq n/2$, let $\partial A$ denote the set of vertices in $V\setminus A$ adjacent to some vertex in $A$. The vertex expansion of a set $A$ is the ratio $\psi(A):=\frac{|\partial A|}{|A|}$. The vertex expansion of a graph $G$ is the minimum vertex expansion over all possible sets $|A|\leq\frac{n}{2}$, and it is denoted as $\psi(G)$.
Computing vertex expansion is known to be NP-hard (see Appendix B for details), so it is usually studied alongside approximation algorithms. Vertex expansion is intimately related to balanced vertex-separators. Notably, approximations of vertex-expansion imply pseudo-approximations of balanced vertex-separators [FHL05, FHL08, LR99]. In our case, a simpler fact is more relevant: Given a graph $G$ with $n$ vertices, if every induced subgraph $H$ of $G$ on $n^{\prime}$ vertices has vertex expansion at most $\frac{\kappa}{\sqrt{n^{\prime}}}$, then $G$ admits a $\frac{2}{3}$-balanced vertex separator of size $2\kappa\sqrt{n}$. This separator can be obtained by repeatedly finding sets $A$ of vertices with small vertex expansion, and then deleting $\partial(A)$ and $A$ from the graph. The final separator is the union of all sets $\partial A$ from this process. This reduction is standard for a number of different formulations of expansion and separators (e.g., see [ST07, Lemma A.1]).
The second quantity relates to our optional partition oracle used for dimension-reduction. Intuitively, the following quantity arises when we aim to randomly partition a metric while trying to avoid cutting large metric balls (in expectation):
Definition 1.2.
Let $(X,d)$ be a metric space. The (weak) modulus of padded decomposability of $(X,d)$ is defined as the smallest value of $\alpha$ so that for every $\Delta>0$, there exists a distribution $\mu$ of partitions of $X$ with parts of diameter at most $\Delta$, and
$\begin{array}[]{c c c c}\sup_{x\in X}&\Pr_{P\sim\mu}\left[B\left(x,\frac{\Delta}{\alpha}\right)\subset P(x)\right]&\geq&\frac{1}{2},\end{array}$
where $B(x,r)$ denotes the metric ball of radius $r$ centred at $x$. We denote this quantity as $\alpha(X,d)$.
Definition 1.3.
For a graph $G=(V,E)$, the extremal vertex modulus of padded decomposability on $G$ is defined as the maximum value of $\alpha(V,d_{\omega})$ for vertex-weight functions $\omega:V\to\mathbb{R}_{\geq 0}$ inducing shortest-path metrics $d_{\omega}$ on $G$. We denote this quantity as $\alpha(G)$.
A similar definition exists for edge-weighted shortest-path metrics, but we will only use this one. $\alpha(G)$ is known to be bounded for a number of graph classes. In particular, for $K_{h}$-minor-free graphs, $\alpha(G)\in\operatorname{\mathcal{O}}(\log h)$ [CF25]. Note that genus-$g$ graphs are $K_{\Theta(\sqrt{g})}$-minor-free.
Finally, the result we will prove is:
Theorem 1.4 (Refined Cheeger-Style Inequality for Vertex Expansion).
For a graph $G$ with $n$ vertices and maximum degree $\Delta$,
$\frac{\psi(G)^{2}}{\min\{\log\Delta,[\alpha(G)]^{2}\}}\lesssim\frac{\gamma^{(1)}(G)}{\min\{\log\Delta,[\alpha(G)]^{2}\}}\lesssim\gamma^{(n)}(G)\lesssim\gamma^{(1)}(G)\lesssim\psi(G).$
Moreover, given an oracle that produces an $\alpha$-padded partition for any vertex-weighting of $G$ in polynomial time, the reweighted spectral partitioning algorithm can compute in (Monte Carlo randomized) polynomial time a set $S$ with $|S|\leq\frac{n}{2}$ and vertex expansion $\psi(S)\lesssim\sqrt{\gamma^{(n)}(G)\cdot\min\{\log\Delta,\alpha^{2}\}}$.
This is referred to as a “Cheeger-style” inequality since the upper bounds on $\psi(G)$ are given in terms of a square root of an expression involving either $\gamma^{(1)}(G)$ or $\gamma^{(n)}(G)$, while the lower bounds on $\psi(G)$ have no such square root operation. This is similar to the well-known Cheeger inequality, which relates some similar quantities [HLW06, Theorem 4.11].
The proof of this inequality will directly use the algorithm of reweighted spectral partitioning (more specifically, Algorithm 5 from Section 3). In fact, this result is a refinement of a result from a long sequence of Cheeger-style inequalities for vertex-expansion [Roc05, OTZ22, JPV22, KLT22]. In particular, the specific algorithm we use is a refinement of the one implicit in the proofs of Kwok, Lau, and Tung [KLT22], incorporating a new dimension-reduction step based on a technique of Biswal, Lee, and Rao [BLR10] that uses random partitions. Their algorithm is not the only one that could form the basis of ours, but it has the benefit of being quite simple.
1.2.2 Vertex Separators for Special Graph Classes
The bulk of this work is dedicated to showing that reweighted spectral partitoning can be used to recover balanced vertex separators for a number of special graph classes that have nearly-optimal size for that class. All of these proofs will make use of part of the Cheeger-style inequality: By directly bounding $\gamma^{(d)}$ for a particular class (and some dimension $d$), we will obtain guarantees for the performance of reweighted spectral partitioning on that class by applying Theorem 1.4. We present several such results.
Theorem 1.5 (Bounds on $\gamma^{(d)}$).
Let $G$ be a graph with $n$ vertices and maximum degree $\Delta$. Then the following conditional bounds hold:
  • If $G$ is planar, then $\gamma^{(1)}(G)\lesssim\frac{1}{n}$.
  • If $G$ has (oriented) genus $g$, then $\gamma^{(1)}(G)\lesssim\frac{g\cdot\min\{(\log g)^{2},\log\Delta\}}{n}$.
  • If $G$ is $K_{h}$-minor-free, then $\gamma^{(1)}(G)\lesssim\frac{(h\log h\log\log h)^{2}}{n}$.
  • If $G$ is a $d$-dimensional $k$-ply ball-intersection graph [^2] Sometimes called a neighbourhood-system. in $\mathbb{R}^{d}$, then $\gamma^{(d+1)}(G)\lesssim\left(\frac{k}{n}\right)^{\frac{2}{d}}$.
  • If $G$ is a $d$-dimensional $k$-nearest-neighbour graph in $\mathbb{R}^{d}$, then $\gamma^{(d+1)}(G)\lesssim\left(\frac{k}{n}\right)^{\frac{2}{d}}$.
All of these classes are hereditary (closed under induced subgraphs), so these bounds can be combined with Theorem 1.4 in several different ways to obtain results for balanced vertex-separators. The first is a purely existential construction that uses the bound $\psi(G)^{2}\lesssim\gamma^{(d)}(G)\cdot\min\{\log\Delta,[\alpha(G)]^{2}\}$:
Corollary 1.6 (Existential Separator Bounds).
Let $G$ be a graph with $n$ vertices and maximum degree $\Delta$. Then the following conditional bounds hold:
  • If $G$ is planar, then $G$ admits a $\frac{2}{3}$-balanced vertex-separator of size $\operatorname{\mathcal{O}}(\sqrt{n})$.
  • If $G$ has (oriented) genus $g$, then $G$ admits a $\frac{2}{3}$-balanced vertex-separator of size
    $\operatorname{\mathcal{O}}(\min\{\log g,\sqrt{\log\Delta}\}\sqrt{gn})$.
  • If $G$ is $K_{h}$-minor-free, then $G$ admits a $\frac{2}{3}$-balanced vertex-separator of size $\operatorname{\mathcal{O}}((h\log h\log\log h)\sqrt{n})$.
  • If $G$ is a $k$-ply ball-intersection graph in $\mathbb{R}^{d}$, then $G$ admits a $\frac{2}{3}$-balanced vertex-separator of size $\operatorname{\mathcal{O}}(\sqrt{\min\{d,\log\Delta\}}\cdot k^{1/d}n^{1-1/d})$.
  • If $G$ is a $k$-nearest-neighbour graph in $\mathbb{R}^{d}$, then $G$ admits a $\frac{2}{3}$-balanced vertex-separator of size $\operatorname{\mathcal{O}}(\sqrt{\min\{d,\log\Delta\}}\cdot k^{1/d}n^{1-1/d})$.
Take note of the fact that these bounds do not result from running reweighted spectral partitioning, and are not inherently constructive. Moreover, the first three of these five bounds are not new, and the bounds for genus-$g$ and $K_{h}$-minor-free graphs are also weaker than the known existential bounds of $\operatorname{\mathcal{O}}(\sqrt{gn})$ [GHT84] and $\operatorname{\mathcal{O}}(h\sqrt{n})$ [KR10], respectively. However, the final two of these bounds for the geometric classes are brand new separator bounds, improving over the previously-known bound of $\operatorname{\mathcal{O}}(dk^{1/d}n^{1-1/d})$ [MTTV97]. While separators of this size would not result from reweighted spectral partitioning, we will discuss their algorithmic aspects in Section 2.3.2.
The second set of bounds we obtain for separators are those that are achieved by reweighted spectral partitioning when supplied with the polynomial-time random padded partition oracle of Conroy and Filtser [CF25]:
Corollary 1.7 (Separator Bounds via Reweighted Spectral Partitioning with Padded Partitions).
Let $G$ be a graph with $n$ vertices and maximum degree $\Delta$. When provided with a known polynomial-time algorithm to (randomly) partition the vertex weights, reweighted spectral partitioning can be used to compute the following sizes of separators in (Monte Carlo randomized) polynomial time:
  • If $G$ is planar, then it produces a $\frac{2}{3}$-balanced vertex-separator of size $\operatorname{\mathcal{O}}(\sqrt{n})$.
  • If $G$ has (oriented) genus $g$, then it produces a $\frac{2}{3}$-balanced vertex-separator of size
    $\operatorname{\mathcal{O}}(\min\{(\log g)^{2},\log\Delta\}\sqrt{gn})$.
  • If $G$ is $K_{h}$-minor-free, then it produces a $\frac{2}{3}$-balanced vertex-separator of size
    $\operatorname{\mathcal{O}}(\min\{\log h,\sqrt{\log\Delta}\}(h\log h\log\log h)\sqrt{n})$.
The final set of bounds we obtain for separators are those that are achieved by reweighted spectral partitioning when no partitioning oracle is provided:
Corollary 1.8 (Separator Bounds via Reweighted Spectral Partitioning with No Partition Oracle).
Let $G$ be a graph with $n$ vertices and maximum degree $\Delta$. Then reweighted spectral partitioning can be used to compute the following sizes of separators in (Monte Carlo randomized) polynomial time:
  • If $G$ is planar, then it produces a $\frac{2}{3}$-balanced vertex-separator of size $\operatorname{\mathcal{O}}(\sqrt{n\log\Delta})$.
  • If $G$ has (oriented) genus $g$, then it produces a $\frac{2}{3}$-balanced vertex-separator of size
    $\operatorname{\mathcal{O}}(\min\{\log g\cdot\sqrt{\log\Delta},\log\Delta\}\cdot\sqrt{gn})$.
  • If $G$ is $K_{h}$-minor-free, then it produces a $\frac{2}{3}$-balanced vertex-separator of size
    $\operatorname{\mathcal{O}}(\sqrt{\log\Delta}\cdot(h\log h\log\log h)\cdot\sqrt{n})$.
  • If $G$ is a $k$-ply ball-intersection graph in $\mathbb{R}^{d}$, then it produces a $\frac{2}{3}$-balanced vertex-separator of size $\operatorname{\mathcal{O}}(\sqrt{\log\Delta}\cdot k^{1/d}n^{1-1/d})$.
  • If $G$ is a $k$-nearest-neighbour graph in $\mathbb{R}^{d}$, then it produces a $\frac{2}{3}$-balanced vertex-separator of size $\operatorname{\mathcal{O}}(\sqrt{\log\Delta}\cdot k^{1/d}n^{1-1/d})\leq\operatorname{\mathcal{O}}(\sqrt{d+\log k}\cdot k^{1/d}n^{1-1/d})$.
Not providing an oracle has the benefit of keeping the algorithm incredibly simple (see Section 3), but the guarantees on the size of the resulting separators are weaker.
1.2.3 Comparison to Other Algorithms
The main benefit of reweighted spectral partitioning is that it is a simple polynomial-time algorithm that provides near-optimal separators for a number of graph classes, and requires nothing but the graph itself as input. However, despite its simplicity, it remains comparable with other known polynomial-time algorithms that do not require anything but the graph itself, and in some cases it produces improved results. We provide comparisons to other work in Table 1.
Table 1: Comparison of our performance guarantees for reweighted spectral partitioning with existing (Monte Carlo randomized) polynomial-time methods. The best bound is always bolded (multiple are bolded if the best bound depends on parameters). Note that one of the previous works’ bounds for genus-$g$ graphs requires a fix given in Section 5.3.
2 Technical Overview, Related Work, and Additional Results
In this section, we will briefly outline some related work in spectral graph theory, and then discuss our techniques for obtaining bounds. In particular, we will use two different families of techniques for obtaining bounds, and we will discuss each one separately.
2.1 Fiedler Bounds and the Cheeger Inequality
For an undirected graph $G$ with $n$ vertices, let its adjacency matrix (or weight matrix) be denoted $A$, and the diagonal matrix formed from its degrees (or sums of incident weights) be denoted $D$. The Laplacian matrix of $G$ is L(G):=DAL(G):=D-A, and its eigenvalues are denoted $0=\lambda_{1}(G)\leq\lambda_{2}(G)\leq\lambda_{3}(G)\leq\cdots\leq\lambda_{n}(G)$. The second-smallest Laplacian eigenvalue ($\lambda_{2}(G)$) was named the algebraic connectivity of a graph by Fiedler [Fie73], and is sometimes called the Fiedler value of a graph. As the name would suggest, algebraic connectivity is related to a combinatorial definition of connectivity:
For a graph $G=(V,E)$, the edge expansion of a set $S\subset V$ with $|S|\leq\frac{n}{2}$ is defined as $\phi(S):=\frac{|\{uv\in E\cap S\times(V\setminus S)\}|}{|S|}$, and the edge expansion of the graph $G$ is defined as $\phi(G):=\min_{|S|\leq\frac{n}{2}}\phi(S)$.
The edge expansion $\phi(G)$ and the Fiedler value $\lambda_{2}(G)$ are related by the following important result:
Theorem 2.1 (Cheeger inequality for edge expansion [AM85, Moh89, JS88]).
For a graph $G$ with maximum degree $\Delta$,
$\frac{\phi(G)^{2}}{2\Delta}\leq\lambda_{2}(G)\leq 2\phi(G).$
This inequality is very similar to Theorem 1.4 in a number of ways. In particular, there is a simple near-linear-time algorithm, called spectral partitioning, that can produce a set $S$ with $\phi(S)\lesssim\sqrt{\Delta\lambda_{2}(G)}$ [KMP11]. Very similarly to our results, one can obtain direct bounds on $\lambda_{2}(G)$ for several special graph classes (see Table 2). Consequently, this algorithm can be used to produce good cuts, and even near-optimal balanced vertex-separators if the maximum degree is constant.
Table 2: Best-known asymptotic upper bounds on $\lambda_{2}(G)$ for several graph classes, for $n$-vertex graphs of maximum degree $\Delta$.
2.2 The Fastest-Mixing Markov Chain Problem and the Cheeger-Style Inequality for Vertex Expansion
Since the Fiedler value $\lambda_{2}(G)$ is associated with higher connectivity, it is natural to ask if different weights along the edges of the graph can result in higher values of $\lambda_{2}(G)$. Without any constraints on the weights, the value of $\lambda_{2}(G)$ is unbounded. A natural normalization constraint is to require that the total weight around each vertex is at most $1$. Boyd, Diaconis and Xiao studied exactly this problem, which they called the fastest-mixing (reversible) Markov chain problem [BDX04]:
$\begin{array}[]{r c c r c l l}\lambda_{2}^{*}(G)&:=&\max_{P\geq 0}&\lambda_{2}(I-P)\\ &&\textup{subject to}&P(u,v)&=&0&\forall uv\not\in E\cup\{vv:v\in V(G)\}\\ &&&\sum_{v\in V}P(u,v)&=&1&\forall u\in V\\ &&&P(u,v)&=&P(v,u)&\forall uv\in E\cup\{vv:v\in V(G)\}\end{array}$
Here $P$ is a reweighted adjacency matrix of $G$ (with self-loops) constrained to produce a reversible Markov chain with a uniform stationary distribution. Hence, $I$ is the degree-matrix for the weighted graph, and IPI-P is the Laplacian matrix, so $\lambda_{2}(I-P)$ is the Fiedler value.
33footnotetext: Due to some subtle issues, this bound does not apply to general genus-$g$ graphs, which is what the references claim. We both discuss and amend this issue in Section 5.3.
There are two simple and important relationships relating $\lambda_{2}(G)$, $\lambda_{2}^{*}(G)$, and $\gamma^{(n)}(G)$: First, it can be shown that $\lambda_{2}^{*}(G)=\gamma^{(n)}(G)$ by a duality argument [BDX04, Roc05]. Second, $\lambda_{2}(G)\leq\Delta\cdot\lambda_{2}^{*}(G)$. Therefore, Theorem 1.5 actually generalizes the known bounds on $\lambda_{2}(G)$. In fact, one of the resulting bounds is new:
Corollary 2.2.
If $G$ is a genus-$g$ graph, then its Fiedler value satisfies $\lambda_{2}(G)\lesssim\frac{g\cdot\Delta\log\Delta}{n}$.
We will henceforth use $\gamma^{(n)}(G)$ instead of $\lambda_{2}^{*}(G)$ since they are equal.
There has been significant prior work towards relating $\gamma^{(n)}$ to vertex-expansion. In particular, our Theorem 1.4 is a refinement of the following result:
Theorem 2.3 (Cheeger inequality for vertex expansion [Roc05, OTZ22, JPV22, KLT22]).
For an undirected graph $G$ with maximum degree $\Delta$
$\frac{\psi(G)^{2}}{\log\Delta}\lesssim\gamma^{(n)}(G)\lesssim\psi(G).$
This line of work is what gives rise to the name reweighted spectral partitioning. In particular, the most recent improvements to this result were in the dimension-reduction step of the algorithm, improving a $\log n$ factor to the current $\log\Delta$ factor [JPV22, KLT22]. Our result further improves this dimension-reduction step to get a $\min\{\alpha(G)^{2},\log\Delta\}$ factor.
2.3 Bounds via Geometry with Ball-Intersection Graphs
Our first family of methods for obtaining upper bounds makes use of ball-intersection graphs. These are a very natural object to consider: For any feasible solution to $\gamma^{(d)}(G)$, we will show that $f,y$ induce a ball-intersection graph $H=(V,F)$ with centre $f(v)$ and radius $2\sqrt{y(v)}$ for each vertex $v\in V$. In fact, we will reformulate $\gamma^{(d)}(G)$ into a new form that more directly relates to such a representation, at only a constant-factor loss.
We will use this form to give upper bounds on $\gamma^{(d)}(G)$ for planar graphs, genus-$g$ graphs, ball-intersection graphs of bounded ply, and $k$-nearest-neighbour graphs. Specifically, we will obtain the following bounds:
Proposition 2.4.
Let $G$ be a planar graph with $n$ vertices and maximum degree $\Delta$. Then,
$\gamma^{(1)}(G)\lesssim\gamma^{(3)}(G)\leq\frac{8}{n}.$
Proposition 2.5.
Let $G$ be the intersection graph of $d$-dimensional balls with ply $k$ and maximum degree $\Delta$. Then,
$\gamma^{(d+1)}(G)\lesssim\left(\frac{k}{n}\right)^{\frac{2}{d}}.$
Corollary 2.6.
Let $G$ be a $d$-dimensional $k$-nearest neighbour graph with $n$ vertices. Then,
$\gamma^{(d+1)}(G)\lesssim\left(\frac{k}{n}\right)^{\frac{2}{d}}.$
Proposition 2.7.
Let $G$ be a graph with $n$ vertices, maximum degree $\Delta$, and genus at most $g$. Then,
$\gamma^{(1)}(G)\lesssim\frac{g\log\Delta}{n}.$
Note that Proposition 2.4 implies the planar separator theorem when applied to Theorem 1.4, while Propositions 2.5 and 2.6 both imply the new separator theorems for their respective classes mentioned in Corollary 1.6.
Many of the techniques for obtaining these bounds will be similar to existing techniques for bounding the Fiedler value $\lambda_{2}(G)$ of graphs in these classes [ST96, ST07, Kel04, Kel06b, Kel06a]. At a high-level, the method is always to map a set of $d$-dimensional balls representing the graph to a $d$-dimensional (hyper)sphere (that is, the (hyper)surface of a $(d+1)$-dimensional ball) so that the measure of the surface of the (hyper)sphere can be related to the average ply of the balls. In most cases, it is easier to bound the maximum ply of the balls, but genus-$g$ graphs require a more robust approach.
2.3.1 New techniques for bounded-genus graphs
While most of these bounds can be obtained primarily by adapting known techniques, we require some very significant novel constructions to prove Proposition 2.7.
First, we show that Kelner’s method for triangulated genus-$g$ graphs that obtains a bound of $\lambda_{2}(G)\lesssim\frac{\text{poly}(\Delta)\cdot g}{n}$ can also be extended to bound $\gamma^{(1)}(G)$. Kelner’s method is quite involved, so there are many aspects that have to be re-checked in the context of our modified problem, but this step uses similar methods overall.
However, this is not enough to obtain our bound of $\gamma^{(n)}(G)\lesssim\frac{g\cdot\log\Delta}{n}$, which applies not just to triangulated genus-$g$ graphs but also to general genus-$g$ graphs. To obtain this bound, we provide a pair of results:
Lemma 2.8.
Let $G$ be a graph with genus $g$, $n$ vertices, and maximum degree $\Delta$. Then there exists a graph $H$ with $n\Delta$ vertices, maximum degree $4$, and genus $g$, so that $\gamma^{(1)}(G)\lesssim\gamma^{(1)}(H)\cdot\Delta\cdot\log\Delta$.
Lemma 2.9.
Let $G$ be a graph of genus $g$ with $n$ vertices and maximum degree $\Delta$. Then there is a triangulated genus-$g$ graph $H$ with $(\Delta+1)\cdot n$ vertices and maximum degree $O(\Delta)$, so that $\gamma^{(1)}(G)\lesssim(\Delta+1)\cdot\gamma^{(1)}(H)$.
Both of these results use a new structure we call uniform shallow minors, which are essentially shallow minors in which all collections of vertices are of exactly equal size, and all vertices are in some collection. One could also think of these as a very special kind of bounded-diameter decomposition. However, the way we use this structure is in some sense opposite to how both minors and bounded-diameter decompositions are typically used: In each result, we construct a graph $H$ that contains $G$ as a uniform shallow minor.
2.3.2 A new separator theorem for ball-intersection graphs of bounded ply
For the case of $d$-dimensional ball-intersection graphs of ply $k$, and $d$-dimensional $k$-nearest-neighbour graphs, the sizes of the separators guaranteed to exist by Corollary 1.6 are actually improvements over the previous bests by a factor of $\frac{\sqrt{\min\{d,\log\Delta\}}}{d}$. The previous best bound for both classes was a proof that $\operatorname{\mathcal{O}}(d)$-balanced vertex-separators exist of size $\operatorname{\mathcal{O}}(k^{1/d}n^{1-1/d})$ [MTTV97], which implies the existence of a $\frac{2}{3}$-balanced vertex separator of size $\operatorname{\mathcal{O}}(dk^{1/d}n^{1-1/d})$. While our separators can’t be constructed via reweighted spectral partitioning, they still can be constructed in polynomial time if the explicit geometric representation of the graph is given. We will obtain the following results:
Corollary 2.10.
Let $G$ be the intersection graph of $d$-dimensional balls with ply $k$ and maximum degree $\Delta$, provided as a set of $d$-dimensional coordinates and radii for each vertex. Then, in polynomial time, we can compute a balanced $2/3$-vertex-separator of size $\operatorname{\mathcal{O}}\left(\sqrt{\min\{d,\log\Delta\}}\cdot\left(\frac{k}{n}\right)^{\frac{1}{d}}\right)$.
Corollary 2.11.
Let $G$ be a $d$-dimensional $k$-nearest neighbour graph with maximum degree $\Delta$, provided as a set of $d$-dimensional coordinates and a value $k$. Then, in polynomial time, we can compute a balanced $2/3$-vertex-separator of size $\operatorname{\mathcal{O}}\left(\sqrt{\min\{d,\log\Delta\}}\cdot\left(\frac{k}{n}\right)^{\frac{1}{d}}\right)$.
2.4 Bounds via Extremal Spread
Our second family of methods for obtaining upper bounds on $\gamma^{(n)}$ and $\gamma^{(1)}$ makes use of a widely-used quantity for approximation algorithms: Extremal spread of (vertex-weighted) shortest-path metrics. Intuitively, extremal spread is a (normalized) weighting of the vertices in a graph so that the pairwise distances are maximized.
Definition 2.12.
For a graph $G$ and a non-negative vertex weighting function $\omega:V(G)\to\mathbb{R}_{\geq 0}$, let $d_{G}^{\omega}:V\times V\to\mathbb{R}_{\geq 0}$ be the vertex-weighted shortest-path semi-metric through $G$ with vertex-weights given by $\omega$ (the vertex-weighted length of a path is the sum of the weights of the vertices along it, halving the contributions of the first and last vertex). The spread of $\omega$ is
$\sum_{u,v\in V(G)}d_{G}^{\omega}(u,v).$
The $L^{p}$-extremal spread of $G$ is
[…]
$L^{1}$-extremal spread is known to be closely related to vertex expansion. However, in this work, we are exclusively interested in the case of $L^{2}$-extremal spread, which has two very important known bounds:
Proposition 2.13 ([BLR10, Theorem 3.1]).
For a genus-$g$ graph $G$ with $n$ vertices, if $n\geq 3\sqrt{g}$, then
$\overline{s}_{2}(G)\gtrsim\frac{n^{2}}{\sqrt{g}}.$
Proposition 2.14 ([BLR10, DP25]).
For a $K_{h}$-minor-free graph $G$ on $n$ vertices, if $n\gtrsim h\log\log h$, then
$\overline{s}_{2}(G)\gtrsim\frac{n^{2}}{h\log\log h}.$
The first of these theorems is optimal up to constant factors, while the second is optimal up to the $\operatorname{\mathcal{O}}(\log\log h)$ factors. These $\operatorname{\mathcal{O}}(\log\log h)$ factors come from the current state of progress towards Hadwiger’s conjecture, a major open problem in graph theory [Die25].
The key result we will obtain is as follows:
Proposition 2.15.
For a graph $G$ with maximum degree $\Delta$,
$\gamma^{(1)}(G)\lesssim\left[\alpha(G)\right]^{2}\cdot\frac{n^{3}}{\left[\overline{s}_{2}(G)\right]^{2}}.$
3 The Algorithm
We will now describe the three steps of the reweighted spectral partitioning algorithm in detail. This is only a reference for the algorithm and its time complexity/variations – the corresponding theorems are stated in the next section.
We start by presenting the simplest possible form of the algorithm, which is Monte Carlo randomized.
Dimension Reduction
We obtain a solution to $\gamma^{(1)}$ from the solution to $\gamma^{(n)}$.
Algorithm 1 ProjectionReduce($f:V\to\mathbb{R}^{n}$, $G=(V,E)$)
for $i=1$ to $\Theta(\log n)$ do
  Sample $r_{i}\sim\mathcal{N}(0,1)^{n}$
  $f^{\prime}_{i}(v)\leftarrow\langle f(v),r_{i}\rangle$ for all $v\in V$
end for
return $f^{\prime}_{i}$ maximizing $\sum_{u,v\in V}[f^{\prime}_{i}(u)-f^{\prime}_{i}(v)]^{2}$
Algorithm 2 DistanceReduce($y:V\to\mathbb{R}$, $G=(V,E)$, RandomPartition)
Compute $d(u,v)$ for all $u,v\in V$ as shortest-path distance in $G$ with vertex weights $y$
$S_{0}\leftarrow\left\{\arg\max_{v\in V}\sum_{u\in V}d(u,v)^{2}\right\}$
for $i=1$ to $\Theta(\log n)$ do
  $\mathcal{P}\leftarrow$ RandomPartition(G,y)(G,y) $\triangleright$ $\mathcal{P}$ is a partition of the vertices
  Let $S_{i}$ be the union of parts $P\in\mathcal{P}$ each sampled with probability $\frac{1}{2}$
end for
for each $S_{i}$ do
  $f^{\prime}_{i}(u)\leftarrow\min_{v\in S_{i}}d(u,v)$ for all $u\in V$
end for
return $f^{\prime}_{i}$ maximizing $\sum_{uv\in E}[f^{\prime}_{i}(u)-f^{\prime}_{i}(v)]^{2}$
Algorithm 3 BestReduce($f:V\to\mathbb{R}^{n}$, $y:V\to\mathbb{R}$, $G$, RandomPartition $\bot$)
$f^{\prime}_{1}\leftarrow$ ProjectionReduce($f$, $G$)
if RandomPartition $\neq\bot$ then $\triangleright$ If optional partition oracle is provided, use it
  $f^{\prime}_{2}\leftarrow$ DistanceReduce($y$, $G$, RandomPartition)
  return $f^{\prime}_{i}$ maximizing $\sum_{uv\in E}[f^{\prime}_{i}(u)-f^{\prime}_{i}(v)]^{2}$
else
  return $f^{\prime}_{1}$
end if
Sweep Algorithm
Extract a set of vertices from the one-dimensional embedding.
Algorithm 4 Sweep($f^{\prime}:V\to\mathbb{R}$, $y:V\to\mathbb{R}$, $G=(V,E)$)
Sort vertices by $f^{\prime}$ value: $v_{1},\dots,v_{n}$
Initialize current set of crossing edges $\mathcal{C}\leftarrow\emptyset$
for $k=1$ to n1n-1 do
  $A_{k}\leftarrow\{v_{1},\dots,v_{k}\}$, $B_{k}\leftarrow\{v_{k+1},\dots,v_{n}\}$
  Update $\mathcal{C}$ by adding edges in $v_{k}\times B_{k}\cap E$ and removing edges in $v_{k}\times A_{k}\cap E$.
  $S_{k}\leftarrow\{u:uv\in\mathcal{C},y(u)\geq y(v)\}\cup\{v:uv\in\mathcal{C},y(v)>y(u)\}$
  $\psi_{k}\leftarrow|S_{k}|/\min\{|A_{k}|,|B_{k}|\}$
end for
return the partition $(A_{k},B_{k},S_{k})$ minimizing $\psi_{k}$
The Complete Algorithms
We can use one pass of all three steps to obtain a cut with small vertex expansion, or use $\operatorname{\mathcal{O}}(n)$ passes to obtain a balanced vertex separator.
Algorithm 5 ReweightedSpectralPartitioning($G=(V,E)$, RandomPartition $\bot$)
Step 1: Solve the SDP $\gamma^{(n)}(G)$, obtaining $f:V\to\mathbb{R}^{n}$ and $y:V\to\mathbb{R}$
Step 2: $f^{\prime}\leftarrow$ BestReduce($f$, $y$, $G$, RandomPartition)
Step 3: $(A,B,S)\leftarrow$ Sweep($f^{\prime}$, $y$, $G$)
return $(A,B,S)$
Algorithm 6 BalancedRSP($G=(V,E)$, RandomPartition $\bot$)
Initialize $S_{\mathrm{union}}\leftarrow\emptyset$
$V_{\mathrm{residual}}\leftarrow V$
while $|V_{\mathrm{residual}}|>\frac{2}{3}|V|$ do
  Let $G_{\mathrm{residual}}$ be the induced subgraph of $G$ on $V_{\mathrm{residual}}$
  $(A,B,S)\leftarrow$ ReweightedSpectralPartitioning($G_{\mathrm{residual}}$, RandomPartition)
  $S_{\mathrm{union}}\leftarrow S_{\mathrm{union}}\cup S$
  $V_{\mathrm{residual}}\leftarrow\arg\max_{X\in\{A,B\}}|X|$
end while
return $S_{\mathrm{union}}$
Note that Algorithm 6 is exactly the algorithm one can apply to get the guarantees in Corollary 1.7 and Corollary 1.8.
3.1 Practical considerations
As previously mentioned, we have implemented this algorithm (specifically, Algorithm 6), although without support for partition oracles. To solve the SDP, we used the Clarabel solver [GC24]. Unfortunately, practical implementations for general SDPs are still quite slow and memory-inefficient (they often involve computing the Hessian), and this was not able to compute $\gamma^{(n)}$ for even planar graphs beyond around 150 vertices. Boyd, Diaconis, and Xiao described and tested a much more practical projected subgradient method for this problem [BDX04], but it works in the formulation of $\lambda_{2}^{*}$ rather than $\gamma^{(n)}$, and so does not natively provide the vectors $f(v)\in\mathbb{R}^{n}$ that we use for dimension-reduction and rounding. Moreover, it has no clear stopping criterion to guarantee any level of optimality, and it is not guaranteed to run in polynomial time.
4 A Refined Cheeger Inequality
In this section, we will prove Theorem 1.4 via the algorithm outlined in Section 3. Fix a graph $G=(V,E)$ with maximum degree $\Delta$. Several of the inequalities stated in Theorem 1.4 are already known:
Theorem 4.1 ([OTZ22, Theorem 2.10]).
For a graph $G$,
$\psi(G)^{2}\lesssim\gamma^{(1)}(G)\lesssim\psi(G).$
The result that $\psi(G)^{2}\lesssim\gamma^{(1)}(G)$ was re-proven by Kwok, Lau, and Tung [KLT22], who gave a slightly simpler algorithm that takes the best of a set of candidate cuts via a sweep. Our Algorithm 4 further simplifies their algorithm and considers a superset of the cuts that they consider.
Theorem 4.2 ([KLT22, Proposition 3.14]).
For a graph $G$ with maximum degree $\Delta$,
$\gamma^{(n)}(G)\lesssim\gamma^{(1)}(G)\lesssim\gamma^{(n)}(G)\cdot\log\Delta.$
The proof that $\gamma^{(1)}(G)\lesssim\gamma^{(n)}(G)\cdot\log\Delta$ uses a random projection and succeeds in expectation, so Algorithm 1 replicates this guarantee with high probability.
Thus, the only remaining inequality in Theorem 1.4 left for us to prove is:
$\gamma^{(1)}(G)\lesssim\gamma^{(n)}(G)\cdot[\alpha(G)]^{2}.$
In order to prove the result, we must also do so in a way that shows Algorithm 2 produces a feasible solution for $\gamma^{(1)}(G)$ obtaining this bound with high probability.
We will do this in several steps. First, we define the following quantity:
Definition 4.3.
For a graph $G$, we define the $d$-dimensional $L^{p}$-extremal embedded spread as:
[…]
For the purposes of this section, $Q^{(1)}_{2}$ will be a bit easier to work with than $\gamma^{(1)}$, and the more general definition will also be useful in a later section when we try to relate $\gamma^{(1)}$ and the $L^{2}$-extremal spread $\overline{s}_{2}$. $Q^{(d)}_{2}(G)$ and $\gamma^{(d)}(G)$ are related as follows:
Lemma 4.4.
For a graph $G$, $\gamma^{(d)}(G)=\frac{2n}{\overline{Q}^{(d)}_{2}(G)}$
Proof.
We will make a sequence of transformations to each optimization problem.
We claim that for any function $f:V\to\mathbb{R}^{d}$ for which $\sum_{v\in V}f(v)=\overline{0}$, it holds that
[…]
To prove this claim, let $f_{i}(v)$ denote the $i$th coordinate of the vector $f(v)$. Then,
[…]
$=\sum_{u,v\in V}\sum_{i\in[d]}\left[f_{i}(u)^{2}+f_{i}(v)^{2}-2f_{i}(u)f_{i}(v)\right]$
$=\sum_{i\in[d]}\sum_{u\in V}\sum_{v\in V}\left[f_{i}(u)^{2}+f_{i}(v)^{2}-2f_{i}(u)f_{i}(v)\right]$
$=\sum_{i\in[d]}\sum_{u\in V}\left[\sum_{v\in V}\left(f_{i}(u)^{2}+f_{i}(v)^{2}\right)-2f_{i}(u)\sum_{v\in V}f_{i}(v)\right]$
$=\sum_{i\in[d]}\sum_{u\in V}\sum_{v\in V}\left[f_{i}(u)^{2}+f_{i}(v)^{2}\right]$
$=\sum_{i\in[d]}2n\sum_{x\in V}f_{i}(x)^{2}$
[…]
Therefore,
[…]
No vector f(u)f(v)f(u)-f(v) is affected by translation, so the constraint $\sum_{v\in V}f(v)=\overline{0}$ is unnecessary, and thus the reciprocal of the inner fractional program here is exactly $\overline{Q}_{2}^{(d)}$. ∎
The key lemma is as follows:
Lemma 4.5.
For any graph $G$, and any $d,p\geq 1$, $\overline{Q}_{p}^{(d)}(G)\leq C_{p}\cdot\left[\alpha(G)\right]^{p}\cdot\overline{Q}_{p}^{(1)}(G),$ where $C_{p}$ is a constant dependent only on $p$. Moreover, given a feasible solution to $\overline{Q}_{p}^{(d)}(G)$ and a padded partition scheme with padding $\alpha$, a corresponding feasible solution to $Q_{p}^{(1)}(G)$ satisfying this inequality can be found using Algorithm 2 with high probability.
Proof.
Let $f,y$ be the optimal solution to $Q_{p}^{(d)}$. Let $\omega(v):=y(v)^{\frac{1}{p}}$. Let $d_{\omega}$ be the shortest-path metric over $G$ that uses $\omega$ as vertex weights, so a path $v_{1},\dots,v_{k}$ has weight $\frac{1}{2}\omega(v_{1})+\omega(v_{2})+\cdots+\omega(v_{k-1})+\frac{1}{2}\omega(v_{k})$. Biswal, Lee, and Rao [BLR10, Theorem 4.4] showed that applying Algorithm 2 (or a generalization of this algorithm for arbitrary $p$) embeds the metric $(V,d_{\omega})$ into $\mathbb{R}$ with a function $f^{\prime}:V\to\mathbb{R}$ so that:
  • For every pair of vertices $u$ and $v$, $|f^{\prime}(u)-f^{\prime}(v)|\leq d_{\omega}(u,v)$.
  • With high probability, $C_{p}^{\prime}\cdot\alpha(V,d_{\omega})^{p}\sum_{u,v\in V}|f^{\prime}(u)-f^{\prime}(v)|^{p}\geq\sum_{u,v\in V}d_{\omega}(u,v)^{p}$, for some positive constant $C_{p}^{\prime}$ depending only on $p$.
To verify that $f^{\prime},y$ forms a feasible solution to $Q_{p}^{(1)}$, consider an edge $uv\in E$. For such an edge, the shortest path is always to use the edge, so $d_{\omega}(u,v)=\frac{1}{2}\left[\omega(u)+\omega(v)\right]$. Therefore, $|f^{\prime}(u)-f^{\prime}(v)|^{p}\leq d_{\omega}(u,v)^{p}\leq\left(\frac{\omega(u)+\omega(v)}{2}\right)^{p}\leq\frac{\omega(u)^{p}+\omega(v)^{p}}{2}<y(u)+y(v)$, where the penultimate inequality follows from Jensen on the convex mapping $x\mapsto x^{p}$.
We now bound the objective value. For each $u,v\in V$, there is some path $u=w_{1},w_{2},\dots,w_{k}=v\in V$ with
$d_{\omega}(u,v)$ $=\frac{1}{2}\omega(w_{1})+\omega(w_{2})+\cdots+\frac{1}{2}\omega(w_{k})$
$=\frac{1}{2}\left[\left(\omega(w_{1})+\omega(w_{2})\right)+\left(\omega(w_{2})+\omega(w_{3})\right)+\cdots+\left(\omega(w_{k-1})+\omega(w_{k})\right)\right].$
[…]
Picking $C_{p}:=2C_{p}^{\prime}$ then gives the stated inequality. ∎
The combination of Lemma 4.4 and Lemma 4.5 therefore completes the proof of Theorem 1.4.
5 Bounds via Geometric Intersection Graphs
In this section, we prove Propositions 5.5, 2.5, 2.6 and 2.7. Each of these is a bound on $\gamma^{(d)}$ for a different graph class (and a few different values of $d$). In particular, each of these bounds uses the relationships these classes have with a few special kinds of $d$-dimensional ball-intersection graphs, and how such ball-intersection graphs relate directly to $\gamma^{(d)}$.
We start by defining a modified form of $\gamma^{(d)}$:
[…]
Lemma 5.1.
For any graph $G$, and any value $d\geq 1$, $\frac{1}{2}\gamma^{(d)}\leq\dot{\gamma}^{(d)}(G)\leq\gamma^{(d)}(G)$.
Proof.
[…]
The special conditions we need for our ball-intersection graphs all involve the concept of ply: For a set of balls $\mathcal{B}$ in $\mathbb{R}^{d}$, a point $p\in\mathbb{R}^{d}$ is said to have ply $k$ if there are exactly $k$ balls in $\mathcal{B}$ that contain $p$. Then, the ply of $\mathcal{B}$ itself is said to be the maximum ply value of any point not contained in the boundary of some ball. Note that, due to this boundary condition, sets of balls with even ply $1$ can still contain intersections. For planar graphs and the geometric graph classes, we will be able to work with bounded ply for the full set of balls. However, for bounded-genus graphs, we instead will use set of balls containing a very large subset that has bounded ply. See Figure 2 for some examples of balls with bounded ply, and their corresponding geometric intersection graphs. A set of $2$-dimensional balls with ply $1$ is also called a (univalent) circle packing.
Figure 2: Two examples sets of balls, and the corresponding geometric intersection graphs. The left example has ply $1$, while the right example has ply $2$.
It will also be helpful for the results in this section to note the following:
Lemma 5.2.
[OTZ22, Proof of Proposition 2.9]. For a graph $G$ and any value $d\geq 1$, $\gamma^{(1)}(G)\lesssim d\gamma^{(d)}(G)$. Moreover, given a feasible solution to $\gamma^{(d)}(G)$, a feasible solution to $\gamma^{(1)}(G)$ satisfying this inequality can be found in polynomial time.
5.1 Planar Graphs
Planar graphs admit a very well-structured representation as a ball-intersection graph. This comes from the celebrated circle packing theorem:
Theorem 5.3 (Planar Circle Packing Theorem/Koebe-Andreev-Thurston Theorem [Koe36, And70, TM79]).
Let $G$ be a simple undirected graph. Then $G$ is the geometric intersection graph of $2$-dimensional balls with ply $1$ (a circle packing) if and only if $G$ is a planar graph. Moreover, if $G$ is a maximal planar graph, then this representation is unique up to Möbius transformations.
Spielman and Teng [ST96, ST07] made use of this theorem to bound $\lambda_{2}(G)$ for planar graphs $G$, hence giving a “spectral” proof of a weaker version of the planar separator theorem via Theorem 2.1. We will use a similar argument to bound $\gamma^{(1)}(G)$, hence also implying a “spectral”-type of proof of the planar separator theorem in full generality via Theorem 1.4. We will make use of one key result of Spielman and Teng to accomplish this:
Theorem 5.4 ([ST96, ST07, Theorem 4.5]).
Let $B_{1},\dots,B_{n}$ be a collection of balls in $\mathbb{R}^{d}$ with centres $c_{1},\dots,c_{n}$, so that no point $x\in\mathbb{R}^{d}$ is contained in $\left\lceil\frac{n}{2}\right\rceil$ of the balls. Then there is a homeomorphism $\alpha$ from $\mathbb{R}^{d}$ to a subset of the sphere $S^{d}$, [^4] We use the standard topological convention of $S^{d}$ as the $d$-dimensional surface of the $(d+1)$-dimensional unit ball embeddable in $\mathbb{R}^{d+1}$. This differs from the notation of Spielman and Teng, who used $S^{d}$ to denote the (d1)(d-1)-dimensional surface of the $d$-dimensional unit ball. so that $\alpha(B_{i})$ is exactly a geodesic ball (or sphere cap) in $S^{d}$ with center $\alpha(c_{i})$, and moreover so that the centroid of the values $\alpha(c_{1}),\dots,\alpha(c_{n})$, in the natural representation of $S^{d}$ as surface of the unit ball of $d+1$ dimensions, is exactly the origin.
This theorem statement is slightly weaker than the statement used by Spielman and Teng, who also described the structure of the homeomorphism as a stereographic projection. However, this statement is sufficient for our purposes. In particular, it gives us a method of normalizing a geometric intersection graph of balls in $\mathbb{R}^{d}$ to be a geometric intersection graph of (geodesic) balls in the sphere $S^{d}$. Importantly, since the theorem gives a homeomorphism, it preserves the ply of each individual point. From this, we can now obtain the desired bound on $\gamma^{(1)}$ for planar graphs:
Proposition 5.5 (Extended restatement of Proposition 2.4).
Let $G$ be a planar graph with $n$ vertices and maximum degree $\Delta$. Then,
$\gamma^{(1)}(G)\lesssim\gamma^{(3)}(G)\leq\frac{8}{n}.$
Hence, $\psi(G)^{2}\lesssim\frac{1}{n}$.
The method to prove this will be analogous to a proof of Spielman and Teng [ST96, ST07, proof of Theorem 3.3].
Proof.
If we can show that $\gamma^{(3)}(G)\leq\frac{8}{n}$, the other inequalities in the statement follow from previous theorem statements.
By Theorem 5.3 and Theorem 5.4, there exists a representation of $G$ as a circle packing on the sphere $S^{2}\subset\mathbb{R}^{3}$. Let the centers be given by $f:V\to S^{2}\subset\mathbb{R}^{3}$ (i.e., $f(v)||_{2}^{2}=1$ for each $v\in V$). For each $v\in V$, let $s(v)$ be the longest Euclidean distance from $f(v)$ to a point in its geodesic ball on $S^{2}$, embedded in $\mathbb{R}^{3}$. See Figure 3 for an example of these values. Note that the surface area of the geodesic ball is given exactly by $\pi s(v)^{2}$. We claim first that $f,s$ form a feasible solution to $\dot{\gamma}^{(3)}(G)$: The constraint $\sum_{v\in V}f(v)=\overline{0}$ is exactly the statement that the centroid of the centres is the origin, which is given by Theorem 5.4. The other constraints follow from the fact that two balls intersect if and only if they share an edge. Next, we bound the objective value: It follows from the statement of Theorem 5.3 that the ply of all points in $S^{2}$ except for a set of measure 0 is at most $1$. Hence, the total area of all balls is bounded by the area of $S^{2}$ itself, which is $4\pi$. That is, $\sum_{v\in V}\pi s(v)^{2}\leq 4\pi$. Therefore, we get a bound on the objective value of
[…]
Finally, the result follows from Lemma 5.1. ∎
$f(v)$$s(v)$
Figure 3: Illustration of a geodesic ball (sphere cap) on the unit sphere $S^{2}$ embedded in $\mathbb{R}^{3}$, along with its corresponding centre $f(v)$. The distance $s(v)$ is the longest Euclidean distance from $f(v)$ to any point in the ball.
5.2 Ball-Intersection Graphs and Nearest-Neighbour Graphs
Spielman and Teng [ST96, ST07] also considered higher-dimensional geometric graphs with similar properties to planar graphs, and bounded $\lambda_{2}$ for these classes. In particular, they obtained bounds for geometric intersection graphs of $d$-dimensional balls with ply $k$, and $d$-dimensional $k$-nearest-neighbour graphs. We can use similar techniques to prove Propositions 2.5 and 2.6.
Unfortunately, their proof also contains a minor error. Specifically, they used an inequality that does not hold for $d\geq 3$ dimensions: Let $A_{d}$ denote the area of the unit $d$-sphere $S^{d}$, and let $V_{d}$ denote the volume of the unit $d$-ball $B^{d}$ (both of these are $d$-dimensional measures). [^5] We again use the topological convention of the $d$-sphere $S^{d}$ as the $d$-dimensional surface of the $(d+1)$-dimensional unit ball. Let $C$ be a sphere cap (geodesic ball) on the unit $d$-sphere $S^{d}$ embedded in $\mathbb{R}^{d+1}$. Let $f$ be the centre of $C$ on $S^{d}$. Let $r$ be the maximum Euclidean distance from $f$ to a point in $C$. They claimed that $V_{d}r^{d}\leq\text{volume}(B)$. This is true for $d\leq 2$. In fact, we implicitly make use of this fact in the previous subsection. However, we will now provide a counter-example for $d=3$:
Example 5.6.
For $d=3$ dimensions, any sphere cap $C$ that covers strictly less than half the hypersurface of $S^{3}$ has $3$-dimensional surface area strictly less than the volume of the $3$-dimensional ball $V_{3}$ of radius $r$, where $r$ is the maximum Euclidean distance through $\mathbb{R}^{4}$ from the centre of the sphere cap to any other point inside it.
Proof.
Define $h,r,\phi$ as in the $2$-dimensional cross-section of the $4$-dimensional unit ball given in Figure 4.
Figure 4: Cross-section of a unit ball, and the different associated quantities.
The $3$-dimensional area of the $3$-dimensional sphere-cap with colatitude angle $\phi$ (equivalently, height $h$) is $\int_{0}^{\phi}A_{2}(\sin\theta)^{2}d\theta=\frac{A_{2}}{2}(\phi-\cos\phi\sin\phi)$. Note that $A_{2}=4\pi$ is the surface area of the unit $2$-sphere, and that $\cos\phi=1-h$, so $\phi=\arccos(1-h)$. For the volume, $r=\sqrt{2h}$, so $V_{3}\cdot r^{3}=V_{3}\cdot(2h)^{\frac{3}{2}}$. Using, say, $h=0.4$, we can compare the values, and verify that the $3$-dimensional surface area of the sphere cap is <2.81<2.81, while $V_{3}\cdot r^{3}>2.99$. ∎
A weaker form of their bound still holds, fortunately. We will prove and use the following inequality:
Lemma 5.7.
For a sphere cap $C$ on the $d$-dimensional unit sphere embedded in $\mathbb{R}^{d+1}$, let $f$ be its centre and let $r$ be the maximum Euclidean distance in $\mathbb{R}^{d+1}$ from $f$ to a point in $C$. Assume $C$ contains at most half the surface area of the sphere. Then, $d$-dimensional surface area of the sphere cap has $\text{area}(C)\geq\frac{A_{d}r^{d}}{4^{d}}$.
We defer the proof of this inequality to Appendix D, since it is a fairly straightforward expansion of some Gamma and Beta functions. Using it, we can now prove the main result of this subsection:
Proposition 5.8 (Restatement of Proposition 2.5).
Let $G$ be the intersection graph of $d$-dimensional balls with ply $k$ and maximum degree $\Delta$. Then,
$\gamma^{(d+1)}(G)\lesssim\left(\frac{k}{n}\right)^{\frac{2}{d}}.$
Proof.
The proof is very similar to that of Proposition 2.4, and again is based on the methods of Spielman and Teng [ST96, ST07, proof of Theorem 5.1]. By Theorem 5.4, we may realize $G$ as the geometric intersection graph of a set of geodesic balls on $S^{d}$, so that every point in $S^{d}$ is contained in at most $k$ of these geodesic balls almost surely, and so that the centroid of the centres of the geodesic balls in $\mathbb{R}^{d+1}$ is exactly the origin. Denote the geodesic balls (sphere caps) as $C_{1},\dots,C_{n}$. For the geodesic ball $C_{i}$, let $f_{i}$ denote its centre and let $r_{i}$ denote the maximum Euclidean distance in $\mathbb{R}^{d+1}$ from $f_{i}$ to a point in $C_{i}$. By Lemma 5.7, $\text{area}(C_{i})\geq\frac{A_{d}r_{i}^{d}}{4^{d}}$. The total sum of all such areas is bounded above by $kA_{d}$, so $\sum_{i=1}^{n}r_{i}^{d}\leq k4^{d}$. By the power-mean inequality, we get $\frac{\sum_{i=1}^{n}r_{i}^{2}}{n}\leq\left(\frac{\sum_{i=1}^{n}r_{i}^{d}}{n}\right)^{\frac{2}{d}}\leq\left(\frac{k4^{d}}{n}\right)^{\frac{2}{d}}\leq 16\left(\frac{k}{n}\right)^{\frac{2}{d}}$. ∎
As previously mentioned, we will also apply this result to $k$-nearest neighbour graphs, again in a similar manner to Spielman and Teng [ST96, ST07, Corollary 5.2]. Let $\tau_{d}$ denote the kissing number in $d$ dimensions, which is the maximum number of non-overlapping unit balls in $\mathbb{R}^{d}$ arranged to all touch a central unit ball. It is known that as $d\to\infty$, $2^{0.2075d(1+\operatorname{o}(1))}\leq\tau_{d}\leq 2^{0.401d(1+\operatorname{o}(1))}$ [Wyn65, KL78]. Hence, $\log\tau_{d}\gtrsim d$ and $\tau_{d}^{\frac{1}{d}}\lesssim 1$. Miller, Teng, Thurston, and Vavasis [MTTV97] observed that every $k$-nearest neighbour graph is the subgraph of an intersection graph of balls with ply $\tau_{d}k$, and moreover that every $k$-nearest neighbour graph has maximum degree bounded by $\tau_{d}k$. Hence, we obtain the following corollary:
Corollary 5.9 (Restatement of Corollary 2.6).
Let $G$ be a $d$-dimensional $k$-nearest neighbour graph with $n$ vertices. Then,
$\gamma^{(d+1)}(G)\lesssim\left(\frac{k}{n}\right)^{\frac{2}{d}}.$
Note that the weaker bound on $\frac{|\partial S|}{|S|}$ stated in Corollary 1.8 for this class comes from the inequality $\log\Delta\lesssim\log\tau_{d}k\lesssim d+\log k$.
We will now give specialized algorithms for these graph classes, since the implied separator results are new.
Corollary 5.10 (Restatement of Corollary 2.10).
Let $G$ be the intersection graph of $d$-dimensional balls with ply $k$ and maximum degree $\Delta$, provided as a set of $d$-dimensional coordinates and radii for each vertex. Then, in polynomial time, we can compute a balanced $2/3$-vertex-separator of size
$\operatorname{\mathcal{O}}\left(\sqrt{\min\{d,\log\Delta\}}\cdot\left(\frac{k}{n}\right)^{\frac{1}{d}}\right).$
Proof.
Instead of applying Theorem 1.4, perform the stereographic projection implicit in the proof of Theorem 5.4, and then directly apply Theorem 4.2 with Algorithm 1, as well as Lemma 5.2 (taking the best of the two). Finally, apply Theorem 4.1 with Algorithm 4. ∎
Corollary 5.11 (Restatement of Corollary 2.11).
Let $G$ be a $d$-dimensional $k$-nearest neighbour graph with maximum degree $\Delta$, provided as a set of $d$-dimensional coordinates and a value $k$. Then, in polynomial time, we can compute a balanced $2/3$-vertex-separator of size
$\operatorname{\mathcal{O}}\left(\sqrt{\min\{d,\log\Delta\}}\cdot\left(\frac{k}{n}\right)^{\frac{1}{d}}\right).$
Proof.
We start by computing the graph itself. By using a simple construction of Kwok, Lau, and Tung [KLT22, Proof of Lemma 3.12], we can use the coordinates to obtain radii for a representation of the graph as a ball-intersection graph, while optimizing the objective function up to constant factors (relative to the fixed choice of coordinates). Then, instead of applying Theorem 1.4, perform the stereographic projection implicit in the proof of Theorem 5.4, and directly apply Theorem 4.2 with Algorithm 1, as well as Lemma 5.2 (taking the best of the two). Finally, apply Theorem 4.1 with Algorithm 4. ∎
Another result of Spielman and Teng for a graph class they call $\alpha$-overlap graphs also relied on the incorrect inequality for the area of sphere caps [ST07, Theorem 6.4]. Lemma 5.7 may also be applied to fix that result in essentially the same manner. The result could be generalized to a bound on $\gamma^{(1)}$ as well.
5.3 Bounded-Genus Graphs
In this subsection, we will prove Proposition 2.7, which states that a genus-$g$ graph $G$ with maximum degree $\Delta$ has $\gamma^{(1)}(G)\lesssim\frac{g\log\Delta}{n}$.
This will involve several steps. At a high-level, we will use two different groups of techniques:
  • Techniques involving uniform shallow-minors, which we will define shortly.
  • Techniques based on the theory of circle packings for genus-$g$ graphs, adapting constructions used by Kelner to give a bound on the Fiedler value $\lambda_{2}$ for triangulated genus-$g$ graphs [Kel04, Kel06b, Kel06a].
We will now briefly outline some important definitions and challenges related to genus-$g$ graphs. Recall that a genus-$g$ graph is one that can be drawn on a genus-$g$ (oriented) surface without crossing edges. Such a drawing is called a surface embedding, and it implies the existence of a rotation system: A (circular) ordering of the edges around each vertex. Like a plane graph, we can define a face of a genus-$g$ graph as any cycle that is formed by a tour induced by always moving counter-clockwise in the rotation system at a vertex to choose the next edge, starting from the previous. A rotation system (and a corresponding embedding) is called cellular if each edge is a part of exactly two such tours. Not every surface embedding is cellular, even for plane graphs: A face on a plane graph can include an edge twice, each time traversed in opposite directions. For instance, this is always true of cut edges in any embedding of a planar graph (see Figure 5). In a surface embedding for a genus-$g$ graph with $g\geq 1$, it can also be the case that an edge is included twice in the same direction (see Figure 5).
Figure 5: Non-cellular embeddings: planar graph (left) and $K_{5}$ in the flat torus (right). A non-cellular face in each is traced.
For simple graphs, we call a graph embedded in a surface with cellular faces a triangulation if all of its faces consist of exactly $3$ edges. Planar triangulations correspond exactly to maximal planar graphs, which are graphs for which the addition of any edge would enforce the invalidation of planarity, regardless of embedding. For graphs of genus-$g$, this is no longer true:
Example 5.12.
$K_{5}$ is a maximal graph of genus $1$, and $K_{5}$ admits no triangulated embedding into a surface of genus $1$.
See Figure 5 for an embedding of $K_{5}$ in the torus.
Proof.
The proof is by contradiction. Suppose this was not the case, and that it had a triangulated embedding in the torus. $K_{5}$ has $5$ vertices and $10$ edges. By the Euler characteristic, an embedding of $K_{5}$ in the torus must have $5$ faces. We will now count the number of edge-face incidences in two different ways. Since the embedding is triangulated (and hence cellular), each edge is incident to exactly two faces, and there are $10$ edges and hence $2\cdot 10=20$ incidences in total. Since the embedding is triangulated, there are exactly $3$ edges incident to each face, and hence there are $3\cdot 5=15$ incidences in total, giving a contradiction. ∎
Kelner claimed to have bounded $\lambda_{2}$ for all genus-$g$ graphs, rather than just triangulations. The method he used to try and prove this was to reduce the case of general genus-$g$ graphs to the case of triangulated genus-$g$ graphs. However, the reduction used for this is actually incorrect. The claimed argument is as follows:
  • Adding edges to a graph can only increase $\lambda_{2}$ (this is also true for $\gamma^{(1)}$).
  • Add edges until the graph is maximal.
  • Apply the bound on triangulated genus-$g$ graphs that Kelner proves in the rest of his paper. This bound relies on a generalized circle packing theorem of He and Schramm [HS93] that applies only to simple triangulations of genus-$g$ surfaces.
This argument has two separate issues:
  • As mentioned, a maximal genus-$g$ graph is not always a triangulation.
  • Even in a planar graph, naively adding edges may increase the maximum degree, possibly up to as large as $\Omega(n)$, and the bound that Kelner obtains for triangulations has a dependence on the maximum degree.
It is worth noting that both of these issues can be sidestepped for planar graphs: Maximal planar graphs are always planar triangulations, and there is a known result that triangulates a planar graph while asymptotically preserving the maximum degree [KB92, KB97]. It is possible that these issues could be amended for genus-$g$ graphs by allowing augmentation to obtain a multigraph, but such a construction is not known, and it would need to be checked that Kelner’s construction can be made to work with a multigraph. Instead, we have both generalized and improved Kelner’s bound via our other techniques.
We now move onto the proof. This will involve several phases. Henceforth, we will fix a rotation system for our initial graph $G$. This rotation system need not be cellular, but we will assume that $G$ is simple (as we do throughout this paper). Importantly for our purposes, so long as we carefully respect this rotation system, we will be able to make certain local transformations to a graph without changing its genus.
5.3.1 Uniform Shallow Minors
For the first phase of our transformations, we will make use of the following structure:
Definition 5.13.
Let $G=(V,E)$ be a graph with $n$ vertices, and let $H=(V^{\prime},E^{\prime})$ be a graph with $r\cdot n$ vertices. Suppose there is a function $p:V^{\prime}\to V$ so that $|p^{-1}(v)|=r$ for each $v\in V$, and $p^{-1}(v)$ is a connected subset of vertices with diameter $L$. Suppose furthermore that for each edge $uv\in E^{\prime}$, $p(u)p(v)\in E$. Then we say that $G$ is a uniform shallow minor of $H$ with depth $L$.
See Figure 6 for some examples.
Figure 6: Two examples of uniform shallow minors: A graph that admits $C_{4}$ as a uniform shallow minor with depth 22 (left), and a graph that admits $P_{4}$ as a uniform shallow minor with depth 11 (right).
Usually when studying minors, we start with a graph, and then consider which minors it contains or does not contain. Often, we even “forbid” minors, resulting in classes like $K_{h}$-minor-free graphs. However, we will use uniform shallow minors in the exact opposite sense: Given a graph $G$, we will aim to find a graph $H$ for which $G$ is a uniform shallow minor. In particular, their use is characterized by the following technical lemma:
Lemma 5.14.
Let $G=(V,E)$ be a graph with $n$ vertices, and let $H=(V^{\prime},E^{\prime})$ be a graph with $|V^{\prime}|=r\cdot n$ vertices. Suppose $G$ is a uniform shallow minor of $H$ with depth $L$. Then $\gamma^{(1)}(G)\lesssim r\cdot L\cdot\gamma^{(1)}(H)$.
This lemma will be useful in the following sense: If we have a graph $G$ of genus $g$, we may be able to find a graph $H$ of genus $g$ by performing local transformations at each vertex, so that $G$ is a uniform shallow minor of $H$. In particular, we may be able to find an $H$ with useful properties, such as a reduced maximum degree or faces that are easier to triangulate. Then, a bound on $\gamma^{(1)}(H)$ will imply a bound on $\gamma^{(1)}(G)$, with “loss” $L$.
Figure 7: Examples of the construction in Lemma 5.14 applied to the uniform shallow minors in Figure 6.
Proof.
We start with a trivial case whose elimination we will use later: If $\gamma^{(1)}(H)\cdot r\cdot L\geq 1$, then the result holds since $\gamma^{(1)}(G)\lesssim 1$: We can obtain a solution validating this bound by picking two vertices $u,v$ of the graph, and setting $y(u)=y(v)=2$, f(u)=1,f(v)=1f(u)=1,f(v)=-1, and $y(w)=f(w)=0$ for any $w\not\in\{u,v\}$.
At a high-level, our construction will randomly sample a representative vertex in $H$ for each vertex in $G$, and then look at the edges along the shortest-paths between these representatives. This is visualized in Figure 7.
Let $p:V^{\prime}\to V$ be the function that certifies the depth-$L$ uniform shallow minor. Let $f_{H},y_{H}$ be a solution to $\gamma^{(1)}(H)$. We will construct a solution $f_{G},y_{G}$ to $\gamma^{(1)}(G)$ with random sampling as follows: Let $\pi(v)$ be a random variable that is uniformly chosen over $p^{-1}(v)$, so that each $\pi(v)$ is independent. Let $m(v)$ be samples of all such $\pi(v)$ so that
$\sum_{u\in V}\sum_{v\in V}|f_{H}(m(u))-f_{H}(m(v))|^{2}\geq\operatorname{\mathbb{E}}\left[\sum_{u\in V}\sum_{v\in V}|f_{H}(\pi(u))-f_{H}(\pi(v))|^{2}\right].$
Let $f_{G}(v):=f_{H}(m(v))-\frac{1}{|V|}\sum_{x\in V}f_{H}(m(x))$, and let $y_{G}(v):=2(2L+1)\cdot\sum_{v^{\prime}\in p^{-1}(v)}y_{H}(v^{\prime})$.
We check that the constraints of $\gamma^{(1)}(G)$ are satisfied by $f_{G},y_{G}$. The normalization constraint is satisfied since $\sum_{v\in V}f_{G}(v)=\sum_{v\in V}f_{H}(m(v))-\sum_{x\in V}f_{H}(m(x))=\overline{0}$. For the edge constraints, let $e=uv\in E$. There is a path of length $k\leq 2L+1$ in $H$ from $m(u)$ to $m(v)$ using only vertices in $p^{-1}(\{u,v\})$. Denote this path as $m(u)=v_{1},\dots,v_{k+1}=m(v)$. Then,
$\begin{array}[]{r c l}y_{G}(u)+y_{G}(v)&=&2(2L+1)\left[\sum_{u^{\prime}\in p^{-1}(u)}y_{H}(u^{\prime})+\sum_{v^{\prime}\in p^{-1}(v)}y_{H}(v^{\prime})\right]\\[8.0pt] &\geq&2(2L+1)\left[y_{H}(v_{1})+\cdots+y_{H}(v_{k+1})\right]\\[8.0pt] &\geq&(2L+1)\sum_{i=1}^{k}\left[y_{H}(v_{i})+y_{H}(v_{i+1})\right]\\[8.0pt] &\geq&(2L+1)\sum_{i=1}^{k}\left|f_{H}(v_{i})-f_{H}(v_{i+1})\right|^{2}\\[8.0pt] &\geq&\frac{2L+1}{k}\left[\sum_{i=1}^{k}\left|f_{H}(v_{i})-f_{H}(v_{i+1})\right|\right]^{2}\\[8.0pt] &\geq&\frac{2L+1}{k}\left|f_{H}(v_{1})-f_{H}(v_{k+1})\right|^{2}\\[8.0pt] &=&\frac{2L+1}{k}\left|f_{G}(u)-f_{G}(v)\right|^{2}\\[8.0pt] &\geq&\left|f_{G}(u)-f_{G}(v)\right|^{2},\end{array}$
where the preantepenultimate step follows from Cauchy-Schwarz, the antepenultimate step follows from the triangle inequality, and the last step uses $k\leq 2L+1$. Thus, the edge constraints are satisfied.
It remains to verify the objective. To obtain a lower bound on $\sum_{v\in V}|f_{G}(v)|^{2}$, we first bound the squared distances within each $p^{-1}(v)$. For any vertex $v\in V$ and any two vertices $u^{\prime},v^{\prime}\in p^{-1}(v)$, there is a path $u^{\prime}=v_{1}^{\prime},\dots,v_{l+1}^{\prime}=v^{\prime}$ of length $l\leq L$ contained entirely within $p^{-1}(v)$. By the triangle inequality, Cauchy-Schwarz, and the constraint from $\gamma^{(1)}(H)$,
$\begin{array}[]{r c l}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}&\leq&\left[|f_{H}(v_{1}^{\prime})-f_{H}(v_{2}^{\prime})|+\cdots+|f_{H}(v_{l}^{\prime})-f_{H}(v_{l+1}^{\prime})|\right]^{2}\\[4.0pt] &\leq&l\left[|f_{H}(v_{1}^{\prime})-f_{H}(v_{2}^{\prime})|^{2}+\cdots+|f_{H}(v_{l}^{\prime})-f_{H}(v_{l+1}^{\prime})|^{2}\right]\\[4.0pt] &\leq&L\sum_{i=1}^{l}\left[y_{H}(v_{i}^{\prime})+y_{H}(v_{i+1}^{\prime})\right].\end{array}$
Summing over all pairs $u^{\prime},v^{\prime}\in p^{-1}(v)$ and noting that each $y_{H}(w^{\prime})$ for $w^{\prime}\in p^{-1}(v)$ appears in at most $2|p^{-1}(v)|^{2}$ terms,
$\sum_{u^{\prime}\in p^{-1}(v)}\sum_{v^{\prime}\in p^{-1}(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}\leq 2Lr^{2}\sum_{w^{\prime}\in p^{-1}(v)}y_{H}(w^{\prime}).$
Using this bound, we obtain a lower bound on $\sum_{v\in V}|f_{G}(v)|^{2}$:
$\begin{array}[]{r c l}2|V|\cdot\sum_{v\in V}|f_{G}(v)|^{2}&=&\sum_{u,v\in V}|f_{G}(u)-f_{G}(v)|^{2}\\[12.0pt] &=&\sum_{u,v\in V}|f_{H}(m(u))-f_{H}(m(v))|^{2}\\[12.0pt] &\geq&\operatorname{\mathbb{E}}\left[\sum_{u,v\in V}|f_{H}(\pi(u))-f_{H}(\pi(v))|^{2}\right]\\[12.0pt] &=&\frac{1}{r^{2}}\sum_{u\neq v\in V}\sum_{u^{\prime}\in p^{-1}(u)}\sum_{v^{\prime}\in p^{-1}(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}\\[12.0pt] &=&\frac{1}{r^{2}}\left[\sum_{u^{\prime},v^{\prime}\in V^{\prime}}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}-\sum_{v\in V}\sum_{u^{\prime},v^{\prime}\in p^{-1}(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}\right]\\[12.0pt] &\geq&\frac{1}{r^{2}}\left[2r|V|\sum_{v^{\prime}\in V^{\prime}}|f_{H}(v^{\prime})|^{2}-2Lr^{2}\sum_{v^{\prime}\in V^{\prime}}y_{H}(v^{\prime})\right]\\[12.0pt] &=&\frac{2\sum_{v^{\prime}\in V^{\prime}}|f_{H}(v^{\prime})|^{2}}{r}\left[|V|-Lr\gamma^{(1)}(H)\right]\\[12.0pt] &\geq&\frac{2(|V|-1)\sum_{v^{\prime}\in V^{\prime}}|f_{H}(v^{\prime})|^{2}}{r},\end{array}$
where the antepenultimate step uses $\sum_{u^{\prime},v^{\prime}\in V^{\prime}}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}=2|V^{\prime}|\sum|f_{H}(v^{\prime})|^{2}$ and the bound on within-set distances, the penultimate step uses $\sum y_{H}(v^{\prime})=\gamma^{(1)}(H)\sum|f_{H}(v^{\prime})|^{2}$, and the last step uses the assumption $Lr\gamma^{(1)}(H)\leq 1$ from the beginning of the proof. Thus, $\sum_{v\in V}|f_{G}(v)|^{2}\gtrsim\frac{\sum_{v^{\prime}\in V^{\prime}}|f_{H}(v^{\prime})|^{2}}{r}$.
Finally, we verify the objective:
$\begin{array}[]{r c l}\frac{\sum_{v\in V}y_{G}(v)}{\sum_{v\in V}|f_{G}(v)|^{2}}&=&\frac{2(2L+1)\sum_{v^{\prime}\in V^{\prime}}y_{H}(v^{\prime})}{\sum_{v\in V}|f_{G}(v)|^{2}}\\[12.0pt] &\lesssim&\frac{2(2L+1)\sum_{v^{\prime}\in V^{\prime}}y_{H}(v^{\prime})}{\sum_{v^{\prime}\in V^{\prime}}|f_{H}(v^{\prime})|^{2}/r}\\[12.0pt] &=&2r(2L+1)\cdot\gamma^{(1)}(H)\\[12.0pt] &\lesssim&r\cdot L\cdot\gamma^{(1)}(H).\end{array}$
We will now apply this result to prove the following lemma:
Lemma 5.15 (Restatement of Lemma 2.8).
Let $G$ be a graph with genus $g$, $n$ vertices, and maximum degree $\Delta$. Then there exists a graph $H$ with $n\Delta$ vertices, maximum degree $4$, and genus $g$, so that $\gamma^{(1)}(G)\lesssim\gamma^{(1)}(H)\cdot\Delta\cdot\log\Delta$.
Proof.
We create a set $V^{\prime}$ that contains $\Delta$ vertices for each vertex in $V(G)$. Let $p:V^{\prime}\to V(G)$ denote the mapping of these vertices. Likewise, for each edge-vertex pair $(e,v)\in E\times V$ with $e$ incident to $v$, map $(e,v)$ to some vertex in $p^{-1}(v)$, in such a manner that no two distinct edge-vertex pairs are mapped to the same vertex in $V^{\prime}$. Use this second mapping to construct a set of endpoint-distinct edges $E^{\prime}$. For each $v\in V(G)$, construct a near-perfect binary tree $T_{v}$ over the $\Delta$ vertices in $p^{-1}(v)$ so that some Euler tour of $T_{v}$ contains the vertices mapped to the rotation system around $v$ as a subsequence (it suffices to construct a near-perfect binary search tree over an arbitrary ordering of $p^{-1}(v)$ containing this subsequence). Note that $T_{v}$ will have max-degree $3$ and diameter at most $2\log_{2}\Delta$. Let $E^{\prime\prime}$ be the union over all such trees. Let $H:=(V^{\prime},E^{\prime}\cup E^{\prime\prime})$. Note that the maximum degree of $H$ is $4$, since each vertex in $V^{\prime}$ is incident to at most one edge in $E^{\prime}$ and at most $3$ edges in $E^{\prime\prime}$.
We claim that $H$ is of genus exactly $g$: $G$ is a minor of $H$ (by contracting all edges in $E^{\prime\prime}$), so clearly the genus of $H$ is at least $g$. Furthermore, a rotation system of $G$ can be extended to a rotation system of $H$ since the trees forming $E^{\prime\prime}$ are planar respecting the ordering induced by the rotation system (see Figure 8), so the genus of $H$ is also at most $g$.
Figure 8: Two examples of the degree-reduction construction for a genus-$g$ graph, respecting a rotation system, local to a vertex. The vertex is replaced with $\Delta$ vertices, at least one vertex for each incident edge, and these new vertices are connected with a near-perfect binary tree. In the left example, the displayed vertex is of maximum degree in its graph, while in the right example it is not (resulting in a degree-11 vertex).
It remains to show that $\gamma^{(1)}(G)\lesssim\gamma^{(1)}(H)\cdot\Delta\cdot\log\Delta$. In fact, $G$ is exactly a uniform shallow minor of $H$ with depth $2\log_{2}\Delta$, so we simply apply Lemma 5.14. ∎
We will be able to leverage this lemma in order to turn a bound on $\gamma^{(1)}$ for bounded-genus graphs of constant maximum degree into a bound on $\gamma^{(1)}$ for bounded-genus graphs of arbitrary maximum degree at only a $\log\Delta$ loss. We will now use a similar technique to further reduce to the case of triangulated bounded-genus graphs with constant maximum degree:
Lemma 5.16 (Restatement of Lemma 2.9).
Let $G$ be a graph of genus $g$ with $n$ vertices and maximum degree $\Delta$. Then there is a triangulated genus-$g$ graph $H$ with $(\Delta+1)\cdot n$ vertices and maximum degree $\operatorname{\mathcal{O}}(\Delta)$, so that $\gamma^{(1)}(G)\lesssim(\Delta+1)\cdot\gamma^{(1)}(H)$.
Proof.
We will assume for simplicity that $G$ is connected. $O(1)$ edges can be added to each vertex to make this the case if it is not true.
Figure 9: A demonstration of the triangulation steps on a face. From left to right, top to bottom: $G$, $H_{1}$, $H_{2}$, $_{3}, HH$. The same steps can also be performed on a non-cellular face.
Let $H_{1}$ be the graph that adds $\Delta$ vertices around each vertex to $G$, and ensure that between any two edges in the rotation system around a vertex in $G$, at least one such edge is added. Observe that $H_{1}$ is a genus-$g$ graph with $(\Delta+1)n$ vertices and maximum degree $2\Delta$. Observe also that $G$ is a uniform shallow minor of $H_{1}$ of depth $2$, and hence Lemma 5.14 shows that $\gamma^{(1)}(G)\lesssim(\Delta+1)\gamma^{(1)}(H_{1})$.
Now, we will add edges to $H_{1}$. In particular, we will construct a sequence of graphs that add edges to $H_{1}$, but never add vertices, and hence $\gamma^{(1)}$ will only increase for this sequence of graphs. Specifically, within each face of $H_{1}$, we create a cycle of all the newly added vertices in order around the face. The result is a graph we call $H_{2}$ (with genus $g$, $(\Delta+1)n$ vertices, and maximum degree $\max\{3,2\Delta\}$) that has no adjacent pair of faces both with size >4>4. Moreover, any vertex incident to a face of size >4>4 (i.e., those in the created cycles) is incident exactly two other faces, and has degree exactly $3$. For each face of exactly size $4$ in $H_{2}$, triangulate it arbitrarily to get another new graph $H_{3}$ (with genus $g$, $(\Delta+1)n$ vertices, and maximum degree $\max\{5,4\Delta\}$). See Figure 9 for examples of each of these graphs around a particular face. It’s worth noting that for non-cellular embeddings of higher genus graphs, it is possible that a non-simple face may include the same edge twice in the same “direction” during a traversal around the face (as opposed to the example in the figure, where the doubly-included edge is included in opposite directions). Fortunately, this does not cause issues for the construction described here either. It is also worth noting that these steps are necessary even if all the faces are simple: Specifically, we would like to handle the remaining non-triangular faces independently, and the separation ensures that we are not able to later produce parallel edges (see [KB97, Figure 6.3(b)] for a demonstration of the issues of adjacent simple faces).
It remains to triangulate the remaining non-triangular faces of $H_{3}$, none of which are adjacent, and all of which are simple. This can be accomplished by adding at most $2$ edges per vertex. In particular, Kant and Bodlaender [KB92, KB97] gave an algorithm called “zig-zagging” that adds edges to a single face using a sequence of ear-cuts, triangulating the face. This algorithm applies in the case of faces on any rotation system, not just a rotation system on a planar graph. Call the final graph $H$, which is a triangulated genus-$g$ graph with maximum degree $\max\{7,4\Delta\}$. See the last diagram in Figure 9 for a demonstration of this algorithm.
Since $H$ has the same vertex set as $H_{1}$, and a strict superset of its edges, $\gamma^{(1)}(H)\geq\gamma^{(1)}(H_{1})$, so the lemma follows from the earlier argument. ∎
5.3.2 Circle Packings for Bounded-Genus Graphs
Because of the lemmas in Section 5.3.1 we henceforth need only consider triangulated genus-$g$ graphs of constant degree. We will now begin generalizing the arguments of Kelner [Kel04, Kel06b, Kel06a] to obtain bounds on $\gamma^{(1)}(G)$ for this class.
A genus-$g$ graph naturally embeds into a genus-$g$ surface. Moreover, a genus-$g$ graph in fact has a circle packing on a genus-$g$ surface. Both the surface and packing is unique up to a certain set of transformations [HS93]. However, to bound $\gamma^{(3)}$ in a similar manner to Proposition 2.4, we would need a circle packing on the unit sphere $S^{2}$, likely of ply $O(g)$. Unfortunately, as Kelner [Kel04, Kel06b, Kel06a] pointed out, such circle packings need-not exist. Instead, in order to bound $\lambda_{2}$, Kelner made use of two steps: First, a specific subdivision procedure is applied to the graph, and it is shown that a bound for the graph obtained by a sequence of subdivision procedures also induces a bound on the initial graph. Second, by applying this procedure a sufficient number of times, it is shown that a continuous analogue of the circle packing theory can be used to obtain a bound. We will apply analogues of each of these steps for $\gamma^{(3)}$ (and $\gamma^{(1)}$, which differs only by a constant factor of at most $3$) instead of $\lambda_{2}$.
The subdivision procedure used by Kelner is the hexagonal subdivision of a triangulation $G$, which produces a graph $G^{\prime}$ by bisecting every edge in $E(G)$, and then connecting all three bisection vertices around each resulting face to form a triangle with new edges. See Figure 10 for an example of this operation.
Refer to caption    Refer to caption    Refer to caption
Figure 10: The hexagonal subdivision procedure applied to small example triangulation subgraphs. The left and centre drawings show a single application of the procedure, while the right drawing shows many.
For a triangulation $G$, let $G^{(k)}$ denote the triangulation that results from $k$ successive applications of hexagonal subdivision. The first step is to show that we can relate $\gamma^{(1)}(G)$ and $\gamma^{(1)}(G^{(k)})$:
Lemma 5.17.
Let $G$ be a triangulation of genus $g$ with maximum degree $\Delta$. Then there is some (universal) constant $c$ so that $|V(G)|\cdot\gamma^{(1)}(G)\lesssim\Delta^{c}\cdot|V(G^{(k)})|\cdot\gamma^{(1)}(G^{(k)})$.
We defer the proof of Lemma 5.17 to Appendix C since it is similar to that of Lemma 5.14, and the details are quite technical. Essentially, it will make use of an argument of Kelner [Kel06a, Kel04, Kel06b] that randomly “embeds” $G$ into $G^{(k)}$. The main distinguishing feature compared to the proof of Lemma 5.14 is that this randomized argument includes not just random vertices, but also random paths between vertices.
The second step is to show that a sufficient number of subdivisions allows us to obtain a bound. This step will make use of the following lemma essentially proved by Kelner:
Lemma 5.18 ([Kel06b, Lemma 5.3, Lemma 5.4, proof of Theorem 2.3]).
Let $G$ be a graph of genus $g$. For each $k$, let $F^{(k)}$ denote a genus-$g$ surface on which $G^{(k)}$ admits a circle packing. For each vertex $v\in G^{(k)}$, let $C_{v}$ denote the disk in the packing corresponding to $v$, and let $r_{v}$ denote its radius. Let $A(C)$ and $D(C)$ denote the area (measure) and diameter (longest geodesic path between a point pair), respectively, of a compact and connected region $C$ either in $F^{(k)}$ or $S^{2}$. Then there is a sequence $f^{(k)}:F^{(k)}\to S^{2}$ of analytic maps so that, for any $\varepsilon>0$, there exists a threshold $N$ such that for any $k\geq N$ a partition of $S^{(k)}$ into $S^{(k)}_{1}\cup S^{(k)}_{2}$ exists with the following properties:
  • For any vertex $v\in G^{(k)}$ where $C_{v}\subset S^{(k)}_{1}$,
    $r_{v}^{2}\lesssim D\left(f^{(k)}(C_{v})\right)^{2}\lesssim A\left(f^{(k)}(C_{v})\right).$
  • The projection of $S^{(k)}_{1}$ with $f^{(k)}$ has ply $O(g)$. That is: For each point $p\in S^{2}$,
    $\left|\left(f^{(k)}\right)^{-1}(p)\right|\lesssim g,$
    almost surely.
  • $\sum_{C_{v}:C_{v}\cap S^{(k)}_{2}\neq\emptyset}D\left(f^{(k)}(C_{v})\right)^{2}\lesssim\varepsilon.$
  • The mapping of the centres of each $C_{v}$ ($v\in V(G^{(k)})$) under $f^{(k)}$, for $S^{2}$ embedded in $\mathbb{R}^{3}$ in the standard manner, is exactly the origin.
This is a significant simplification of the sequence of results used by Kelner, tailored for our purposes. We note, for the interested reader, that Kelner gives a complete exposition of the required steps that can be used to prove the above statement only in the journal version of his work [Kel06b] and his thesis [Kel06a, Chapter 11]. Only a high-level outline is given in the original paper [Kel04].
This lemma gives us an analogous approximate circle packing construction for bounded-genus graphs after a sufficient number of hexagonal subdivisions. We will now use it to prove the following lemma:
Lemma 5.19.
Let $G$ be a triangulation of genus $g$. Then, there is some threshold $N\geq 0$ so that for all $k\geq N$, $\gamma^{(1)}(G^{(k)})\lesssim\gamma^{(3)}(G^{(k)})\lesssim\frac{g}{|V(G^{(k)})|}$.
Proof.
Apply Lemma 5.18 with $\varepsilon=O(1)$ to get $N$. Fix any $k\geq N$. Denote $n^{(k)}:=|V(G^{(k)})|$. We will now choose the values $y,f$ for the formulation of $\dot{\gamma}^{(3)}$ to obtain a similar argument to the proof of Proposition 2.4. For each $v\in V(G^{(k)})$, choose $s(v):=D(f^{(k)}(C_{v}))$. Choose $f(v)$ as the image of the centre of $C_{v}$ under $f^{(k)}$. Note that under this choice, $\sum_{v\in V(G^{(k)})}f^{(k)}(v)$ is exactly the origin.
For each $uv\in E(G^{(k)})$,
$f(v)-f(u)||_{2}\leq D(f^{(k)}(C_{v}))+D(f^{(k)}(C_{u}))=s(u)+s(v).$
Hence, all the constraints are satisfied.
Note that $\sum_{v\in V(G^{(k)})}|f^{(k)}(v)|^{2}=n^{(k)}$. Let $V_{1}=\{v\in V(G^{(k)}):C_{v}\subset S_{1}\}$ and $V_{2}=\{v\in V(G^{(k)}):C_{v}\cap S_{2}\neq\emptyset\}$. Then,
$n^{(k)}\cdot\dot{\gamma}^{(3)}(G^{(k)})=\sum_{v\in V(G^{(k)})}D(f^{(k)}(C_{v}))^{2}=\sum_{v\in V_{1}}D(f^{(k)}(C_{v}))^{2}+\sum_{v\in V_{2}}D(f^{(k)}(C_{v}))^{2}$
$\lesssim\sum_{v\in V_{1}}A(f^{(k)}(C_{v}))+\varepsilon\lesssim g\cdot A(S^{2})+\varepsilon\lesssim g,$
so $\dot{\gamma}^{(3)}(G^{(k)})\lesssim\frac{g}{n^{(k)}}$. ∎
5.3.3 Combining the Lemmas
The combination of Lemma 2.8, Lemma 2.9, Lemma 5.17 and Lemma 5.19 together imply the main result:
Proposition 5.20 (Restatement of Proposition 2.7).
Let $G$ be a graph with $n$ vertices, maximum degree $\Delta$, and genus at most $g$. Then,
$\gamma^{(1)}(G)\lesssim\frac{g\log\Delta}{n}.$
Proof.
Let $G^{\prime}$ be the genus-$g$ graph on $n\Delta$ vertices with maximum degree at most $4=O(1)$ produced by Lemma 2.8. Since the maximum degree of $G^{\prime}$ is constant, a triangulated genus-$g$ graph $G^{\prime\prime}$ with $O(n\Delta)$ vertices and constant maximum degree can be obtained by Lemma 2.9 so that $\gamma^{(1)}(G^{\prime})\lesssim\gamma^{(1)}(G^{\prime\prime})$. By Lemma 5.17, and Lemma 5.19 $\gamma^{(1)}(G^{\prime\prime})\lesssim\frac{g}{n\Delta}$, since the degree of $G^{\prime\prime}$ is constant. Hence,
$\gamma^{(1)}(G)\lesssim\gamma^{(1)}(G^{\prime})\cdot\Delta\log\Delta\leq\gamma^{(1)}(G^{\prime\prime})\cdot\Delta\log\Delta\lesssim\frac{g\log\Delta}{n}.$
6 Bounds via $1$-dimensional Extremal $L_{p}$-Embedded Spread
In this section, we will prove Proposition 2.15, which we restate here:
Proposition 6.1 (Restatement of Proposition 2.15).
For a graph $G$ with maximum degree $\Delta$,
$\gamma^{(1)}(G)\lesssim\left[\alpha(G)\right]^{2}\cdot\frac{n^{3}}{\left[\overline{s}_{2}(G)\right]^{2}}.$
When combined with Proposition 2.13 and Proposition 2.14, this will imply the remaining bounds in Theorem 1.5.
To do this, we will use the definition of $L^{p}$-embedded spread, $\overline{Q}_{p}^{(d)}$, given in Definition 4.3. Recall the lemma relating $\overline{Q}_{2}^{(d)}$ and $\gamma^{(d)}$:
Lemma 6.2 (Restatement of Lemma 4.4).
For a graph $G$, $\gamma^{(d)}(G)=\frac{2n}{\overline{Q}^{(d)}_{2}(G)}$.
Given this equivalence, Proposition 2.15 is equivalent to the following lemma:
Lemma 6.3.
Let $G$ be a graph,
$\left[\alpha(G)\right]^{2}\cdot\overline{Q}_{2}^{(1)}(G)\gtrsim\frac{\left[\overline{s}_{2}(G)\right]^{2}}{n^{2}}.$
To prove this, we will apply Cauchy-Schwarz directly to $\overline{s}_{2}(G)$, and then apply a metric embedding to the resulting quantity. The specific metric embedding result we will use was proven by Biswal, Lee, and Rao [BLR10]:
Proposition 6.4 ([BLR10, Theorem 4.4]).
For a metric $(X,d)$ with padded partition modulus $\alpha(X,d)$, there exists an embedding $f:X\to\mathbb{R}$ such that:
  • For every $x,y\in X$, $|f(x)-f(y)|\leq d(x,y)$ (i.e., $f$ is non-expansive).
  • $\alpha(X,d)^{2}\sum_{x,y\in X}|f(x)-f(y)|^{2}\gtrsim\sum_{x,y\in X}d(x,y)^{2}$.
We previously used a more general algorithmic form of this result in the proof of Lemma 4.5.
The first step is given by the following lemma:
Lemma 6.5.
For a graph $G$,
[…]
Proof.
The result follows directly from Cauchy-Schwarz. ∎
The second step is then given by the following lemma:
Lemma 6.6.
For a graph $G$,
[…]
Proof.
[…]
$y(u)+y(v)=\omega(u)^{2}+\omega(v)^{2}\geq\frac{1}{2}\left[\omega(u)+\omega(v)\right]^{2}\geq\frac{1}{2}|f^{\prime}(u)-f^{\prime}(v)|^{2}\geq|f(u)-f(v)|^{2}.$
Thus, $y,f$ forms a feasible solution to $\overline{Q}_{2}^{(1)}(G)$. Finally, for the objective value, it follows that
$\sum_{u,v\in V(G)}\left[d_{G}^{\omega}(u,v)\right]^{2}\lesssim\left[\alpha(G)\right]^{2}\cdot\sum_{u,v\in V(G)}|f^{\prime}(u)-f^{\prime}(v)|^{2}\lesssim\left[\alpha(G)\right]^{2}\cdot\sum_{u,v\in V(G)}|f(u)-f(v)|^{2}.$
Finally, we can prove Lemma 6.3:
Proof of Lemma 6.3.
Combine Lemma 6.5 and Lemma 6.6. ∎
Consequently, we have completed the proof of Proposition 2.15, and therefore we have proved the remaining bounds stated in Theorem 1.5.
7 Conclusion and Future Work
In order to obtain guarantees on the performance of reweighted spectral partitioning, we have provided bounds on $\gamma^{(n)}(G)$ for a number of graph classes. However, very few of them are tight, even up to constant factors. We pose the following open questions:
  • For a graph $G$ of $n$ vertices with genus $g$, is it true that $\gamma^{(n)}(G)\lesssim\frac{g}{n}$? A positive answer would also imply a conjecture of Spielman and Teng [ST96, ST07, Conjecture 1].
  • For a graph $G$ of $n$ vertices with no $K_{h}$-minor, is it true that $\gamma^{(n)}(G)\lesssim\frac{h^{2}}{n}$? If true, this would be an even stronger result than the first question.
  • For a graph $G$ of $n$ vertices with no $K_{h}$-minor, can it be shown that $\gamma^{(1)}(G)\lesssim\gamma^{(n)}(G)\log h$? If true, this would strengthen Theorem 1.4. Such a result would likely involve a slightly different construction than padded partitions.
We also ask if Algorithm 2 could be replaced by an algorithm that does not directly depend on the partition oracle, but still obtains the same guarantees. By analyzing the argument of Biswal, Lee, and Rao [BLR10, Theorem 4.4], it can be shown that it would suffice to obtain a constant-factor pseudo-approximation algorithm for the obnoxious $k$-median problem (specifically with $k=\Omega(n)$), which asks for a set $S$ of $k$ points in an approximate metric space $(X,d)$ (satisfying the triangle inequality up to a constant factor) maximizing $\sum_{x\in X}\min_{s\in S}d(x,s)$. The intuition for this problem is that proximity to some types of facilities may be undesirable (e.g., garbage dumps). Here, “pseudo-approximation” means that it would suffice to obtain a set of $k^{\prime}=\Omega(n)<k$ points that produces a constant-factor approximation for the best solution with $k$ points. This problem is well-studied in practice [LC21, CMD18, Gok20, LG18, BLMN07, SCC25], but has seen only limited study in theory [KD22].
8 Acknowledgements
Thank you to Therese Biedl for pointing out that not every embedded graph in the torus can be simply triangulated, which led to the $K_{5}$ example discussed at the start of Section 5.3.
Thank you also to Lap Chi Lau for his excellent course in 2022, in which the some of the preliminary results of this work began as a course project.
In parallel to this work, and also as a part of the same course, Kam Chuen Tung independently derived essentially the same result as Proposition 5.5, and similar bounds for $\gamma^{(1)}$ based on $L^{2}$-extremal spread for bounded-genus graphs and $K_{h}$-minor-free graphs. These results have appeared in his recent PhD thesis [Tun25, Chapter 7], which also includes some interesting generalizations.
After the initial preprint of this work, Lap Chi Lau notified the author that he was also independently aware of the bound on $\gamma^{(1)}$ for planar graphs stated in Proposition 5.5 prior to either work.
Appendix A Graph Class Definitions
There are a number of graph classes that are relevant to our results. The most notable is the well-known class of planar graphs, which are graphs that can be drawn on the plane without crossing edges. Most of the other classes we will consider generalize planar graphs in some manner.
Definition A.1.
A graph $G$ is said to have genus $g$ if it can be embedded into an orientable surface of genus $g$, and it cannot be embedded into an orientable surface of smaller genus. Planar graphs are exactly those that have genus 0.
Up to homeomorphism, there is only one orientable surface of genus $g$, for any particular $g$. For $g=0$, it is the sphere. The orientable surface of genus $g>0$ can be obtained from the orientable surface of genus g1g-1 by attaching a “handle”. Alternatively, it can be expressed as $g$ “doughnuts” merged together. See Figure 11 for examples of these surfaces, and see Figure 12 for an example of a graph drawn on a genus-$3$ surface.
Figure 11: Examples of (left to right) a genus-$1$ surface, a genus-$2$ surface, a genus-$3$ surface, and a genus $10$ surface.
Figure 12: An example of a graph (left) and a drawing of the graph on a genus-$3$ surface (right).
Definition A.2.
For a graph $G$ and a graph $H$, we say that $H$ is a minor of $G$ if it can be obtained from $G$ by a sequence of edge deletions, vertex deletions, and edge contractions (that is, merging the two vertices incident to an edge into one). Equivalently, $H$ is a minor of $G$ if there exists a mapping $f:V(G)\to V(H)$ so that if $uv\in E(H)$, then $\left(f^{-1}(u)\times f^{-1}(v)\right)\cap E(G)\neq\emptyset$.
We say a graph $G$ is $H$-minor-free if it does not contain $H$ as a minor.
We will also consider two much more fundamentally geometric graph classes:
Definition A.3.
A graph $G$ is said to be the intersection graph of $d$-dimensional balls with ply $k$ (sometimes called a $d$-dimensional $k$-ply neighbourhood system) if there is a collection of $d$-dimensional balls $B_{1},\dots,B_{n}$ with ply at most $k$ for all points almost surely, so that each ball corresponds to a vertex, and each edge corresponds to a pair of vertices whose balls intersect.
Definition A.4.
A $d$-dimensional $k$-nearest neighbour graph is a graph $G$ with $n$ vertices $v_{1},\dots,v_{n}$ corresponding to points $p_{1},\dots,p_{n}\in\mathbb{R}^{d}$ so that an edge $v_{i}v_{j}$ exists if and only if $p_{i}$ is among the $k$-nearest neighbours of $p_{j}$, or $p_{j}$ is among the $k$-nearest neighbours of $p_{i}$.
Appendix B Vertex Expansion is NP-hard
Although vertex-expansion is widely-known to be NP-hard (see e.g. [LRV13]), we have not been able to find an explicit proof in the literature, so we will provide one here for completeness. Kaibel [Kai04, Theorem 2] faced a similar issue with edge expansion, and presented a proof of hardness of edge expansion based on the hardness of maximum cut.
In this section, we will give a proof of the hardness of vertex expansion based on the hardness of edge expansion. Specifically, we consider the decision problem variants of each, where we ask if there is a subset $S$ of the vertices with edge or vertex expansion at most $\alpha$.
Theorem B.1.
The decision problem form of vertex expansion is NP-hard.
Proof.
We will show that edge expansion can be reduced to vertex expansion. Let $G=(V,E)$ be a graph with $n=|V|$ vertices, and let $\alpha>0$. These form the input to the edge expansion problem, and hence to our reduction. We will construct a graph $G^{\prime}$ and a value $\beta$ so that $G$ has edge expansion at most $\alpha$ if and only if $G^{\prime}$ has vertex expansion at most $\beta$. At a high-level, $G^{\prime}$ will be formed by bisecting each edge of $G$, and then replacing each vertex that was not a bisection vertex with a clique of size $k$. For a sufficiently large value of $k$ (which is still polynomial in $n$), we will be able to show that an upper bound on the vertex expansion of $G^{\prime}$ can be “rounded” to an upper bound on the edge expansion of $G$. In particular, for any $\varepsilon>0$, we may assume $k$ to be a sufficiently large polynomial in $n$ so that for any $S\subset V$, $|S|k\leq|S|k+|E[S]|\leq|S|k+|E|\leq(|S|+\varepsilon)k$, We will place further requirements on our choice of $k$ (including the choice of $\varepsilon$) later.
We now give the explicit construction of $G^{\prime}$ and $\beta$, for a parameter $k$. Let $E[S]$ denote the set of edges in the induced subgraph of $G$ with the vertices $S$. Let $V_{1}$ be a set with $k$ copies of each $v\in V$ labelled $v^{1},\dots,v^{k}$. Let $V_{2}=E$. Let $E_{1}=\{v^{i}v^{j}:v\in V,1\leq i,j\leq k\}$ and $E_{2}=\{ev^{i}:e\in E,v\in V,1\leq i\leq k\}$. Let $G^{\prime}$ be a graph with vertices $V^{\prime}=V_{1}\cup V_{2}$ and edges $E^{\prime}:=E_{1}\cup E_{2}$. We choose $\beta:=\frac{\alpha}{k}$.
We will henceforth use the notation $\partial A:=\{N_{G^{\prime}}(A)\cap V^{\prime}\setminus A\}$ for a set of vertices $A\subset V^{\prime}$, and the notation $\delta B:=E\cap(B\times(V\setminus B))$ for a set of edges $B\subset E$.
Before proving that $\phi(G)\leq\alpha$ if and only if $\psi(G^{\prime})\leq\beta$, we will prove some useful structure.
We first claim that, if there is a subset $S\subset V$, then there is a subset $S^{\prime}\subset V^{\prime}$ so that $|\partial(S^{\prime})|=|\delta(S)|$ and $|S^{\prime}|=|S|k+|E[S]|$, where $E[S]$ denotes the edges in the induced subgraph $G[S]$. In particular, if we let $S^{\prime}$ be the subset of $V^{\prime}$ formed by the union of $E[S]\subset V_{2}$ and $\{v^{i}:v\in S,1\leq i\leq k\}\subset V_{1}$, then we actually have the stronger condition that $\partial(S^{\prime})=\delta(S)$.
We next claim that for any subset $S^{\prime}\subset V^{\prime}$, there are subsets $S^{\prime\prime}\subset V^{\prime}$ and $S\subset V$ so that $\frac{|\partial(S^{\prime\prime})|}{|S^{\prime\prime}|}\leq\frac{|\partial(S^{\prime})|}{|S^{\prime}|}$, $|\delta(S)|=|\partial(S^{\prime})|$ and $|S|k+|E[S]|=|S^{\prime\prime}|$. Specifically, we choose $S^{\prime\prime}$ to be the set that contains, for each $v\in V$, every element of $\{v^{1},\dots,v^{k}\}$ if and only if $S$ contains any element of this set, and furthermore $S^{\prime\prime}$ contains $e=uv\in V_{2}$ if and only if $S$ contains some $u^{i}$ and $v^{j}$. Note that this choice of $S^{\prime\prime}$ corresponds to a set that can be mapped back to sets $S\subset V$ with the desired correspondence via the inverse of the transformation in the previous paragraph.
We now prove that $\frac{|\partial(S^{\prime\prime})|}{|S^{\prime\prime}|}\leq\frac{|\partial(S^{\prime})|}{|S^{\prime}|}$: We can make all the modifications for vertices in $V_{1}$ “first”, and count the changes to the fraction, and then make the modifications for the vertices in $V_{2}$ and count the remaining changes. For the vertices $v^{i}\in V_{1}$ that are in $S^{\prime\prime}$ but not in $S^{\prime}$, they contributed at least $1$ to the numerator and 0 to the denominator under $S^{\prime}$, and under $S^{\prime\prime}$ they contribute 0 to the numerator and $1$ to the denominator. Henceforth, we may assume without loss of generality that $S^{\prime}$ does not “cut” individual sets $\{v_{1},\dots,v_{k}\}$. For each vertex $e=uv\in V_{2}=E$ in $S^{\prime\prime}$ but not in $S^{\prime}$, its neighbours $N_{G^{\prime}}(e)$ are a subset of $S^{\prime\prime}$. Hence, it contributed at least $1$ to the numerator and 0 to the denominator under $S^{\prime}$, and under $S^{\prime\prime}$ they contribute 0 to the numerator and $1$ to the denominator. For the vertices $e=uv\in V_{2}$ in $S^{\prime}$ but not in $S^{\prime\prime}$, such vertices contributed $1$ to the denominator and $k$ to the numerator under $S^{\prime}$, and under $S^{\prime\prime}$ they contribute $1$ to the numerator and 0 to the denominator. We claim that each of these replacements (whose sequence eventually results in $S^{\prime\prime}$) decreases the value of the fraction. Let $a$ be the total size of $|\partial(S^{\prime})|$ excluding a clique of vertices in $V_{1}$, and let $b\geq 1$ be the total size of $|S^{\prime}|$ excluding a bisection vertex in $V_{2}$. We need to prove that $\frac{a+k}{b+1}\geq\frac{a+1}{b}$. If we assume $k\geq n^{2}+n+1$ (which is a polynomial in $n$) then, $a\leq n^{2}$ and $b\leq n$, so $k\geq n^{2}+n+1\implies kb\geq n^{2}+n+1\geq a+b+1\implies ab+kb\geq ab+a+b+1\implies(a+k)b\geq(a+1)(b+1)\implies\frac{a+k}{b+1}\geq\frac{a+1}{b}$.
We will now finally prove that $\phi(G)\leq\alpha$ if and only if $\psi(G^{\prime})\leq\beta=\frac{\alpha}{k}$. In particular, we claim that, for every $\alpha\geq 0$, there exists $S\subset V$ so that $\frac{|\delta(S)|}{|S|}\leq\alpha$ in $G$ if and only if there exists $S^{\prime}$ so that $\frac{|\partial(S^{\prime})|}{|S^{\prime}|}\leq\frac{\alpha}{k}$ in $G^{\prime}$. Note that there are only a polynomial number of possible values of $\frac{|\delta(S)|}{|S|}$ (specifically, there are at most $|E|\cdot|V|$ possible values), so we may assume $\alpha$ is one such value, and that $k$ is chosen so that $\alpha\cdot(1+\varepsilon)$ is strictly less than any larger such value.
In the forward direction, assume $\frac{|\delta(S)|}{|S|}\leq\alpha$. Use $S^{\prime}$ so that $|\partial(S^{\prime})|=|\delta(S)|$ and $|S^{\prime}|=|S|k+|E[S]|$. Then, $\frac{|\partial(S^{\prime})|}{|S^{\prime}|}=\frac{|\delta(S)|}{|S|k+|E[S]|}\leq\frac{|\delta(S)|}{|S|k}\leq\frac{\alpha}{k}$.
In the backward direction, assume $\frac{|\partial(S^{\prime})|}{|S^{\prime}|}\leq\frac{\alpha}{k}$. Then, use the prior construction to find $S$ so that $\frac{|\delta(S)|}{|S|k+|E[S]|}\leq\frac{|\partial(S^{\prime})|}{|S^{\prime}|}\leq\frac{\alpha}{k}$. Recall that $|S|k+|E[S]|\leq(|S|+\varepsilon)k$, so $\frac{|\delta(S)|}{(|S|+\varepsilon)k}\leq\frac{\alpha}{k}$ and thus $\frac{|\delta(S)|}{|S|}\leq\alpha(1+\varepsilon)$. Hence, by the earlier choice of $k$, we also get $\frac{|\delta(S)|}{|S|}\leq\alpha$.
Appendix C Deferred Proof of Subdivision Lemma
We now prove Lemma 5.17 by adapting an argument of Kelner [Kel04, Kel06b, Kel06a].
Lemma C.1 (Restatement of Lemma 5.17).
Let $G$ be a triangulation of genus $g$ with maximum degree $\Delta$. Then there is some (universal) constant $c$ so that $|V(G)|\cdot\gamma^{(1)}(G)\lesssim\Delta^{c}\cdot|V(G^{(k)})|\cdot\gamma^{(1)}(G^{(k)})$.
The proof will be similar to, but not quite the same as, that of Lemma 5.14. The primary differences will be that the paths corresponding to edges will be randomly sampled, and that the “uniformity” is only approximate. It seems quite plausible that a more general form of Lemma 5.14 could be extracted from the below proof, but the most straightforward proof method using these random paths results in an extra factor of $\Delta$. This factor is perfectly fine for proving Lemma 5.17, but poses issues for a generalized form of Lemma 5.14, so we have kept this proof and construction separate.
Proof.
Denote $G=(V,E)$, $n:=|V|$, $H:=G^{(k)}=(V^{\prime},E^{\prime})$, and $n^{\prime}:=|V^{\prime}|$. For the remainder of this proof, we will use the notation $\operatorname{\mathcal{O}}_{\Delta}()$ and $\Theta_{\Delta}()$ to hide polynomial factors in $\Delta$. Note that the maximum degree of $H$ is $\max\{6,\Delta\}$, since no new vertices of degree >6>6 are added from a hexagonal subdivision, nor does any existing vertex have its degree changed. Each edge in $G$ is split into $2^{k}$ pieces in $H$, and each triangle in $G$ is partitioned into $4^{k}$ triangles in $H$. The number of triangles incident to any vertex of degree >6>6 remains constant during subdivision. Hence, $\frac{n^{\prime}}{n}\in\Theta_{\Delta}(4^{k})$. Let $y_{H},f_{H}$ denote the optimal solution to $\gamma^{(1)}(H)$. Assume without loss of generality that $\sum_{x^{\prime}\in V^{\prime}}|f_{H}|^{2}=1$. This assumption will allow us to slightly simplify some later steps.
Kelner [Kel04, Kel06a, Proof of Lemma 5.2] is able to show that there exist random variables $\pi_{V}:V\to V^{\prime}$ and $\pi_{E}:E\to\{\text{paths through $H$}\}$ with the following properties:
  1. 1.
    For each $uv\in E$, $\pi_{E}(uv)$ is a path in $H$ from $\pi_{V}(u)$ to $\pi_{V}(v)$.
  2. 2.
    For each $v\in V$, $\pi_{V}(v)$ is a uniform distribution over its support, which we denote $p(v)$. Moreover, every vertex $v^{\prime}\in V^{\prime}$ is contained in some $p(v)$ for a $v\in V$.
  3. 3.
    For each $u\neq v\in V$, $\pi_{V}(u)$ and $\pi_{V}(v)$ are independent and have disjoint supports.
  4. 4.
    For each $uv\in E$, $\pi_{E}(uv)$ is dependent only on $\pi_{V}(u)$ and $\pi_{V}(v)$.
  5. 5.
    Each path $\pi_{E}(e)$ (for $e\in E$) has length at most $\operatorname{\mathcal{O}}_{\Delta}\left(2^{k}\right)$.
  6. 6.
    Each vertex in $H$ appears in the image of $\pi_{V}$ with probability $\Theta_{\Delta}(1/4^{k})$. That is, $v^{\prime}\in V^{\prime}$ appears in the image of $\pi_{V}(p(v^{\prime}))$ with this probability.
  7. 7.
    Each edge in $H$ appears in the image of $\pi_{E}$ with probability $\operatorname{\mathcal{O}}_{\Delta}(1/2^{k})$, and moreover each edge (and hence also vertex endpoint) in $H$ appears in the support of $\operatorname{\mathcal{O}}_{\Delta}(1)$ random variables $\pi_{E}(e)$ for $e\in E$. Denote the support of edges in $\pi_{E}(e)$ as $p(e)$.
  8. 8.
    Each vertex $v^{\prime}\in V^{\prime}$ appears as an endpoint in the support of some $\pi_{E}(uv)$ only if it is contained in $p(u)\cup p(v)$ (i.e., the vertices incident to elements of $p(uv)$ are a subset of $p(u)\cup p(v)$).
Essentially, this construction is similar to our earlier uniform shallow minors, except that the uniformity is approximate and ignores factors of $\Delta$. Additionally, this construction also uses randomized paths that use only a small number of vertices in expectation. In contrast, our earlier construction for uniform shallow minors used an upper bound that accounted for every vertex in every path.
Using these properties, we obtain that for each pair $u,v\in V$,
$\operatorname{\mathbb{E}}\left[|f_{H}(\pi_{V}(u))-f_{H}(\pi_{V}(v))|^{2}\right]\in\Omega_{\Delta}\left(\frac{1}{16^{k}}\sum_{u^{\prime}\in p(u),v^{\prime}\in p(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}\right),$
and if $uv\in E$ then
$\operatorname{\mathbb{E}}\left[\sum_{u^{\prime}v^{\prime}\in\pi_{E}(uv)}y_{H}(u^{\prime})+y_{H}(v^{\prime})\right]\in\operatorname{\mathcal{O}}_{\Delta}\left(\frac{1}{2^{k}}\sum_{u^{\prime}v^{\prime}\in p(uv)}y_{H}(u^{\prime})+y_{H}(v^{\prime})\right).$
Hence, there exists some deterministic choices $\pi_{V}^{*},\pi_{E}^{*}$ so that
$\frac{\sum_{uv\in E}\sum_{u^{\prime}v^{\prime}\in\pi_{E}^{*}(uv)}y_{H}(u^{\prime})+y_{H}(v^{\prime})}{\sum_{u,v\in V}|f_{H}(\pi_{V}^{*}(u))-f_{H}(\pi_{V}^{*}(v))|^{2}}$ $\leq\frac{\operatorname{\mathbb{E}}\left[\sum_{uv\in E}\sum_{u^{\prime}v^{\prime}\in\pi_{E}(uv)}y_{H}(u^{\prime})+y_{H}(v^{\prime})\right]}{\operatorname{\mathbb{E}}\left[\sum_{u,v\in V}|f_{H}(\pi_{V}(u))-f_{H}(\pi_{V}(v))|^{2}\right]}$
$\in\operatorname{\mathcal{O}}_{\Delta}\left(8^{k}\frac{\sum_{uv\in E}\sum_{u^{\prime},v^{\prime}\in p(uv)}y_{H}(u^{\prime})+y_{H}(v^{\prime})}{\sum_{u\neq v\in V}\sum_{u^{\prime}\in p(u),v^{\prime}\in p(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}}\right)$
$=\operatorname{\mathcal{O}}_{\Delta}\left(8^{k}\frac{\sum_{v^{\prime}\in V^{\prime}}y_{H}(v^{\prime})}{\sum_{u\neq v\in V}\sum_{u^{\prime}\in p(u),v^{\prime}\in p(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}}\right).$
Let $\rho:=\max_{uv\in E}|\pi_{E}^{*}(uv)|$ be the maximum length (in terms of edges) of a sampled path, so $\rho\in\operatorname{\mathcal{O}}_{\Delta}(2^{k})$. Choose $f_{G}(v):=f_{H}(\pi_{V}^{*}(v))-\frac{1}{n}\sum_{x\in V}f_{H}(\pi_{V}^{*}(x))$, and $y_{G}(v):=2\rho\sum_{u:uv\in E}\sum_{x\in p(v)\cap\pi_{E}^{*}(uv)}y_{H}(x)$.
We start by showing that the constraints are satisfied. By definition, $\sum_{v\in V}f_{G}(v)=\overline{0}$. Consider some edge $uv\in E$, and the corresponding path $\pi_{E}^{*}(uv)$ through $H$. By Cauchy-Schwarz and the triangle inequality (similar to the argument in Lemma 4.5),
$y_{G}(u)+y_{G}(v)\geq|f_{G}(u)-f_{G}(v)|^{2},$
so the constraints are satisfied.
It remains only to check the objective value. First, note that
$2n\sum_{x\in V}|f_{G}(x)|^{2}=\sum_{u,v\in V}|f_{G}(u)-f_{G}(v)|^{2}=\sum_{u,v\in V}|f_{H}(\pi_{V}^{*}(u))-f_{H}(\pi_{V}^{*}(v))|^{2}.$
Moreover,
$\sum_{u\neq v\in V}\sum_{u^{\prime}\in p(u),v^{\prime}\in p(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}$
$\sum_{u^{\prime},v^{\prime}\in V^{\prime}}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}-\sum_{v\in V}\sum_{u^{\prime},v^{\prime}\in p(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}.$
For any $v\in V$, any two $u^{\prime},v^{\prime}\in p(v)$, and any $w^{\prime}\in V^{\prime}\setminus p(v)$, the triangle inequality gives $|f_{H}(u^{\prime})-f_{H}(v^{\prime})|\leq|f_{H}(u^{\prime})-f_{H}(w^{\prime})|+|f_{H}(v^{\prime})-f_{H}(w^{\prime})|$, so by Jensen’s inequality on the convex function $x\mapsto x^{2}$, $\frac{1}{2}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}\leq|f_{H}(u^{\prime})-f_{H}(w^{\prime})|^{2}+|f_{H}(v^{\prime})-f_{H}(w^{\prime})|^{2}$. In other words, every squared distance within $p(v)$ can be bounded above by two squared distances crossing out of $p(v)$. So long as there are at least two vertices, there are always more of the latter type (up to factors in $\Delta$), so it follows that
$\sum_{u\neq v\in V}\sum_{u^{\prime}\in p(u),v^{\prime}\in p(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}\in\Omega_{\Delta}\left(\sum_{u^{\prime},v^{\prime}\in V^{\prime}}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}\right).$
Finally,
[…]
$\in\operatorname{\mathcal{O}}_{\Delta}\left(n\rho\cdot 8^{k}\frac{\sum_{v^{\prime}\in V^{\prime}}y_{H}(v^{\prime})}{\sum_{u\neq v\in V}\sum_{u^{\prime}\in p(u),v^{\prime}\in p(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}}\right)$
$=\operatorname{\mathcal{O}}_{\Delta}\left(n\rho\cdot 8^{k}\frac{\sum_{v^{\prime}\in V^{\prime}}y_{H}(v^{\prime})}{\sum_{u^{\prime},v^{\prime}\in V^{\prime}}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}}\right)$
$=\operatorname{\mathcal{O}}_{\Delta}\left(n\rho\cdot 8^{k}\frac{\sum_{v^{\prime}\in V^{\prime}}y_{H}(v^{\prime})}{2n^{\prime}\sum_{v^{\prime}\in V^{\prime}}|f_{H}(v^{\prime})|^{2}}\right)$
$=\operatorname{\mathcal{O}}_{\Delta}\left(4^{k}\frac{\sum_{v^{\prime}\in V^{\prime}}y_{H}(v^{\prime})}{\sum_{v^{\prime}\in V^{\prime}}|f_{H}(v^{\prime})|^{2}}\right).$
Appendix D Deferred Proof of Area Inequality
In this section, we prove Lemma 5.7.
Lemma D.1 (Restatement of Lemma 5.7).
For a sphere cap $C$ on the $d$-dimensional unit sphere embedded in $\mathbb{R}^{d+1}$, let $f$ be its centre and let $r$ be the maximum Euclidean distance in $\mathbb{R}^{d+1}$ from $f$ to a point in $C$. Assume $C$ contains at most half the surface area of the sphere. Then, $d$-dimensional surface area of the sphere cap has $\text{area}(C)\geq\frac{A_{d}r^{d}}{4^{d}}$.
The proof will use three important well-understood functions: The Gamma function $\Gamma$, the Beta function $B$, and the regularized incomplete Beta function $I$. We will not require the explicit representation of the Gamma function, but we will use Gautschi’s inequality:
Proposition D.2 (Gautschi’s inequality [Gau59, DLMF 5.6.4]).
For $x\geq 0,s\in(0,1)$, it holds that
$x^{1-s}<\frac{\Gamma(x+1)}{\Gamma(x+s)}<(x+1)^{1-s}.$
We will use the explicit forms of the Beta and regularized incomplete Beta functions:
Definition D.3.
For $z_{1},z_{2},x>0$, the incomplete Beta function is defined as $B(x;z_{1},z_{2}):=\int_{0}^{x}t^{z_{1}-1}(1-t)^{z_{2}-1}dt$. The Beta function is defined as $B(z_{1},z_{2}):=B(1;z_{1},z_{2})$, and it is also known to be equal to $\frac{\Gamma(z_{1})\Gamma(z_{2})}{\Gamma(z_{1}+z_{2})}$. The regularized incomplete Beta function is defined as $I_{x}(z_{1},z_{2}):=\frac{B(x;z_{1},z_{2})}{B(z_{1},z_{2})}$.
We also have the following important result for our purposes:
Proposition D.4 ([Li11, Equation (1)]).
For a sphere cap $C$ with height $h\leq 1$ of the $d$-dimensional unit sphere, $\text{area}(C)=\frac{1}{2}A_{d}\cdot I_{2h-h^{2}}\left(\frac{d}{2},\frac{1}{2}\right)$.
Note that the convention for dimension differs in the reference.
Proof of Lemma 5.7.
Let $h=\frac{r^{2}}{2}$ be the height of the sphere cap (see Figure 4). Then,
$B\left(2h-h^{2};\frac{d}{2},\frac{1}{2}\right)=\int_{0}^{2h-h^{2}}t^{\frac{d}{2}-1}(1-t)^{-\frac{1}{2}}dt\geq\int_{0}^{2h-h^{2}}t^{\frac{d}{2}-1}dt=\frac{2(2h-h^{2})^{\frac{d}{2}}}{d}.$
Next,
$B\left(\frac{d}{2},\frac{1}{2}\right)=\frac{\Gamma\left(\frac{d}{2}\right)\cdot\Gamma\left(\frac{1}{2}\right)}{\Gamma\left(\frac{d}{2}+\frac{1}{2}\right)}=\frac{\Gamma\left(\frac{d}{2}\right)\cdot\sqrt{\pi}}{\Gamma\left(\frac{d}{2}+\frac{1}{2}\right)}<\sqrt{\pi}\left(\frac{d}{2}-\frac{1}{2}\right)^{-\frac{1}{2}},$
where the last inequality follows from Gautschi’s inequality.
Now, $2h-h^{2}=h(2-h)\geq h$ (since $h\leq 1$), so
$I_{2h-h^{2}}\left(\frac{d}{2},\frac{1}{2}\right)\geq\frac{\frac{2(2h-h^{2})^{\frac{d}{2}}}{d}}{\sqrt{\pi}\left(\frac{d}{2}-\frac{1}{2}\right)^{-\frac{1}{2}}}\geq\frac{2h^{\frac{d}{2}}}{\sqrt{\pi}d}\left(\frac{d}{2}-\frac{1}{2}\right)^{\frac{1}{2}}=\frac{2h^{\frac{d}{2}}}{\sqrt{2\pi}}\frac{\sqrt{d-1}}{d}.$
Finally,
$\text{area}(C)=\frac{1}{2}A_{d}\cdot I_{2h-h^{2}}\left(\frac{d}{2},\frac{1}{2}\right)\geq\frac{1}{2}A_{d}\cdot\frac{2h^{\frac{d}{2}}}{\sqrt{2\pi}}\frac{\sqrt{d-1}}{d}=A_{d}\cdot\frac{r^{d}}{2^{\frac{d}{2}}\sqrt{2\pi}}\frac{\sqrt{d-1}}{d}\geq A_{d}\cdot\frac{r^{d}}{4^{d}},$
where the last inequality holds since $d\geq 2$. ∎